### v1

# The Value of Information in Human-AI Decision-making

Abstract
Humans and AIs are often paired on decision tasks with the expectation of achieving complementary performance, where the combination of human and AI outperforms either one alone. However, how to improve performance of a human-AI team is often not clear without knowing more about what particular information and strategies each agent employs. We provide a decision-theoretic framework for characterizing the value of information—and consequently, opportunities for agents to better exploit available information–in AI-assisted decision workflow. We demonstrate the use of the framework for model selection, empirical evaluation of human-AI performance, and explanation design. We propose a novel information-based instance-level explanation technique that adapts a conventional saliency-based explanation to explain information value in decision making.
1 Introduction
As the performance of artificial intelligence (AI) models improves, workflows in which human and AI model-based judgments are combined to make decisions are sought in medicine, finance, and other domains. Though statistical models often make more accurate predictions than human experts on average (Ægisdóttir et al., 2006; Grove et al., 2000; Meehl, 1954), whenever humans have access to additional information over the AI, there is potential to achieve complementary performance by pairing the two, i.e., better performance than either the human or AI alone. For example, a physician may have access to additional information that may not be captured in tabular electronic health records or other structured data (Alur et al., 2024b).
However, evidence of complementary performance between humans and AI is limited, with many studies showing that human-AI teams underperform an AI alone (Buçinca et al., 2020; Bussone et al., 2015; Green and Chen, 2019; Jacobs et al., 2021; Lai and Tan, 2019; Vaccaro and Waldo, 2019; Kononenko, 2001). A solid understanding of such results is limited by the fact that most analyses of human-AI decision-making focus on ranking the performance of human-AI teams or each individually using measures like posthoc decision accuracy. This approach is problematic for several reasons. First, it does not account for the best achievable performance based on the information available at the time of the decision (Kleinberg et al., 2015; Guo et al., 2024; Rambachan, 2024). Second, it cannot provide insight into the potential for available information to improve the decisions, making it difficult to design interventions that improve the team’s performance.
In contrast, identifying information complementarities that contribute to the maximum achievable decision performance of a human and AI model—such as when one of the agents has access to information not contained in the other’s judgments, or has not fully integrated information available in the environment into their judgments—provides more actionable information for intervening to improve the decision pipeline. For example, if human experts are found to possess decision-relevant information over the AI, we might collect further data to improve the AI model. If the model predictions contain decision-relevant information not contained in human decisions, we might design more targeted explanations to help humans integrate under-exploited information.
We contribute a decision-theoretic framework for characterizing the value of information available within an AI-assisted decision workflow. In our framework, information is considered valuable to a decision-maker to the extent that it is possible to in theory incorporate it into their decisions to improve performance. Specifically, our approach analyzes the expected marginal payoff gain from best case (Bayes rational) use of additional information over best case use of the information already encoded in agent decisions for a given decision problem. Based on the intuition that any information that is used by the agents will eventually reveal itself through variation in their decisions, we identify the value of the information in agent (human, AI, or human-AI) decisions by offering them as a signal to a Bayesian rational decision-maker.
We introduce two metrics for evaluating information value in human-AI collaboration. The first—global human-complementary information value—calculates the value of a new piece of information to an agent over the data-generating distribution. The second—instance-level human-complementary information value—supports analyses of how humans or AI systems use information on a instance-by-instance basis.
We apply the framework by applying it to three decision-making tasks where AI models serve as human decision-making assistants [^1] Code to replicate our experiments is availabel at https://osf.io/p2qzy/?view_only=ec06600d06cd4e59bb6051f992e54c08: chest X-ray diagnosis (Rajpurkar et al., 2018; Johnson et al., 2019), deepfake detection (Dolhansky et al., 2020; Groh et al., 2022), and recidivism prediction (Angwin et al., 2022; Dressel and Farid, 2018). First, we demonstrate its utility in model selection by evaluating how well different AI models complement human decision-makers, showing how even among models with similar accuracy, some models strictly offer more complementary information than others across decision problems. Next, we use our framework to empirically evaluate how providing AI assistance (alongside instance-level features) helps human exploit available information for decision-making. Lastly, we demonstrate use of the framework to design explanations by extending SHAP Lundberg and Lee (2017) to highlight the portion of an AI’s prediction that complements human information.
2 Related work
Human-AI complementarity
Many empirical studies of human-AI collaboration focus on AI-assisted human decision-making for legal, ethical or safety reasons (Bo et al., 2021; Boskemper et al., 2022; Bondi et al., 2022; Schemmer et al., 2022). However, a recent meta-analysis by Vaccaro et al. (2024) finds that on average, human–AI teams perform worse than the better of either humans or AI alone. In response, a growing body of work seeks to evaluate and enhance complementarity in human–AI systems (Bansal et al., 2021b, 2019, a; Wilder et al., 2021; Rastogi et al., 2023; Mozannar et al., 2024b). The present work differs from much of this prior work by approaching human-AI complementarity from the perspective of information value and use, including whether the human and AI decisions provide additional information that is not used by the other.
Evaluation of human decision-making with machine learning
Our work contributes to the development of methods for evaluating decisions of human-AI teams (Kleinberg et al., 2015, 2018; Lakkaraju et al., 2017; Mullainathan and Obermeyer, 2022; Rambachan, 2024; Guo et al., 2024; Ben-Michael et al., 2024; Shreekumar, 2025). Kleinberg et al. (2015) first proposed that evaluations of human-AI collaboration should be based on what information is available at the time of decisions. Our work contributes to definition of Bayesian best-attainable-performance benchmarks (Hofman et al., 2021; Wu et al., 2023; Agrawal et al., 2020; Fudenberg et al., 2022). Closest to our work, Guo et al. (2024) use a rational Bayesian agent faced with deciding between the human and AI recommendations as the theoretical upper bound on expected performance of any human-AI team. This benchmark provides a basis for identifying informational “opportunities” within a decision problem.
Human information in machine learning
One popular approach for developing machine learning models is to incorporate human information or expertise in model predictions (Alur et al., 2024a, b; Corvelo Benz and Rodriguez, 2023; Mozannar et al., 2024a; Bastani et al., 2021; Madras et al., 2018; Raghu et al., 2019; Keswani et al., 2022, 2021; Okati et al., 2021). Corvelo Benz and Rodriguez (2023) propose multicalibration over human and AI model confidence information to guarantee the existence of an optimal monotonic decision rule. Alur et al. (2023) propose a hypothesis testing framework to evaluate the added value of human expertise over AI forecasts. Our work shares the motivation of incorporating human expertise but targets a broader scope by quantifying the information value for all available signals and agent decisions implied in a human–AI decision pipeline, rather than focusing solely on improving model performance.
3 Methodology
Our framework takes input as a decision problem associated with an information model and outputs the value of information in any available signals to any agent, conditioning on the existing information in their decisions within a Bayesian decision theoretic framework. Our framework provides two separate functions to quantify the value of information globally across the data-generating process and locally in a realization drawn from the data-generating process. We also introduce a robust analysis approach to information order, which enables to compare the agent-complementary information in signals for all possible decision problems.
Decision Problem
A decision problem consists of three key elements. We illustrate with an example of a weather decision.
  • A payoff-relevant state $\omega$ from a space $\mathbf{\Omega}$. For example, $\omega\in\mathbf{\Omega}=\{0,1\}=\{\text{no rain},\text{rain}\}$.
  • A decision $d$ from the decision space $\mathbf{D}$ characterizing the decision-maker (DM)’s choice. For example, $d\in\mathbf{D}=\{0,1\}=\{\text{not take umbrella},\text{take umbrella}\}$.
  • A payoff function $S:\mathbf{D}\times\mathbf{\Omega}\to\mathbb{R}$, used to assess the quality of a decision given a realization of the state. For example, $S(d=0,\omega=0)=0,S(d=0,\omega=1)=-100,S(d=1,\omega=0)=-50,S(d=1,\omega=1)=0$, which punishes the DM for selecting an action that does not match the weather.
Information Model
We cast the information available to a DM as a signal defined within an information model. We use the definition of an information model in Blackwell et al. (1951). The information model can be represented by a data-generating model with a set of signals.
  • Signals. There are $n$ “basic signals” represented as random variables $\Sigma_{1},\ldots,\Sigma_{n}$, from the signal spaces $\mathbf{\Sigma}_{1},\ldots,\mathbf{\Sigma}_{n}$. These represent information obtained by a decision-maker, e.g., $\mathbf{\Sigma}_{1}=\{\text{cloudy},\text{not cloudy}\}$, $\mathbf{\Sigma}_{2}\in\{0,\ldots,100\}$ for temprature Celsius, etc. The decision-maker observes a signal, which is a combination of the basic signals, represented as a set $V\subseteq 2^{\{\Sigma_{1},\dots,\Sigma_{n}\}}$. For example, a signal representing a combination of two basic signals $V=\{\Sigma_{1},\Sigma_{2}\}$ observed by the decision-maker might consist of cloudiness $\Sigma_{1}$ and the temperature $\Sigma_{2}$ of the day. Given a signal composed of $m$ basic signals, we write the realization of $V$ as $v=(\sigma_{j_{1}},\dots,\sigma_{j_{m}})$, where the realizations $\sigma_{j_{i}}\in\mathbf{\Sigma}_{j_{i}}$ are sorted by the index of the basic signals $j_{i}\in[n]$. The union $V$ of two signals $V_{1},V_{2}$ takes the set union, i.e., $V=V_{1}\cup V_{2}$. Though $V$ is initially defined as a set of random variables, we will slightly abuse notation $V$ to represent a random variable that is drawn from the joint distribution of the basic signals in it.
  • Data-generating process. A data-generating process is a joint distribution $\pi\in\Delta(\mathbf{\Sigma}_{1}\times\ldots\times\mathbf{\Sigma}_{n}\times\mathbf{\Omega})$ over the basic signals and the payoff-relevant state. However, the DM may only observe a subset $V$ of the $n$ basic signals. We can define the Bayesian posterior belief upon receiving a signal $V=v$ from the data-generating model as
    $\Pr[\omega|v]=\frac{\pi(v,\omega)}{\pi(\omega)}$
    where we slightly abuse notation to write $\pi(v,\omega)$ as the marginal probability of the signal realized to be $v$ and the state being $\omega$ with expectation over unobserved signals.
Information value
Our framework quantifies the value of information in a signal $V$ as the extent to which the payoff could be improved by the ideal use of $V$ over a baseline information set. We suppose a rational Bayesian DM who knows the data-generating process, observes a signal realization, updates their prior to arrive at posterior beliefs, and then chooses a decision to maximize their expected payoff based on the posterior belief. Formally, the rational DM’s expected payoff given a (set of) signal(s) $V$ is
$\mathrm{R}^{\pi,S}(V)=\text{\bf E}_{v\sim\pi}\![{\max_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\Pr(\omega|v)}\![{S(d,\omega)}]}]$
We use $\emptyset$ to represent a null signal, such that $\mathrm{R}^{\pi,S}(\emptyset)$ is the expected payoff of a Bayesian rational DM who has no access to a signal but only uses their prior belief to make decisions. In this case, the Bayesian rational DM will take the best fixed action under the prior, and their expected payoff is:
$\mathrm{R}^{\pi,S}(\emptyset)=\max_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\pi}\![{S(d,\omega)}]$
$\mathrm{R}^{\pi,S}(\emptyset)$ defines the maximum expected payoff that can be achieved with no information. Bayesian decision theory quantifies the information value of $V$ by the payoff improvement of $V$ over the payoff obtained without information.
Definition 3.1.
Given a decision task with payoff function $S$ and an information model $\pi$, we define the information value of $V$ as
$IV^{\pi,S}(V)=\mathrm{R}^{\pi,S}(V)-\mathrm{R}^{\pi,S}(\emptyset)$
We adopt the same idea to define the agent-complementary information values in our framework.
3.1 Agent-Complementary Information Value
Given the above definitions, it becomes possible to measure the additional value that new signals can provide over the information already captured by an agent’s decisions. Here, agent may refer to a human, an AI system, or a human–AI team. The intuition behind our approach is that any information that is used by decision-makers should eventually reveal itself through variation in their behaviors. We recover the information value in agent decisions by offering the decisions as a signal to the Bayesian rational DM. We model the agent decisions as a random variable $D^{b}$ from the action space $\mathbf{D}$, which follows a joint distribution $\pi^{b}\in\Delta(\mathbf{\Omega}\times\mathbf{\Sigma}_{1}\times\ldots\times\mathbf{\Sigma}_{n}\times\mathbf{D})$ with the state and signals. The expected payoff of a Bayesian rational DM who knows $\pi^{b}$ is given by the function:
$\mathrm{R}(D^{b})=\text{\bf E}_{d^{b}\sim\pi^{b}}\![{\max_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\Pr(\omega|D^{b}=d^{b})}\![{S(d,\omega)}]}]$
We seek to identify signals $V$ that can potentially improve agent decisions by analyzing the information value in the combined signal $D^{b}\cup V$ and the information value in $D^{b}$, which we define as agent-complementary information value.
Definition 3.2.
Given a decision task with payoff function $S$ and an information model $\pi$, we define the agent-complementary information value of $V$ on agent decisions $D^{b}$ as
$ACIV^{\pi,S}(V;D^{b})=\mathrm{R}^{\pi,S}(D^{b}\cup V)-\mathrm{R}^{\pi,S}(D^{b})$
If the $ACIV$ of a signal $V$ is low, this means either that the $IV$ of $V$ is low (e.g., it is not correlated with $\omega$), or that the agent has already exploited the information in $V$ (e.g., the agent relies on $V$ to make their decisions such that their decisions correlate with $\omega$ in the same way as $V$ correlates with $\omega$). If, however, the $ACIV$ of $V$ is high, then at least in theory, the agent can improve their payoff by incorporating $V$ in their decision making.
Furthermore, $ACIV$ can reveal complementary information between different types of agents. For instance, if we view AI predictions as $V$ and treat human decisions as the existing agent signal $D^{b}$, a large $ACIV$ indicates that AI predictions add considerable value beyond what humans alone achieve. In the reverse scenario, if human decisions serve as $V$ and AI predictions are $D^{b}$, we can measure how much humans contribute on top of the AI. We demonstrate further usage of $ACIV$ in Section 4 and Section 5.
3.2 Instance-level Agent-Complementary
Instance-level Agent-Complementary Information Value ($ILIV$) evaluates the additional information contributed by a single realization of a signal rather than the entire joint distribution. This finer-grained view is critical for tasks where we need to understand the information value on individual instances, such as when asking whether a human expert should trust an individual prediction from an AI or how to help a human expert understand how the AI model is exploiting information about an instance.
To quantify the information value of a realization $V=v$, we construct $Z_{v}$ as a binary variable indicating whether signal $V$ is realized as $v$, i.e., $Z_{v}=\mathbb{1}[V=v]$. The data-generating model defining the joint distribution of $Z_{v}$, $\pi_{v}\in\Delta(\mathbf{\Omega}\times\{0,1\})$, can be constructed through transforming the original data-generating model $\pi$. Because $Z_{v}$ is a garbling of $V$ (Blackwell et al., 1951), there always exists a a Markov matrix $\Gamma:\mathbf{V}\rightarrow\{0,1\}$ such that the new data-generating process $\pi_{v}$ can be constructed through $\pi_{v}=\Gamma\pi$.
Definition 3.3.
Given a decision task with payoff function $S$ and an information model $\pi$, we define the instance-level agent-complementary information value of the realization $V=v$ as:
$ILIV^{\pi,S}(v;D^{b})=\mathrm{R}^{\pi_{v},S}(D^{b}\cup Z_{v})-\mathrm{R}^{\pi_{v},S}(D^{b})$
This local measure captures how much additional payoff can be gained by incorporating the specific realization $v$ into the agent decisions $D^{b}$. Summing $ILIV$ over all possible realizations of $V$ recovers the global agent-complementary information value ($ACIV$).
Proposition 3.4.
$ACIV^{\pi,S}(V;D^{b})=\sum_{v\in\mathbf{V}}ILIV^{\pi,S}(v;D^{b})$
We apply $ILIV$ to define an information-based explanation technique (ILIV-SHAP) that extends SHAP to explains how the information value of AI predictions complements human decisions $D^{\text{Human}}$ for specific instances. Vanilla SHAP (Lundberg and Lee, 2017) defines a saliency-based explanation with a set of effect variables $\phi_{i}$ representing the influence of the realization of basic signal $\Sigma_{i}=\sigma_{i}$ on the model output $f(v)$, where $v=(\sigma_{1},\ldots,\sigma_{n})$ and $f:\mathbf{V}\rightarrow\mathbb{R}$. Lundberg and Lee (2017) show that the only possible explanation model fulfills properties of local accuracy, missingness and consistency for model $f$ and instance $v$ is following:
Definition 3.5 (SHAP (Lundberg and Lee, 2017)).
$\phi_{i}(f,v)=\sum_{v^{\prime}\subseteq v}\frac{|v^{\prime}|!(n-|v^{\prime}|-1)!}{n!}[g_{f}(v^{\prime})-g_{f}(v^{\prime}\backslash\sigma_{i})]$
where $g_{f}(v^{\prime})$ denotes the expectation of model output conditioned on the subset of signals with $V^{\prime}=v^{\prime}$, i.e., $\text{\bf E}\![{f(V)|V^{\prime}=v^{\prime}}]$.
Instead of quantifying changes in the AI’s raw predictions, ILIV-SHAP clarifies how each feature contributes to the agent-complementary information value of predictions.
Proposition 3.6 (ILIV-SHAP).
An explanation model whose effective variable defined as
$\phi_{i}(ILIV(f(\cdot);D^{b}),v)$
fulfill properties of local accuracy, missingness and consistency for $ILIV$ of $f(\cdot)$ on instance $v$.
Intuitively, when $\phi_{i}(ILIV(f(\cdot);D^{b}),v)$ for a feature $\sigma_{i}$ is relatively large, it means that the model $f(\cdot)$ is extracting information that $D^{b}$ lacks from $\Sigma_{i}$ (or other features with equivalent information to it). Otherwise, when $\sigma_{i}$ makes a small relative contribution, it means that it makes the model ignore information in other features that have higher information value for human decision-makers (such that on average lossing information value).
3.3 Robustness Analysis of Information Order
Ambiguity about the appropriate payoff function is not uncommon in human-AI decision settings due to challenges of eliciting utility functions and potential variance in these functions across decision-makers or groups of instances; e.g., doctors penalize certain false negative results differently when diagnosing younger versus older patients (Mclaughlin and Spiess, 2023). We therefore define the partial order of complementary information value using Blackwell order. Blackwell’s comparison of signals (Blackwell et al., 1951) defines a (set of) signal $V_{1}$ as more informative than $V_{2}$ if $V_{1}$ has a higher information value on all possible decision problems. We identify this partial order by decomposing the space of decision problems via a basis of proper scoring rules  (Li et al., 2022; Kleinberg et al., 2023).
Definition 3.7 (Blackwell Order of Information).
A (set of) signal $V_{1}$ is Blackwell more informative than $V_{2}$ if $V_{1}$ achieves a higher payoff on any decision problem,
$\mathrm{R}^{\pi,S}(V_{1})\geq\mathrm{R}^{\pi,S}(V_{2}),\forall S$
We test the Blackwell order between signals on a basis of proper scoring rules induced from decision problems. The basis is the set of V-shaped scoring rules, parameterized by the kink of the piecewise-linear utility function.
Definition 3.8.
(V-shaped scoring rule) A V-shaped scoring rule with kink $\mu\in(0,\frac{1}{2}]$ is defined as
$S_{\mu}(d,\omega)=\left\{\begin{array}[]{cc}\frac{1}{2}-\frac{1}{2}\cdot\frac{\omega-\mu}{1-\mu}&\text{if }d\leq\mu\\ \frac{1}{2}+\frac{1}{2}\cdot\frac{\omega-\mu}{1-\mu}&\text{else},\end{array}\right.$
When $\mu^{\prime}\in(\frac{1}{2},1)$, the V-shaped scoring rule can be symmetrically defined by $S_{\mu^{\prime}}=S_{1-\mu^{\prime}}(1-y,\omega)$.
Intuitively, the kink $\mu$ represents the threshold belief where the decision-maker switches between two actions. Larger $\mu$ means that the decision-makers will prefer $d=1$ more. The closer $\mu$ is to $0.5$, the more indifferent the decision-maker is to $d=0$ or $d=1$.
Proposition 3.9 shows that if $V_{1}$ achieves a higher information value on the basis of V-shaped proper scoring rules than $V_{2}$, then $V_{1}$ is Blackwell more informative than $V_{2}$. Proposition 3.9 follows from the fact that any best-responding payoff can be linearly decomposed into the payoff on V-shaped scoring rules.
Proposition 3.9 (Hu and Wu 2024).
If
$\mathrm{R}^{\pi,S_{\mu}}(V_{1})\geq\mathrm{R}^{\pi,S_{\mu}}(V_{2}),\forall\mu\in(0,1),$
then $V_{1}$ is Blackwell more informative than $V_{2}$.
Extending this to agent-complementary information value, we say that $V_{1}$ offers a higher complementary value than $V_{2}$ under the Blackwell order if
$ACIV^{\pi,S_{\mu}}(V_{1};D^{b})\geq ACIV^{\pi,S_{\mu}}(V_{2};D^{b}),\forall\mu\in(0,1)$
This definition allows us to rank signals (or sets of signals) universally, without needing to pin down any specific payoff function prior to the analysis. See the use case in Section 4.
4 Experiment I: Model Comparison on Chest Radiographs Diagnosis
We apply our framework to a well-known cardiac dysfunction diagnosis task (Rajpurkar et al., 2018; Tang et al., 2020; Shreekumar, 2025). We apply the framework to analyze how much complementary information value a set of possible AI models offer to human decision-makers.
Figure 1: Information value of all deep-learning models calculated under our framework using Brier score as payoff function.
4.1 Data and Model
We use data from the MIMIC dataset (Goldberger et al., 2000), which contains anonymized electronic health records from Beth Israel Deaconess Medical Center (BIDMC), a large teaching hospital in Boston, Massachusetts, affiliated with Harvard Medical School. Specifically, we utilize chest x-ray images and radiology reports from the MIMIC-CXR database (Johnson et al., 2019) merged with patient and visit information from the broader MIMIC-IV database (Johnson et al., 2023). The payoff-related state, cardiac dysfunction $\omega\in\{0,1\}$, is coded based on two common tests, the NT-proBNP and the troponin, using the age-specific cutoffs from Mueller et al. (2019) and Heidenreich et al. (2022). We use the labels from Irvin et al. (2019) as the human decisions (without AI’s assistance) in the diagnosis task, which is a rule-based tool labeling the symptoms as positive, negative, or uncertain, i.e., $d\in\{+,?,-\}$ [^2] The three symbols represents the encoding we use for signal construction, not an assertion of how radiologists communicate.. We fune-tuned five deep-learning models on the cardiac dysfunction diagnosis task, VisionTransformer (Alexey, 2020), SwinTransformer (Liu et al., 2021), ResNet (He et al., 2016), Inception-v3 (Szegedy et al., 2016), and DenseNet (Huang et al., 2017). Our training set contains 12,228 images and validation set contains 6,115 images. On a hold-out test set with 12,229 images, the AUC achieved by the five models are: DenseNet with $0.77$, Inception v3 with $0.76$, ResNet with $0.77$, SwinTransformer with $0.78$, VisionTransformer with $0.80$.
We consider Brier score as a payoff function, and also conduct a robust analysis considering various V-shaped payoff functions with different kinks on a discretized grid of $[0,1]$ with a step $0.01$. We use the hold-out test set to estimate the data-generating process, which defines the joint distribution of state, human decisions and AI models’ predictions.
4.2 Results
Can the AI models complement human judgment? We first analyze the agent-complementary information values in Figure 1, using the Brier score as the payoff function. We find that all AI models provide complementary information value to human judgment. This highlights the same takeaways as section 5, that the AI model has considerable potential to improve human decisions. As shown in Figure 1 (comparison between $\mathrm{R}^{\pi,S}(D^{\text{Human}}\cup D^{\text{AI}})$ and $\mathrm{R}^{\pi,S}(D^{\text{Human}})$), all AI models capture at least $20\%$ of the total available information value that is not exploited by human decisions. This motivates deploying at AI to assist humans in this scenario.
From the other direction, the human decisions also provide complementary information to all AI models, comparing $\mathrm{R}^{\pi,S}(D^{\text{AI}}\cup D^{\text{AI}})$ with $\mathrm{R}^{\pi,S}(D^{\text{AI}})$ in Figure 1. In scenarios where partial automation is possible, this observation might inspire, for example, further investigation into what information the humans may have access to that is not represented in AI training data.
Which AI model offers the most decision-relevant information over human judgments? Figure 1 shows that VisionTransformer contains slightly higher information value than the other models and Inception v3 contains slightly lower information value than the other models. However, these differences are slight. If the payoff function were questionable in any way, an organization may not want to trust the model rankings. This motivates evaluating model performance over many possible losses to test if there is a Blackwell ordering of models. Across all the V-shaped payoff functions, we find that VisionTransformer is Blackwell more informative and Inception v3 is Blackwell less informative than all other models. By Proposition 3.9, we test the payoff of models on all V-shaped scoring rules, shown in Figure 4. The VisionTransformer achieves a higher information value on all V-shaped scoring rules, implying a higher information value on all decision problems. This analysis highlights the insufficiency of accuracy-based model comparisons to account for all downstream decision problems: 1) while accuracy may rank models in one way, there may exist decision problems where the order between models is reversed; 2) while two models may seem comparable in accuracy, one can be more informative than the other for all decision problems, which is robustly good.
5 Experiment II: Behavioral Analysis on Deepfake Detection
We analyze a deepfake video detection task (Dolhansky et al., 2020), where participants are asked to judge whether a video was created by generative AI. We apply the framework to benchmark the use of available information by the human, AI, and human-AI team.
Figure 2: Information value calculated under our framework in the information model defined by experiment data from Groh et al. (2022). The baseline on the left represents the expected payoff given no information, i.e., $\mathrm{R}^{\pi,S}(\emptyset)$, and the benchmark on the right represents the expected payoff given all available information, i.e., $\mathrm{R}^{\pi,S}(\{\Sigma_{1},\ldots,\Sigma_{n},D^{\text{Human}},D^{\text{Human-AI}},D^{\text{AI}}\})$. All the payoffs are calculated by $\mathrm{R}^{\pi,S}(\cdot)$, where $\cdot$ is the signal on the y-axis.
5.1 Data and Model
We define the information model on the experiment data of Groh et al. (2022). They recruited 5,524 non-expert participants through Prolific. Participants were asked to examine the videos, and provided with assistance from a computer vision model, which achieved $65\%$ accuracy on a holdout dataset. They reported their decisions in two rounds. They first reviewed the video and reported an initial decision ($D^{\text{Human}}$) without access to the model. Then, in a second round, they were told the AI’s recommendation ($D^{\text{AI}}$) and chose whether to change their initial decision, resulting in a final decision ($D^{\text{Human-AI}}$). Participants’ decisions (both initial and final) were elicited as a percentage indicating how confident they were that the video was a deepfake, measured in $1\%$ increments: $d\in\{0\%,1\%,\ldots,100\%\}$. We round the predictions from the AI to the same 100-scale probability scale available to study participants.
We use the Brier score as the payoff function: $S(\omega,d)=1-(\omega-d)^{2}$, with the binary payoff-related state: $\omega\in\{0,1\}=\{\text{genuine},\text{fake}\}$. The scale of Brier score is $[0,1]$ and a random guess ($d\sim\text{Bernoulli}(0.5)$) achieves $0.75$ payoff under Brier score. We choose the Brier score instead of the norm-1 accuaracy used by Groh et al. (2022) because Brier score is a proper scoring rule where truthfully reporting the belief maximizes the payoff [^3] We prefer a proper scoring rule so that the rational decision-maker’s strategy is to reveal their true belief, ensuring that the signal’s information value accurately reflects its role in forming beliefs. If the goal is merely to evaluate the outcome performance of real decision-makers, a non-proper scoring rule may still suffice..
We identify a set of features that were implicitly available to all three agents (human, AI, and human-AI). Because the video signal is high dimensional, we make use of seven seven video-level features using manually coded labels by Groh et al. (2022): graininess, blurriness, darkness, presence of a flickering face, presence of two people, presence of a floating distraction, and the presence of an individual with dark skin, all of which are labeled as binary indictors. These are the basic signals in our framework. We estimate the data-generating process $\pi$ using the realizations of signals, state, first round human decisions, AI predictions, and second round human-AI decisions.
5.2 Results
How much decision-relevant information do each agent’s decisions offer? We first compare the information value of the AI predictions to the decision problem to that of the human decisions in the first round (without AI assistance). To contextualize the value of each, we first construct a scale ranging from the expected payoff of rational DM with no information, i.e., $\mathrm{R}^{\pi,S}(\emptyset)$, to that of the rational agent who has access to all information, i.e., $\mathrm{R}^{\pi,S}(\Sigma_{1}\cup\ldots\cup\Sigma_{n}\cup D^{\text{Human}}\cup D^{\text{Human-AI}}\cup D^{\text{AI}})$. The lower-bound represented by the rational DM with no information is $0.75$ in Brier score, which is equivalent to the payoff achieved by a random guess drawn from $\text{Bernoulli}(0.5)$.
Using this scale, Figure 2(a) shows that AI predictions provide about $65\%$ of the total possible information value over the no information baseline, while human decisions only provide about $15\%$. Hence, human decisions are only weakly informative for the problem.
We next consider the human-AI decisions. Given that the AI predictions contained a significant portion of the total possible information value, we might hope that when participants are provided with AI predictions, their performance comes close to the full information baseline. However, the information value contained in human-AI decisions is also only take a small proportion of the total possible information value ($30\%$). This aligns with the findings by Guo et al. (2024) that humans tend to rely on the AI but are bad at distinguishing when AI predictions are correct.
How much additional decision-relevant information do the available features offer over each agent’s decisions? To understand what information might improve human decisions, we assess the $ACIV$s of different signals over different agents. This describes the additional information value in the signal after conditioning on the exisiting information in the agents’ decisions. As shown on the fifth row in Figure 2, the presence of a flickering face offers larger $ACIV$ over human decisions than over AI predictions, meaning that human decisions could improve by a greater amount if they were to incorporate this information. Meanwhile, as shown on the fourth row in Figure 2, the presence of a individual with dark skin offers larger $ACIV$ over AI predictions than over human decisions. This suggests that the AI and human rely on differing information to make their initial predictions, where the AI relies more on information associated with the presence of a flickering face while human participants rely more on information associated with the presence of an individual with dark skin.
By comparing the $ACIV$s of different signals over human decisions and human-AI decisions, we also find that simply displaying AI predictions to human did not help human-AI to exploit the observed signals in their decisions. As shown in Figure 2, with the assistance of AI, the human-AI teams’ decisions have similar $ACIV$ compared to human decisions (note that the $ACIV$ over human decisions shows significant difference on same signal from the $ACIV$ over AI predictions), except a little improvement on the presence of a flickering face. This finding further comfirms the hypothsis on human simply relying on AI predictions without processing the information contained in them.
6 Experiment III: Information-based Local Explanation on Recidivism Prediction
We apply our framework to a recidivism prediction task, where the decision-maker decides whether to release a defendent. We apply the framework to evaluate and augment saliency-based explanations.
Figure 3: (A) Summary of SHAP explanation and (B) summary of ILIV-SHAP explanation. (C) The waterfall plot of an example of SHAP on instance 4 and (D) the waterfall plot of an example of ILIV-SHAP on instance 4. See more examples of ILIV-SHAP in Appendix C.
6.1 Data and Model
We use the dataset from COMPAS dataset (Angwin et al., 2022), which contains 11,758 defendents with associated features capturing demographics, information about their current charges, and their prior criminal history. We merge the dataset with the experimental data from Lin et al. (2020) to represent the human decisions, which contains recidivism decisions on a subset of instances from Angwin et al. (2022)’s experiment on laypeople [^4] Here the experimental decisions avoid the problem of performative prediction (Perdomo et al., 2020). Human decisions were elicited on a 30-point probability scale. Merging the two datasets produces 9,269 instances (defendants). For the AI model, we trained an extreme gradient boosting (XGBoost) model (Chen and Guestrin, 2016) on a training set with 6,488 instances, achieving an AUC of $0.84$ on a hold-out test set with 2,781 instances. We round the model predictions to the same 30-point probability scale available to study participants in Lin et al. (2020).
We use Brier score as the payoff function, with the binary payoff-related state $\omega$ indicating whether the defendent gets rearrested within two years. We use the features of defendent contained in COMPAS dedaset as the signals in our demonstration: demographic features (age, sex, and race), information about the current charge (type of offense and whether it is a misdemeanor or a felony), prior criminal history (e.g., past number of arrest and charges for several offense categories) and the predicted score from COMPAS system [^5] https://doc.wi.gov/Pages/AboutDOC/COMPAS.aspx. We use the hold-out test set to estimate the data-generating process, which defines the joint distribution of state, signals, human decisions and AI predictions.
Constructing instance-level signals.
Denote the XGBoost model’s predictions as a random variable $Y^{pred}=f(X)$, where $X$ denotes the random variable of data features and $f(\cdot)$ denotes the model’s predictive function. We construct $Z_{f(x)}$ as $\mathbb{1}[Y^{pred}=f(x)]$ for every $x$ and use it to calculate the instance-level agent-complementary information value ($ILIV$) of $Y^{pred}$ on instances $x$ (Definition 3.3). We generate the information-based explanations by ILIV-SHAP in this demonstration.
6.2 Results
How well does SHAP explain complementary information offered by the AI prediction over what humans already know? Figure 3 (A) and (B) compares the distribution of feature-attribution scores from SHAP to those from our ILIV-SHAP. We observe multiple discrepancies in feature importance across the methods. For instance, in Figure 3(A), the age feature negatively correlates with the AI’s prediction, indicating that younger defendants tend to receive higher predictions. However, Figure 3(B) ILIV-SHAP shows no such association between age and the AI’s information value over human decisions. This implies that, although age influences the AI’s prediction, on average it does not provide additional information that humans lack.
Conversely, some features that correlate strongly with the $ILIV$ of the AI prediction are relatively unimportant for the AI’s raw prediction. For example, in Figure 3(A), low decile scores predicted by the COMPAS (colored in blue) have a relatively small impact on the AI’s numerical output compared with other features. In contrast, Figure 3(B) reveals that some low decile scores yield large positive $ILIV$, suggesting that this feature might offer valuable information for human decision-makers despite not significantly changing the AI’s raw prediction.
How can an information-based explanation enhance understanding of AI predictions? We argue that an information-based explanation (via ILIV-SHAP) can serve as a supplement to saliency-based explanations, providing users a sense of whether the individual prediction can help them make decisions. First, an information-based explanation conveys whether the prediction correlates with the payoff-related state. For example, Figure 3(C) shows that the model prediction $f(x)$ deviates $+0.383$ from the prior prediction $\text{\bf E}\![{f(X)}]$ on instance 4. However, the information-based explanation of the same instance in Figure 3(D) shows that the $ILIV$ of the AI prediction over human decisions does not change much from the $ILIV$ of the prior prediction (which is very low at $0.005$). This suggests to the user that focusing on the AI prediction on this instance is not necessary to good decision-making.
Second, the information-based explanation conveys to the user whether the AI gets useful information beyond their own information from a certain feature. In Figure 3(A), the defendent’s prior count record (priors_count=3) changes the AI prediction the most ($+0.186$). However, Figure 3(D) shows that priors_count = 3 contributes marginally to the $ILIV$ ($+0.00004$). Therefore, ILIV-SHAP conveys that even if priors_count=3 changes the AI prediction significantly, the AI does not necessarily better predict the payoff-related state as a result of it.
7 Discussion and Limitations
We propose a decision-theoretic framework for assigning value to information in human-AI decision-making. Our methods quantify the additional information value of any signals over an agent’s decisions. The three demonstrations show how quantified information value can be used in model selection, empricial evaluation, and explanation design. These demonstrations are just a few of many possible uses cases. For example, alternative explanation approaches could be compared via their information value to human decisions. Information value analysis could drive elicitation of human private signals or decision rules to further improve the pairing. New explanation strategies could integrate visualized information value with conventional depictions of feature importance.
Information value cannot definitively establish that particular signals were used by a human. It is always possible that the human has other private signals offering equivalent information to a feature being analyzed. Our framework cannot account for private unobservable signals even though they might have strong correlation with the payoff-related state and agent deicions. However, quantifying the value of observed information is a tool toward learning about information that may exist beyond a problem definition.
Our framework quantifies the best-attainable performance improvement from integrating signals in decisions. This does not necessarily mean that empirically those signals will lead to agents better performing. However, the motivating idea is that Bayesian decision theory provides a theoretical basis that can be adapted to support informative comparisons to human behavior. For example, if we suspect a human decision-maker uses AI predictions and their own predictions strictly monotonically, we could constrain the Bayesian decision-maker to only make monotonic decisions with AI prediction and their own predictions.
Appendix A The Combinatorial Nature of the Value of Signals
We model the information value of a single signal over the existing information in agent decisions. When decision-makers are provided with multiple signals, they might use them in combination. Therefore, our definition of information value in Definition 3.2 may overlook the signals’ value in combination with other signals. Signals can be complemented [Chen and Waggoner, 2016], i.e, they contain no information value by themselves but a considerable value when combined with other signals. For example, two signals $\Sigma_{1}$ and $\Sigma_{2}$ ight be uniformly random bits and the state $\omega=\Sigma_{1}\oplus\Sigma_{2}$, the XOR of $\Sigma_{1}$ and $\Sigma_{2}$. In this case, neither of the signals offers information value on its own, but knowing both leads to the maximum payoff. To consider this complementation between signals, we use the Shapley value $\phi$ [Shapley, 1953] to interpret the contribution to information gain of each basic signal. The Shapley value calculates the average of the marginal contribution of a basic signal $\mathbf{\Sigma}_{i}$ in every combination of signals.
$\phi(\Sigma_{i})=\frac{1}{n}\sum_{V\subseteq\{V_{1},\ldots,V_{n}\}/\{\Sigma_{i}\}}{(n-1)\choose|V|}^{-1}(ACIV(V\cup\{\Sigma_{i}\};D^{b})-ACIV(V;D^{b}))$ (1)
The Shapley value suggests how much information value of the basic signal is unexploited by the human decision-maker on average in all combinations.
The following algorithm provides a polynomial-time approximation of the Shapley value of $ACIV$. Under the assumption of submodularity, it orders the signals the same as the Shapley value.
Algorithm 1 Algorithm 1
1:  $V^{*}=\{D^{b}\}$
2:  $\Phi^{*}=\{\}$
3:  for $i=1$ to $n$ do
4:     $\phi^{\prime}_{j}=ACIV(\Sigma_{j};V^{*})\ for\ each\ j$
5:     $j^{*}=\arg\max_{j\ s.t.\ \Sigma_{j}\notin V^{*}}\phi^{\prime}_{j}$
6:     $\phi_{j^{*}}=\max_{j\ s.t.\ \Sigma_{j}\notin V^{*}}\phi^{\prime}_{j}$
7:     $add\ \Sigma_{j^{*}}\ to\ V^{*}$
8:     $add\ \phi_{j^{*}}\ to\ \Phi^{*}$
9:  end for
10:  $output\ \phi_{j^{*}}$
Appendix B Robustness analysis in Experiment I
Figure 4: Robust analysis for experiment I on all V-shaped payoff functions. The kink $\mu$ is shown on the x-axis. Each subplot displays the difference between the $ACIV$ on the row model over human decisions and the $ACIV$ on the column model over human decisions. The subplots are symmetric along the diagonal, e.g., $(1,2)$ subplot and $(2,1)$ subplot display the same distribution with opposite signs.
Appendix C More examples of ILIV-SHAP
(a) SHAP on instance 0
(b) ILIV-SHAP on instance 0
(a) SHAP on instance 1
(b) ILIV-SHAP on instance 1
(a) SHAP on instance 2
(b) ILIV-SHAP on instance 2

### v3

@@
 Abstract
-Humans and AIs are often paired on decision tasks with the expectation of achieving complementary performance, where the combination of human and AI outperforms either one alone.
-However, how to improve performance of a human-AI team is often not clear without knowing more about what particular information and strategies each agent employs.
-We provide a decision-theoretic framework for characterizing the value of information—and consequently, opportunities for agents to better exploit available information–in AI-assisted decision workflow.
+Multiple agents—including humans and AI models—are often paired on decision tasks with the expectation of achieving complementary performance, where the combined performance of both agents outperforms either one alone.
+However, knowing how to improve the performance of a human-AI team is often difficult without knowing more about what particular information and strategies each agent employs.
+We provide a decision-theoretic framework for characterizing the value of information—and consequently, opportunities for agents to better exploit available information–in AI-assisted decision workflows.
 We demonstrate the use of the framework for model selection, empirical evaluation of human-AI performance, and explanation design.
-We propose a novel information-based instance-level explanation technique that adapts a conventional saliency-based explanation to explain information value in decision making.
+We propose a novel information-based explanation technique that adapts SHAP, a saliency-based explanation to explain information value in decision making.
 1 Introduction
-As the performance of artificial intelligence (AI) models improves, workflows in which human and AI model-based judgments are combined to make decisions are sought in medicine, finance, and other domains.
-Though statistical models often make more accurate predictions than human experts on average (Ægisdóttir et al., 2006; Grove et al., 2000; Meehl, 1954), whenever humans have access to additional information over the AI, there is potential to achieve complementary performance by pairing the two, i.e., better performance than either the human or AI alone.
-For example, a physician may have access to additional information that may not be captured in tabular electronic health records or other structured data (Alur et al., 2024b).
-However, evidence of complementary performance between humans and AI is limited, with many studies showing that human-AI teams underperform an AI alone (Buçinca et al., 2020; Bussone et al., 2015; Green and Chen, 2019; Jacobs et al., 2021; Lai and Tan, 2019; Vaccaro and Waldo, 2019; Kononenko, 2001).
-A solid understanding of such results is limited by the fact that most analyses of human-AI decision-making focus on ranking the performance of human-AI teams or each individually using measures like posthoc decision accuracy.
+As the performance of artificial intelligence (AI) for domain-specific predictions improves, workflows in which human and AI models are combined to make decisions in medicine, finance, and law, among other domains.
+Often, statistical models can be developed that exceed the accuracy of human experts on average (Ægisdóttir et al., 2006; Grove et al., 2000; Meehl, 1954).
+However, whenever humans have access to additional information over the AI, there is potential to achieve complementary performance by pairing the two, i.e., better performance than either the human or AI alone.
+For example, a physician may have access to additional information that may not be captured in tabular electronic health records or other structured data (Alur et al., 2024).
+Evidence of complementary performance between humans and AI nonetheless remains limited, with many studies showing that human-AI teams underperform an AI alone (Buçinca et al., 2020; Bussone et al., 2015; Green and Chen, 2019b; Jacobs et al., 2021; Lai and Tan, 2019; Vaccaro and Waldo, 2019; Kononenko, 2001).
+A solid understanding of such results is limited by the fact that most analyses of human-AI decision-making focus on ranking the performance of human-AI teams or each individually using measures based on post hoc decision accuracy (Passi and Vorvoreanu, 2022).
 This approach is problematic for several reasons.
 First, it does not account for the best achievable performance based on the information available at the time of the decision (Kleinberg et al., 2015; Guo et al., 2024; Rambachan, 2024).
-Second, it cannot provide insight into the potential for available information to improve the decisions, making it difficult to design interventions that improve the team’s performance.
-In contrast, identifying information complementarities that contribute to the maximum achievable decision performance of a human and AI model—such as when one of the agents has access to information not contained in the other’s judgments, or has not fully integrated information available in the environment into their judgments—provides more actionable information for intervening to improve the decision pipeline.
-For example, if human experts are found to possess decision-relevant information over the AI, we might collect further data to improve the AI model.
-If the model predictions contain decision-relevant information not contained in human decisions, we might design more targeted explanations to help humans integrate under-exploited information.
+Second, it cannot provide insight into the potential for available information to improve decision performance, making it difficult to design interventions that improve the team’s performance.
+In contrast, identifying information complementarities that contribute to the maximum achievable decision performance of a human and AI model—such as when one of the agents has access to information not contained in the other’s judgments, or has not fully integrated information available in the environment into their judgments—provides actionable information for improving the decision pipeline.
+For example, the degree to which a model provides information that complements human judgments can guide AI model selection.
+Evidence that human experts possess decision-relevant information over AI predictions motivates further data collection to improve the AI model.
+Evidence that model predictions contain decision-relevant information that humans do not exploit can guide the design of explanations highlighting complementary information.
 We contribute a decision-theoretic framework for characterizing the value of information available within an AI-assisted decision workflow.
-In our framework, information is considered valuable to a decision-maker to the extent that it is possible to in theory incorporate it into their decisions to improve performance.
+In our framework, information is considered valuable to a decision-maker to the extent that it is possible, in theory, to incorporate it into their decisions to improve performance.
 Specifically, our approach analyzes the expected marginal payoff gain from best case (Bayes rational) use of additional information over best case use of the information already encoded in agent decisions for a given decision problem.
-Based on the intuition that any information that is used by the agents will eventually reveal itself through variation in their decisions, we identify the value of the information in agent (human, AI, or human-AI) decisions by offering them as a signal to a Bayesian rational decision-maker.
+The intuition behind this approach is that any information that is used by the agents will eventually reveal itself through variation in their decisions.
+We identify the value of the information in agent (human, AI, or human-AI) decisions by offering them as a signal to a Bayesian rational decision-maker.
+While we focus on combinations of humans and AI agents, the methods we propose can be applied to any combination of agents.
 We introduce two metrics for evaluating information value in human-AI collaboration.
-The first—global human-complementary information value—calculates the value of a new piece of information to an agent over the data-generating distribution.
-The second—instance-level human-complementary information value—supports analyses of how humans or AI systems use information on a instance-by-instance basis.
-We apply the framework by applying it to three decision-making tasks where AI models serve as human decision-making assistants [^1] Code to replicate our experiments is availabel at https://osf.io/p2qzy/?view_only=ec06600d06cd4e59bb6051f992e54c08: chest X-ray diagnosis (Rajpurkar et al., 2018; Johnson et al., 2019), deepfake detection (Dolhansky et al., 2020; Groh et al., 2022), and recidivism prediction (Angwin et al., 2022; Dressel and Farid, 2018).
-First, we demonstrate its utility in model selection by evaluating how well different AI models complement human decision-makers, showing how even among models with similar accuracy, some models strictly offer more complementary information than others across decision problems.
-Next, we use our framework to empirically evaluate how providing AI assistance (alongside instance-level features) helps human exploit available information for decision-making.
-Lastly, we demonstrate use of the framework to design explanations by extending SHAP Lundberg and Lee (2017) to highlight the portion of an AI’s prediction that complements human information.
+The first—global human-complementary information value—calculates the value of a new piece of information to an agent over all of its possible realizations among all instances.
+The second—instance-level human-complementary information value—identifies opportunities for decision-makers to better use instance-level information.
+We demonstrate the framework by applying the metrics to three decision-making tasks where AI models serve as human decision-making assistants [^1] Code to replicate our experiments is available at https://osf.io/p2qzy/?view_only=ec06600d06cd4e59bb6051f992e54c08: chest X-ray diagnosis (Rajpurkar et al., 2018; Johnson et al., 2019), deepfake detection (Dolhansky et al., 2020; Groh et al., 2022), and recidivism prediction (Angwin et al., 2022; Dressel and Farid, 2018).
+First, we demonstrate the framework’s utility in model selection by evaluating how well different AI models complement human decision-makers, showing how even among models with similar accuracy, some models strictly offer more complementary information than others across decision problems.
+Next, we use the framework to empirically evaluate how providing AI assistance (alongside instance-level features) helps humans exploit available information for decision-making.
+Lastly, we demonstrate use of the framework to design explanations by extending SHAP (Lundberg and Lee, 2017) to highlight the portion of an AI’s explanation that complements human information.
 2 Related work
-Human-AI complementarity
-Many empirical studies of human-AI collaboration focus on AI-assisted human decision-making for legal, ethical or safety reasons (Bo et al., 2021; Boskemper et al., 2022; Bondi et al., 2022; Schemmer et al., 2022).
+Human-AI complementarity.
+Many empirical studies of human-AI collaboration focus on AI-assisted human decision-making for legal, ethical, or safety reasons (Bo et al., 2021; Boskemper et al., 2022; Bondi et al., 2022; Schemmer et al., 2022).
 However, a recent meta-analysis by Vaccaro et al.
-(2024) finds that on average, human–AI teams perform worse than the better of either humans or AI alone.
+(2024) finds that, on average, human–AI teams perform worse than the better of the two agents alone.
 In response, a growing body of work seeks to evaluate and enhance complementarity in human–AI systems (Bansal et al., 2021b, 2019, a; Wilder et al., 2021; Rastogi et al., 2023; Mozannar et al., 2024b).
-The present work differs from much of this prior work by approaching human-AI complementarity from the perspective of information value and use, including whether the human and AI decisions provide additional information that is not used by the other.
-Evaluation of human decision-making with machine learning
-Our work contributes to the development of methods for evaluating decisions of human-AI teams (Kleinberg et al., 2015, 2018; Lakkaraju et al., 2017; Mullainathan and Obermeyer, 2022; Rambachan, 2024; Guo et al., 2024; Ben-Michael et al., 2024; Shreekumar, 2025).
+The present work differs from much of this prior work by approaching human-AI complementarity from the perspective of information value and use, including asking whether the human and AI decisions provide additional information that is not used by the other.
+Evaluation of human decision-making with machine learning.
+Our work contributes methods for evaluating the decisions of human-AI teams (Kleinberg et al., 2015, 2018; Lakkaraju et al., 2017; Mullainathan and Obermeyer, 2022; Rambachan, 2024; Guo et al., 2024; Ben-Michael et al., 2024; Shreekumar, 2025).
 Kleinberg et al.
-(2015) first proposed that evaluations of human-AI collaboration should be based on what information is available at the time of decisions.
-Our work contributes to definition of Bayesian best-attainable-performance benchmarks (Hofman et al., 2021; Wu et al., 2023; Agrawal et al., 2020; Fudenberg et al., 2022).
+(2015) proposed that evaluations of human-AI collaboration should be based on the information that is available at the time of decisions.
+According to this view, our work defines Bayesian best-attainable-performance benchmarks similar to several prior works (Guo et al., 2024; Wu et al., 2023; Agrawal et al., 2020; Fudenberg et al., 2022).
 Closest to our work, Guo et al.
-(2024) use a rational Bayesian agent faced with deciding between the human and AI recommendations as the theoretical upper bound on expected performance of any human-AI team.
-This benchmark provides a basis for identifying informational “opportunities” within a decision problem.
-Human information in machine learning
-One popular approach for developing machine learning models is to incorporate human information or expertise in model predictions (Alur et al., 2024a, b; Corvelo Benz and Rodriguez, 2023; Mozannar et al., 2024a; Bastani et al., 2021; Madras et al., 2018; Raghu et al., 2019; Keswani et al., 2022, 2021; Okati et al., 2021).
+(2024) model the expected performance of a rational Bayesian agent faced with deciding between the human and AI recommendations as the theoretical upper bound on the expected performance of any human-AI team.
+This benchmark provides a basis for identifying exploitable information within a decision problem.
+Human information in machine learning.
+Some approaches focus on automating the decision pipeline by explicitly incorporating human expertise in developing machine learning models, such as by learning to defer (Mozannar et al., 2024a; Madras et al., 2018; Raghu et al., 2019; Keswani et al., 2022, 2021; Okati et al., 2021).
 Corvelo Benz and Rodriguez (2023) propose multicalibration over human and AI model confidence information to guarantee the existence of an optimal monotonic decision rule.
@@
 (2023) propose a hypothesis testing framework to evaluate the added value of human expertise over AI forecasts.
-Our work shares the motivation of incorporating human expertise but targets a broader scope by quantifying the information value for all available signals and agent decisions implied in a human–AI decision pipeline, rather than focusing solely on improving model performance.
+Our work shares the motivation of incorporating human expertise, but targets a slightly broader scope by quantifying the information value for all available signals and agent decisions in a human–AI decision pipeline.
 3 Methodology
-Our framework takes input as a decision problem associated with an information model and outputs the value of information in any available signals to any agent, conditioning on the existing information in their decisions within a Bayesian decision theoretic framework.
-Our framework provides two separate functions to quantify the value of information globally across the data-generating process and locally in a realization drawn from the data-generating process.
-We also introduce a robust analysis approach to information order, which enables to compare the agent-complementary information in signals for all possible decision problems.
-Decision Problem
+Our framework takes as input a decision problem associated with an information model and outputs the value of information of any available signals to any agent, conditioning on the existing information in their decisions within a Bayesian decision theoretic framework.
+Our framework provides two separate functions to quantify the value of information: one globally across the data-generating process, and one locally in a realization drawn from the data-generating process.
+We also introduce a robust analysis approach to information order, which enables us to compare the agent-complementary information in signals for all possible decision problems.
+Decision Problem.
 A decision problem consists of three key elements. We illustrate with an example of a weather decision.
@@
 For example, $S(d=0,\omega=0)=0,S(d=0,\omega=1)=-100,S(d=1,\omega=0)=-50,S(d=1,\omega=1)=0$, which punishes the DM for selecting an action that does not match the weather.
-Information Model
+In decision problems corresponding to prediction tasks, the decision space is a probabilistic belief over the state space, i.e., $\mathbf{D}=\Delta(\mathbf{\Omega})$.
+For such problems, a payoff function is said to be a proper scoring rule if the optimal action is to predict the true distribution, i.e., $p=\arg\max_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim p}\![{S(d,\omega)}]$.
+For any decision problem with payoff function $S:\mathbf{D}\times\mathbf{\Omega}\to\mathbb{R}$, there is an equivalent proper scoring rule $\hat{S}:\Delta(\mathbf{\Omega})\times\mathbf{\Omega}\to\mathbb{R}$ defined by choosing the optimal decision under the reported belief.
+Formally,
+$\hat{S}(p,\omega)=S(\arg\max_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim p}\![{S(d,\omega)}],\omega).$ (1)
+Equation 1 shows a reduction from the payoff function $S$ to the proper scoring rule $\hat{S}$, i.e., any decision $p$ under $\hat{S}$ represents a decision $d$ in $S$ that best-responds to the distribution $p$.
+Therefore, the best-attainable performance defined in the proper scoring rule $\hat{S}$ is equivalent to the best-attainable performance defined in any payoff function $S$ that can be reduced to $\hat{S}$.
+Throughout our framework, we prefer proper scoring rules over non-proper scoring rules, since the best-attainable performance defined in the former implies the best-attainable performance in the latter.
+Information Model.
 We cast the information available to a DM as a signal defined within an information model.
@@
 There are $n$ “basic signals” represented as random variables $\Sigma_{1},\ldots,\Sigma_{n}$, from the signal spaces $\mathbf{\Sigma}_{1},\ldots,\mathbf{\Sigma}_{n}$.
-These represent information obtained by a decision-maker, e.g., $\mathbf{\Sigma}_{1}=\{\text{cloudy},\text{not cloudy}\}$, $\mathbf{\Sigma}_{2}\in\{0,\ldots,100\}$ for temprature Celsius, etc.
-The decision-maker observes a signal, which is a combination of the basic signals, represented as a set $V\subseteq 2^{\{\Sigma_{1},\dots,\Sigma_{n}\}}$.
+Basic signals represent information available to a decision-maker, e.g., $\mathbf{\Sigma}_{1}=\{\text{cloudy},\text{not cloudy}\}$, $\mathbf{\Sigma}_{2}\in\{0,\ldots,100\}$ for temperature in Celsius, etc.
+The decision-maker observes a signal, which is a subset of the basic signals, $V\subseteq 2^{\{\Sigma_{1},\dots,\Sigma_{n}\}}$.
+Specifically, we use $V=\{\Sigma_{j_{1}},\ldots,\Sigma_{j_{k}}\}$ for a signal having $k$ basic signals and denote the signal space as $\mathbf{V}=\mathbf{\Sigma}_{j_{1}}\times\ldots\times\mathbf{\Sigma}_{j_{k}}$.
 For example, a signal representing a combination of two basic signals $V=\{\Sigma_{1},\Sigma_{2}\}$ observed by the decision-maker might consist of cloudiness $\Sigma_{1}$ and the temperature $\Sigma_{2}$ of the day.
-Given a signal composed of $m$ basic signals, we write the realization of $V$ as $v=(\sigma_{j_{1}},\dots,\sigma_{j_{m}})$, where the realizations $\sigma_{j_{i}}\in\mathbf{\Sigma}_{j_{i}}$ are sorted by the index of the basic signals $j_{i}\in[n]$.
+Given a signal composed of $m$ basic signals, we write the realization of $V$ as $v=(\sigma_{j_{1}},\dots,\sigma_{j_{k}})$, where the realizations $\sigma_{j_{i}}\in\mathbf{\Sigma}_{j_{i}}$ are sorted by the index of the basic signals $j_{i}\in[n]$.
 The union $V$ of two signals $V_{1},V_{2}$ takes the set union, i.e., $V=V_{1}\cup V_{2}$.
-Though $V$ is initially defined as a set of random variables, we will slightly abuse notation $V$ to represent a random variable that is drawn from the joint distribution of the basic signals in it.
   • Data-generating process.
 A data-generating process is a joint distribution $\pi\in\Delta(\mathbf{\Sigma}_{1}\times\ldots\times\mathbf{\Sigma}_{n}\times\mathbf{\Omega})$ over the basic signals and the payoff-relevant state.
-However, the DM may only observe a subset $V$ of the $n$ basic signals.
+$\pi$ can be viewed as the combination of two distributions: the prior distribution of the state $\Pr[\omega]$ and the signal-generating distribution $\Pr[v|\omega]$ defining the conditional distribution of signals.
+The DM may only observe a subset $V$ of the $n$ basic signals.
 We can define the Bayesian posterior belief upon receiving a signal $V=v$ from the data-generating model as
-    $\Pr[\omega|v]=\frac{\pi(v,\omega)}{\pi(\omega)}$
-    where we slightly abuse notation to write $\pi(v,\omega)$ as the marginal probability of the signal realized to be $v$ and the state being $\omega$ with expectation over unobserved signals.
-Information value
-Our framework quantifies the value of information in a signal $V$ as the extent to which the payoff could be improved by the ideal use of $V$ over a baseline information set.
-We suppose a rational Bayesian DM who knows the data-generating process, observes a signal realization, updates their prior to arrive at posterior beliefs, and then chooses a decision to maximize their expected payoff based on the posterior belief.
-Formally, the rational DM’s expected payoff given a (set of) signal(s) $V$ is
-$\mathrm{R}^{\pi,S}(V)=\text{\bf E}_{v\sim\pi}\![{\max_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\Pr(\omega|v)}\![{S(d,\omega)}]}]$
-We use $\emptyset$ to represent a null signal, such that $\mathrm{R}^{\pi,S}(\emptyset)$ is the expected payoff of a Bayesian rational DM who has no access to a signal but only uses their prior belief to make decisions.
+    $\pi(\omega|v):=\Pr[\omega|v]=\frac{\pi(v,\omega)}{\pi(\omega)}$
+    where we slightly abuse notation to write $\pi(v,\omega)$ as the marginal probability of the signal realized to be $v$ and the state being $\omega$ with expectation over other signals.
+The choice of basic signals directly impacts how many observed samples are required to get a good estimate of the data-generating process.
+When high-dimensional signals such as images or text are used, it may not be computationally feasible to estimate the data-generating process.
+In such cases, one can preprocess the high-dimensional signals to get low-dimensional representations which (ideally) capture any important structure.
+These lower-dimensional signals can be defined by humans, such as when concepts are identified and then used to label the high-dimensional signals (e.g., images or parts of images).
+Alternatively, they can be defined algorithmically, by strategically applying dimensionality reduction to generate low-dimensional embeddings.
+We introduce an algorithm for identifying an “optimal” reduction model in decision problems in Appendix B.
+We demonstrate these two methods in Section 4 and Section 5 respectively.
+Information value.
+Our framework quantifies the value of information in a signal $V$ as the expected payoff improvement of an idealized agent who has access to $V$ in addition to some baseline information set.
+We suppose a rational Bayesian DM who knows the prior probability of the state and conditional distribution of signals (i.e., the data-generating process), observes a signal realization, updates their prior to arrive at posterior beliefs, and then chooses a decision to maximize their expected payoff given their posterior belief.
+Formally, given a decision task with payoff function $S$ and an information model $\pi$, the rational DM’s expected payoff given a (set of) signal(s) $V$ is
+$\mathrm{R}^{\pi,S}(V)=\text{\bf E}_{(v,\omega)\sim\pi}\![{S(d^{r}(v),\omega)}]$ (2)
+where $d^{r}(\cdot):\mathbf{V}\rightarrow\mathbf{D}$ denotes the decision rule adopted by the rational DM.
+$d^{r}(v)=\arg\max_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\pi(\omega|v)}\![{S(d,\omega)}]$ (3)
+We use $\emptyset$ to represent a null signal, such that $\mathrm{R}^{\pi,S}(\emptyset)$ is the expected payoff of a Bayesian rational DM who only uses the prior distribution to make decisions.
 In this case, the Bayesian rational DM will take the best fixed action under the prior, and their expected payoff is:
-$\mathrm{R}^{\pi,S}(\emptyset)=\max_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\pi}\![{S(d,\omega)}]$
-$\mathrm{R}^{\pi,S}(\emptyset)$ defines the maximum expected payoff that can be achieved with no information.
+$\mathrm{R}^{\pi,S}(\emptyset)=\max_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\pi}\![{S(d,\omega)}]$ (4)
+This baseline defines the maximum expected payoff that can be achieved with no information.
 Bayesian decision theory quantifies the information value of $V$ by the payoff improvement of $V$ over the payoff obtained without information.
 Definition 3.1.
-Given a decision task with payoff function $S$ and an information model $\pi$, we define the information value of $V$ as
-$IV^{\pi,S}(V)=\mathrm{R}^{\pi,S}(V)-\mathrm{R}^{\pi,S}(\emptyset)$
-We adopt the same idea to define the agent-complementary information values in our framework.
+Given a decision task with payoff function $S$ and an information model $\pi$, the information value of $V$ is defined as
+$\operatorname{IV}^{\pi,S}(V)=\mathrm{R}^{\pi,S}(V)-\mathrm{R}^{\pi,S}(\emptyset)$ (5)
+We adopt the same idea to define the agent-complementary information value in our framework.
 3.1 Agent-Complementary Information Value
-Given the above definitions, it becomes possible to measure the additional value that new signals can provide over the information already captured by an agent’s decisions.
+With the above definitions, it is possible to measure the additional value that new signals can provide over the information already captured by an agent’s decisions.
 Here, agent may refer to a human, an AI system, or a human–AI team.
-The intuition behind our approach is that any information that is used by decision-makers should eventually reveal itself through variation in their behaviors.
+The intuition behind our approach is that any information that is used by decision-makers should eventually reveal itself through variation in their decisions.
 We recover the information value in agent decisions by offering the decisions as a signal to the Bayesian rational DM.
-We model the agent decisions as a random variable $D^{b}$ from the action space $\mathbf{D}$, which follows a joint distribution $\pi^{b}\in\Delta(\mathbf{\Omega}\times\mathbf{\Sigma}_{1}\times\ldots\times\mathbf{\Sigma}_{n}\times\mathbf{D})$ with the state and signals.
-The expected payoff of a Bayesian rational DM who knows $\pi^{b}$ is given by the function:
-$\mathrm{R}(D^{b})=\text{\bf E}_{d^{b}\sim\pi^{b}}\![{\max_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\Pr(\omega|D^{b}=d^{b})}\![{S(d,\omega)}]}]$
-We seek to identify signals $V$ that can potentially improve agent decisions by analyzing the information value in the combined signal $D^{b}\cup V$ and the information value in $D^{b}$, which we define as agent-complementary information value.
+We model the agent decisions as a random variable $D^{b}$ from the action space $\mathbf{D}$, which follows a joint distribution $\pi\in\Delta(\mathbf{\Sigma}_{1}\times\ldots\times\mathbf{\Sigma}_{n}\times\mathbf{\Omega}\times\mathbf{D})$ with the state and signals.
+The expected payoff of a Bayesian rational DM who knows $\pi$ is given by the function:
+$\mathrm{R}^{\pi,S}(D^{b})=\text{\bf E}_{(d^{b},\omega)\sim\pi}\![{S(d^{r}(d^{b}),\omega)}]$
+We seek to identify signals $V$ that can potentially improve agent decisions by analyzing the information value in the combined signal $D^{b}\cup V$ over the information value in $D^{b}$, which we define as the agent-complementary information value.
 Definition 3.2.
-Given a decision task with payoff function $S$ and an information model $\pi$, we define the agent-complementary information value of $V$ on agent decisions $D^{b}$ as
-$ACIV^{\pi,S}(V;D^{b})=\mathrm{R}^{\pi,S}(D^{b}\cup V)-\mathrm{R}^{\pi,S}(D^{b})$
-If the $ACIV$ of a signal $V$ is low, this means either that the $IV$ of $V$ is low (e.g., it is not correlated with $\omega$), or that the agent has already exploited the information in $V$ (e.g., the agent relies on $V$ to make their decisions such that their decisions correlate with $\omega$ in the same way as $V$ correlates with $\omega$).
-If, however, the $ACIV$ of $V$ is high, then at least in theory, the agent can improve their payoff by incorporating $V$ in their decision making.
-Furthermore, $ACIV$ can reveal complementary information between different types of agents.
-For instance, if we view AI predictions as $V$ and treat human decisions as the existing agent signal $D^{b}$, a large $ACIV$ indicates that AI predictions add considerable value beyond what humans alone achieve.
-In the reverse scenario, if human decisions serve as $V$ and AI predictions are $D^{b}$, we can measure how much humans contribute on top of the AI.
-We demonstrate further usage of $ACIV$ in Section 4 and Section 5.
-3.2 Instance-level Agent-Complementary
-Instance-level Agent-Complementary Information Value ($ILIV$) evaluates the additional information contributed by a single realization of a signal rather than the entire joint distribution.
-This finer-grained view is critical for tasks where we need to understand the information value on individual instances, such as when asking whether a human expert should trust an individual prediction from an AI or how to help a human expert understand how the AI model is exploiting information about an instance.
-To quantify the information value of a realization $V=v$, we construct $Z_{v}$ as a binary variable indicating whether signal $V$ is realized as $v$, i.e., $Z_{v}=\mathbb{1}[V=v]$.
-The data-generating model defining the joint distribution of $Z_{v}$, $\pi_{v}\in\Delta(\mathbf{\Omega}\times\{0,1\})$, can be constructed through transforming the original data-generating model $\pi$.
-Because $Z_{v}$ is a garbling of $V$ (Blackwell et al., 1951), there always exists a a Markov matrix $\Gamma:\mathbf{V}\rightarrow\{0,1\}$ such that the new data-generating process $\pi_{v}$ can be constructed through $\pi_{v}=\Gamma\pi$.
+Given a decision task with payoff function $S$ and an information model $\pi$, we define the agent-complementary information value ($\operatorname{ACIV}$) of $V$ on agent decisions $D^{b}$ as
+$\operatorname{ACIV}^{\pi,S}(V;D^{b})=\mathrm{R}^{\pi,S}(D^{b}\cup V)-\mathrm{R}^{\pi,S}(D^{b})$ (6)
+If the $\operatorname{ACIV}$ of a signal $V$ is small relative to the baseline (4), this means either that the information value of $V$ to the decision problem is low (e.g., it is not correlated with $\omega$), or that the agent has already exploited the information in $V$ (e.g., the agent relies on $V$ or equivalent information to make their decisions such that their decisions correlate with $\omega$ in the same way as $V$ correlates with $\omega$).
+If, however, the $\operatorname{ACIV}$ of $V$ is large relative to the baseline, then at least in theory, the agent can improve their payoff by incorporating $V$ in their decision making.
+Furthermore, $\operatorname{ACIV}$ can reveal complementary information between different types of agents.
+For instance, if we view AI predictions as $V$ and treat human decisions as the agent signal $D^{b}$, a large $\operatorname{ACIV}$ indicates that AI predictions add considerable value beyond what humans alone achieve.
+In the reverse scenario, if human decisions serve as $V$ and AI predictions are $D^{b}$, we can measure how much humans can contribute over the information captured in the AI predictions.
+We demonstrate uses of $\operatorname{ACIV}$ in Section 4 and Section 5.
+3.2 Instance-level Agent-complementary Information Value
+$\operatorname{ACIV}$ quantifies the value of the decision-relevant information in a signal $V$ across all possible realizations defined by the data-generating model.
+To provide analogous instance-level quantification of information value, we define Instance-Level agent-complementary Information Value ($\operatorname{ILIV}$) to quantify the value of the decision-relevant information encoded by a single realization of the signal.
+This finer-grained view makes it possible to analyze how much an agent can benefit in theory from better incorporating instance-level information in their decision.
+Given a realization of signal $v=\{\sigma_{j_{1}},\ldots,\sigma_{j_{k}}\}$, we want to know the maximum expected payoff gain from gaining access to $v$ on the instances where $v$ is realized over the existing information encoded in agent decisions.
+Intuitively, this captures how much “room” there is for specific decisions to be improved through better use of the signal.
+To calculate instance-level information value, we rely on the performance of the Bayesian rational agent on the conditional distribution $\pi(\sigma_{1},\ldots,\sigma_{n},\omega,d^{b}|v)$ instead of the whole distribution $\pi(\sigma_{1},\ldots,\sigma_{n},\omega,d^{b})$.
+Formally, given a decision task with payoff function $S$ and information model $\pi$, the expected payoff of the rational DM given signal $V=v$ on instances where $V=v$ is
+$\mathrm{r}^{v,\pi,S}(v)=\text{\bf E}_{\omega\sim\pi(\omega|v)}\![{S(d^{r}(v),\omega)}]$ (7)
+where $d^{r}(v)$ is the Bayesian optimal decision on receiving $v$ as defined in Equation 3.
+If we consider the agent decisions in addition to the realization $v$, the rational DM’s expected payoff on instances where $V=v$ can be written as
+$\mathrm{r}^{v,\pi,S}(v;D^{b})=\text{\bf E}_{(d^{b},\omega)\sim\pi(d^{b},\omega|v)}\![{S(d^{r}(v\cup d^{b}),\omega)}]$ (8)
 Definition 3.3.
-Given a decision task with payoff function $S$ and an information model $\pi$, we define the instance-level agent-complementary information value of the realization $V=v$ as:
-$ILIV^{\pi,S}(v;D^{b})=\mathrm{R}^{\pi_{v},S}(D^{b}\cup Z_{v})-\mathrm{R}^{\pi_{v},S}(D^{b})$
-This local measure captures how much additional payoff can be gained by incorporating the specific realization $v$ into the agent decisions $D^{b}$.
-Summing $ILIV$ over all possible realizations of $V$ recovers the global agent-complementary information value ($ACIV$).
-Proposition 3.4.
-$ACIV^{\pi,S}(V;D^{b})=\sum_{v\in\mathbf{V}}ILIV^{\pi,S}(v;D^{b})$
-We apply $ILIV$ to define an information-based explanation technique (ILIV-SHAP) that extends SHAP to explains how the information value of AI predictions complements human decisions $D^{\text{Human}}$ for specific instances.
-Vanilla SHAP (Lundberg and Lee, 2017) defines a saliency-based explanation with a set of effect variables $\phi_{i}$ representing the influence of the realization of basic signal $\Sigma_{i}=\sigma_{i}$ on the model output $f(v)$, where $v=(\sigma_{1},\ldots,\sigma_{n})$ and $f:\mathbf{V}\rightarrow\mathbb{R}$.
-Lundberg and Lee (2017) show that the only possible explanation model fulfills properties of local accuracy, missingness and consistency for model $f$ and instance $v$ is following:
-Definition 3.5 (SHAP (Lundberg and Lee, 2017)).
-$\phi_{i}(f,v)=\sum_{v^{\prime}\subseteq v}\frac{|v^{\prime}|!(n-|v^{\prime}|-1)!}{n!}[g_{f}(v^{\prime})-g_{f}(v^{\prime}\backslash\sigma_{i})]$
-where $g_{f}(v^{\prime})$ denotes the expectation of model output conditioned on the subset of signals with $V^{\prime}=v^{\prime}$, i.e., $\text{\bf E}\![{f(V)|V^{\prime}=v^{\prime}}]$.
-Instead of quantifying changes in the AI’s raw predictions, ILIV-SHAP clarifies how each feature contributes to the agent-complementary information value of predictions.
-Proposition 3.6 (ILIV-SHAP).
-An explanation model whose effective variable defined as
-$\phi_{i}(ILIV(f(\cdot);D^{b}),v)$
-fulfill properties of local accuracy, missingness and consistency for $ILIV$ of $f(\cdot)$ on instance $v$.
-Intuitively, when $\phi_{i}(ILIV(f(\cdot);D^{b}),v)$ for a feature $\sigma_{i}$ is relatively large, it means that the model $f(\cdot)$ is extracting information that $D^{b}$ lacks from $\Sigma_{i}$ (or other features with equivalent information to it).
-Otherwise, when $\sigma_{i}$ makes a small relative contribution, it means that it makes the model ignore information in other features that have higher information value for human decision-makers (such that on average lossing information value).
+Given a decision task with payoff function $S$ and an information model $\pi$, we define the instance-level agent-complementary information value ($\operatorname{ILIV}$) of signal $V=v$ on instances where $V=v$ as:
+$\operatorname{ILIV}^{v,\pi,S}(v;D^{b})=\mathrm{r}^{v,\pi,S}(v;D^{b})-\mathrm{r}^{v,\pi,S}(\emptyset;D^{b})$ (9)
+where $\mathrm{r}^{v,\pi,S}(\emptyset;D^{b})$ represents the rational DM’s expected payoff on instances where $V=v$ with only the exisiting information encoded in agent decisions.
+Taking the expectation of $\operatorname{ILIV}$ over $V$ recovers the global agent-complementary information value ($\operatorname{ACIV}$), i.e.,
+$\operatorname{ACIV}^{\pi,S}(V;D^{b})=\text{\bf E}_{v\sim\pi(v)}\![{\operatorname{ILIV}^{v,\pi,S}(v;D^{b})}]$
+3.2.1 ILIV-SHAP Information-based Explanation
+We apply $\operatorname{ILIV}$ to define an information-based explanation technique (ILIV-SHAP) that extends the canonical SHAP feature saliency explanation of a model’s prediction.
+ILIV-SHAP communicates how the data features lead to changes in the information value of AI predictions.
+While traditional SHAP summarizes the average contribution of each feature to a specific prediction over the baseline (prior) prediction, ILIV-SHAP summarizes the average contribution of each feature to the decision-relevant information value contained in the prediction.
+Suppose a model $f$ that takes as input $m$ features and outputs a real number.
+Given an instance $\mathbf{x}=(x_{1},\ldots,x_{m})$, the importance of one feature $x_{i}$ to the model output $f(\mathbf{x})$ is encoded by the expected difference of model outputs when $x_{i}$ is marginalized out.
+Specifically, this is quantified by $f(\mathbf{x})-\text{\bf E}\![{f(X)|X_{-i}=\mathbf{x}_{-i}}]$, where $X_{-i}$ denotes all features except $X_{i}$.
+Considering the interaction between features, SHAP (Lundberg and Lee, 2017) uses the Shapley value to quantify the importance scores averaged on different combinations of features:
+$\phi_{i}(f,\mathbf{x})=\sum_{\mathbf{x}^{\prime}\subseteq\mathbf{x}}\frac{|\mathbf{x}^{\prime}|!(m-|\mathbf{x}^{\prime}|-1)!}{m!}[g_{f}(\mathbf{x}^{\prime})-g_{f}(\mathbf{x}^{\prime}\backslash x_{i})]$
+where $g_{f}(\mathbf{x}^{\prime})$ denotes the expected model output conditioned on $\mathbf{x}^{\prime}$, i.e., $\text{\bf E}\![{f(X)|X^{\prime}=\mathbf{x}^{\prime}}]$ for any $\mathbf{x}^{\prime}\subseteq\mathbf{x}$.
+The scores $\phi_{i}(f,\mathbf{x})$ output by SHAP construct an explanation model for a model output, which quantifies the expected counterfactual change in the model output caused by the feature $x_{i}$.
+ILIV-SHAP extends SHAP to give an explanation model on how features impact the decision-relevant information value of an individual model output.
+Definition 3.4 (ILIV-SHAP).
+Given a model $f$ and data features $\mathbf{x}=(x_{1},\ldots,x_{m})$, the importance score of the i-th feature by ILIV-SHAP is
+$\phi_{i}^{\operatorname{ILIV}}(f,\mathbf{x})=\sum_{\mathbf{x}^{\prime}\subseteq\mathbf{x}}\frac{|\mathbf{x}^{\prime}|!(m-|\mathbf{x}^{\prime}|-1)!}{m!}[\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(g_{f}(\mathbf{x}^{\prime});D^{b})-\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(g_{f}(\mathbf{x}^{\prime}\backslash x_{i});D^{b})]$
+where $\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(g_{f}(\mathbf{x}^{\prime});D^{b})$ denotes a counterfactual evaluation of $\operatorname{ILIV}$, which quantifies the expected payoff gain from additionally knowing $g_{f}(\mathbf{x}^{\prime})$ on the instances where $f(\mathbf{x})$ is realized.
+This counterfactual version of $\operatorname{ILIV}$ is guaranteed to achieve the maximum at $\mathbf{x}^{\prime}$ where $f(\mathbf{x})=g_{f}(\mathbf{x}^{\prime})$, i.e., the features missing from $\mathbf{x}^{\prime}$ have no impact on $f(\mathbf{x})$.
+The explanation model offered by ILIV-SHAP is grounded in the following two properties.
+First, $\phi^{\operatorname{ILIV}}_{i}(f,\mathbf{x})$ are consistent with the extent to which feature $x_{i}$ contributes to the information value in $f(\mathbf{x})$.
+Specifically, for any two models $f$ and $f^{\prime}$ and any $\mathbf{x}^{\prime}\subseteq\mathbf{x}$, if $x_{i}$ contributes more to the information value in $f^{\prime}$ than the information value in $f$, i.e., $\operatorname{ILIV}^{f^{\prime}(\mathbf{x}),\pi,S}(g_{f^{\prime}}(\mathbf{x}^{\prime});D^{b})-\operatorname{ILIV}^{f^{\prime}(\mathbf{x}),\pi,S}(g_{f^{\prime}}(\mathbf{x}^{\prime}\backslash x_{i});D^{b})\geq\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(g_{f}(\mathbf{x}^{\prime});D^{b})-\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(g_{f}(\mathbf{x}^{\prime}\backslash x_{i});D^{b})$, then $\phi^{\operatorname{ILIV}}_{i}(f^{\prime},\mathbf{x})\geq\phi^{\operatorname{ILIV}}_{i}(f,\mathbf{x})$.
+This property of ILIV-SHAP follows the consistency property of SHAP (Lundberg and Lee, 2017).
+Second, summing up the scores $\phi^{\operatorname{ILIV}}_{i}(f,\mathbf{x})$ recovers the information value of model output, $\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(f(\mathbf{x});D^{b})$.
+Formally, for any $f$ and any $\mathbf{x}$, $\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(f(\mathbf{x});D^{b})=\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(\text{\bf E}\![{f(X)}];D^{b})+\sum_{i=1}^{m}\phi^{\operatorname{ILIV}}_{i}(f,\mathbf{x})$.
+We demonstrate use of ILIV-SHAP in Section 6.
 3.3 Robustness Analysis of Information Order
-Ambiguity about the appropriate payoff function is not uncommon in human-AI decision settings due to challenges of eliciting utility functions and potential variance in these functions across decision-makers or groups of instances; e.g., doctors penalize certain false negative results differently when diagnosing younger versus older patients (Mclaughlin and Spiess, 2023).
-We therefore define the partial order of complementary information value using Blackwell order.
-Blackwell’s comparison of signals (Blackwell et al., 1951) defines a (set of) signal $V_{1}$ as more informative than $V_{2}$ if $V_{1}$ has a higher information value on all possible decision problems.
-We identify this partial order by decomposing the space of decision problems via a basis of proper scoring rules  (Li et al., 2022; Kleinberg et al., 2023).
+Our approach assumes the specification of a decision problem on which agents’ decisions and use of information are evaluated.
+However, ambiguity around the appropriate decision problem, and in particular, the appropriate scoring rule, is not uncommon in human-AI decision settings.
+Ambiguity can arise as a result of challenges in eliciting utility functions and/or variance in these functions across decision-makers or groups of instances; for example, doctors penalize certain false negative results differently when diagnosing younger versus older patients (Mclaughlin and Spiess, 2023).
+Blackwell’s comparison of signals (Blackwell et al., 1951) is an appropriate tool for addressing ambiguity about the payoff function, as it defines a (set of) signal $V_{1}$ as more informative than $V_{2}$ if $V_{1}$ has a higher information value on all possible decision problems.
+We identify this partial order by decomposing the space of decision problems via a basis of proper scoring rules [^2] For rational DMs, any decision problem can be represented by an equivalent proper scoring rule in Equation 1, such that the partial order defined via proper scoring rules also applies to the corresponding decisin tasks.
+(Li et al., 2022; Kleinberg et al., 2023).
 Definition 3.5 (Blackwell Order of Information).
-A (set of) signal $V_{1}$ is Blackwell more informative than $V_{2}$ if $V_{1}$ achieves a higher payoff on any decision problem,
+A signal $V_{1}$ is Blackwell more informative than $V_{2}$ if $V_{1}$ achieves a higher best-attainable payoff on any decision problems:
 $\mathrm{R}^{\pi,S}(V_{1})\geq\mathrm{R}^{\pi,S}(V_{2}),\forall S$
-We test the Blackwell order between signals on a basis of proper scoring rules induced from decision problems.
-The basis is the set of V-shaped scoring rules, parameterized by the kink of the piecewise-linear utility function.
+The Blackwell order is evaluated over all possible decision problems, which cannot be tested directly.
+Fortunately, we only need to test over all proper scoring rules since any decision problem can be represented by an equivalent proper scoring rule, and the space of proper scoring rules can be characterized by a set of V-shaped scoring rules.
+A V-shaped scoring rule is parameterized by the kink of the piecewise-linear utility function.
 Definition 3.6.
@@
 Proposition 3.7 (Hu and Wu 2024).
-If
-$\mathrm{R}^{\pi,S_{\mu}}(V_{1})\geq\mathrm{R}^{\pi,S_{\mu}}(V_{2}),\forall\mu\in(0,1),$
+If $\forall\mu\in(0,1)$
+$\mathrm{R}^{\pi,S_{\mu}}(V_{1})\geq\mathrm{R}^{\pi,S_{\mu}}(V_{2}),$
 then $V_{1}$ is Blackwell more informative than $V_{2}$.
 Extending this to agent-complementary information value, we say that $V_{1}$ offers a higher complementary value than $V_{2}$ under the Blackwell order if
-$ACIV^{\pi,S_{\mu}}(V_{1};D^{b})\geq ACIV^{\pi,S_{\mu}}(V_{2};D^{b}),\forall\mu\in(0,1)$
-This definition allows us to rank signals (or sets of signals) universally, without needing to pin down any specific payoff function prior to the analysis. See the use case in Section 4.
-4 Experiment I: Model Comparison on Chest Radiographs Diagnosis
+$\operatorname{ACIV}^{\pi,S_{\mu}}(V_{1};D^{b})\geq\operatorname{ACIV}^{\pi,S_{\mu}}(V_{2};D^{b}),\forall\mu\in(0,1)$
+This definition allows us to rank signals (or sets of signals) without needing to commit to a specific payoff function. We present a use case in Section 4.
+4 Experiment I: Model Comparison on Chest Radiograph Diagnosis
 We apply our framework to a well-known cardiac dysfunction diagnosis task (Rajpurkar et al., 2018; Tang et al., 2020; Shreekumar, 2025).
-We apply the framework to analyze how much complementary information value a set of possible AI models offer to human decision-makers.
-Figure 1: Information value of all deep-learning models calculated under our framework using Brier score as payoff function.
+We demonstrate how our framework can be used in model evaluation for analyzing how much complementary information value a set of possible AI models offers to the radiology reports written by experts.
+Figure 1: Information value of all deep-learning models calculated under our framework.
+The first row represents the human-alone decisions (without considering any AI predictions as additional signals).
+The other rows are the combinations of the human-alone decisions and the AI predictions from different pre-trained models.
+We list the AI predictions alone to show the AI-complementary information value offered by human decisions.
 4.1 Data and Model
@@
 (2022).
-We use the labels from Irvin et al.
-(2019) as the human decisions (without AI’s assistance) in the diagnosis task, which is a rule-based tool labeling the symptoms as positive, negative, or uncertain, i.e., $d\in\{+,?,-\}$ [^2] The three symbols represents the encoding we use for signal construction, not an assertion of how radiologists communicate..
-We fune-tuned five deep-learning models on the cardiac dysfunction diagnosis task, VisionTransformer (Alexey, 2020), SwinTransformer (Liu et al., 2021), ResNet (He et al., 2016), Inception-v3 (Szegedy et al., 2016), and DenseNet (Huang et al., 2017).
-Our training set contains 12,228 images and validation set contains 6,115 images.
-On a hold-out test set with 12,229 images, the AUC achieved by the five models are: DenseNet with $0.77$, Inception v3 with $0.76$, ResNet with $0.77$, SwinTransformer with $0.78$, VisionTransformer with $0.80$.
-We consider Brier score as a payoff function, and also conduct a robust analysis considering various V-shaped payoff functions with different kinks on a discretized grid of $[0,1]$ with a step $0.01$.
-We use the hold-out test set to estimate the data-generating process, which defines the joint distribution of state, human decisions and AI models’ predictions.
+We label the radiology reports by a rule-based tool (Irvin et al., 2019) and use the labels as the human decisions (without AI assistance) in the diagnosis task to solve the problem of computational feasibility with high-dimensional textual reports.
+The labels are represented by the symptoms as positive, negative, or uncertain, i.e., $d\in\{+,?,-\}$.
+We fine-tuned five deep-learning models on the cardiac dysfunction diagnosis task, VisionTransformer (Alexey, 2020), SwinTransformer (Liu et al., 2021), ResNet (He et al., 2016), Inception-v3 (Szegedy et al., 2016), and DenseNet (Huang et al., 2017).
+Our training set contains 12,228 images, and the validation set contains 6,115 images.
+On a hold-out test set with 12,229 images, the AUC achieved by the five models is: DenseNet with $0.77$, Inception v3 with $0.76$, ResNet with $0.77$, SwinTransformer with $0.78$, and VisionTransformer with $0.80$.
+We consider Brier score, a.k.a., quadratic score, as the payoff function: $S(\omega,d)=1-(\omega-d)^{2}$.
+The scale of the quadratic score is $[0,1]$ and a random guess ($d\sim\text{Bernoulli}(0.5)$) achieves $0.75$ payoff.
+We use the quadratic score instead of the mean absolute error that is usually used in cardiac dysfunction diagnosis task because the quadratic score is a proper scoring rule where truthfully reporting the belief maximizes the payoff [^3] We prefer a proper scoring rule so that the rational decision-maker’s strategy is to reveal their true belief, ensuring that the signal’s information value accurately reflects its role in forming beliefs..
+We also conduct a robust analysis considering various V-shaped payoff functions with different kinks on a discretized grid of $[0,1]$ with a step of $0.01$.
+We use the hold-out test set to estimate the data-generating process, which defines the joint distribution of state, human decisions, and AI models’ predictions.
+We construct the scale of performances by a no-information bound and a full-information bound.
+The no-information bound is $\mathrm{R}^{\pi,S}(\emptyset)$, the baseline as we define the information value.
+The full-information bound is defined as the expected payoff of a rational DM who has access to all signals, human label from radiology report and predictions from five AI models.
+We do not consider the radiology images as a basic signal in this task because of the problem of computational feasibility with high-dimensional signals.
+We use the five vision models’ predictions as embeddings of the original images.
+See more detail on the algorithm of choosing the optimal embeddings in Appendix B.
 4.2 Results
@@
 We first analyze the agent-complementary information values in Figure 1, using the Brier score as the payoff function.
-We find that all AI models provide complementary information value to human judgment.
-This highlights the same takeaways as section 5, that the AI model has considerable potential to improve human decisions.
-As shown in Figure 1 (comparison between $\mathrm{R}^{\pi,S}(D^{\text{Human}}\cup D^{\text{AI}})$ and $\mathrm{R}^{\pi,S}(D^{\text{Human}})$), all AI models capture at least $20\%$ of the total available information value that is not exploited by human decisions.
-This motivates deploying at AI to assist humans in this scenario.
-From the other direction, the human decisions also provide complementary information to all AI models, comparing $\mathrm{R}^{\pi,S}(D^{\text{AI}}\cup D^{\text{AI}})$ with $\mathrm{R}^{\pi,S}(D^{\text{AI}})$ in Figure 1.
-In scenarios where partial automation is possible, this observation might inspire, for example, further investigation into what information the humans may have access to that is not represented in AI training data.
+We find that all AI models provide complementary information value to the aforementioned human judgment.
+As shown in Figure 1 (comparison between $\mathrm{R}^{\pi,S}(D^{\text{Human}}\cup D^{\text{AI}})$ and $\mathrm{R}^{\pi,S}(D^{\text{Human}})$), all AI models capture at least $20\%$ of the total available information value (across all AI model and human decisions) that is not exploited by human decisions.
+This motivates deploying an AI to assist humans in this scenario.
+In the other direction, the human decisions also provide complementary information to all AI models, comparing $\mathrm{R}^{\pi,S}(D^{\text{Human}})$ with $\mathrm{R}^{\pi,S}(D^{\text{AI}})$ in Figure 1.
+This observation might inspire, for example, further investigation of the information the humans can access to that is not represented in AI training data.
 Which AI model offers the most decision-relevant information over human judgments?
-Figure 1 shows that VisionTransformer contains slightly higher information value than the other models and Inception v3 contains slightly lower information value than the other models.
-However, these differences are slight.
-If the payoff function were questionable in any way, an organization may not want to trust the model rankings.
-This motivates evaluating model performance over many possible losses to test if there is a Blackwell ordering of models.
+Figure 1 shows that VisionTransformer contains slightly higher information value than the other models, and Inception v3 contains slightly lower information value than the other models.
+We assess the stability of VisionTransformer’s superiority over the other AI models across many possible losses to test if there is a Blackwell ordering of models.
+By Proposition 3.7, we test the payoff of models on all V-shaped scoring rules, shown in Figure 4.
 Across all the V-shaped payoff functions, we find that VisionTransformer is Blackwell more informative and Inception v3 is Blackwell less informative than all other models.
-By Proposition 3.9, we test the payoff of models on all V-shaped scoring rules, shown in Figure 4.
 The VisionTransformer achieves a higher information value on all V-shaped scoring rules, implying a higher information value on all decision problems.
-This analysis highlights the insufficiency of accuracy-based model comparisons to account for all downstream decision problems: 1) while accuracy may rank models in one way, there may exist decision problems where the order between models is reversed; 2) while two models may seem comparable in accuracy, one can be more informative than the other for all decision problems, which is robustly good.
 5 Experiment II: Behavioral Analysis on Deepfake Detection
-We analyze a deepfake video detection task (Dolhansky et al., 2020), where participants are asked to judge whether a video was created by generative AI.
-We apply the framework to benchmark the use of available information by the human, AI, and human-AI team.
-Figure 2: Information value calculated under our framework in the information model defined by experiment data from Groh et al.
+We apply our framework to analyze a deepfake video detection task (Dolhansky et al., 2020), where participants are asked to judge whether a video was created by generative AI, including with the assistance of an AI model.
+Figure 2: Information value calculated under our framework in the information model defined by the experiment of Groh et al.
 (2022).
+Basic signals include the seven video level features and three types of agent decisions.
 The baseline on the left represents the expected payoff given no information, i.e., $\mathrm{R}^{\pi,S}(\emptyset)$, and the benchmark on the right represents the expected payoff given all available information, i.e., $\mathrm{R}^{\pi,S}(\{\Sigma_{1},\ldots,\Sigma_{n},D^{\text{Human}},D^{\text{Human-AI}},D^{\text{AI}}\})$.
@@
 (2022).
-They recruited 5,524 non-expert participants through Prolific.
-Participants were asked to examine the videos, and provided with assistance from a computer vision model, which achieved $65\%$ accuracy on a holdout dataset.
+Non-expert participants (n=5,524) were recruited through Prolific and asked to examine the videos.
 They reported their decisions in two rounds.
-They first reviewed the video and reported an initial decision ($D^{\text{Human}}$) without access to the model.
-Then, in a second round, they were told the AI’s recommendation ($D^{\text{AI}}$) and chose whether to change their initial decision, resulting in a final decision ($D^{\text{Human-AI}}$).
-Participants’ decisions (both initial and final) were elicited as a percentage indicating how confident they were that the video was a deepfake, measured in $1\%$ increments: $d\in\{0\%,1\%,\ldots,100\%\}$.
-We round the predictions from the AI to the same 100-scale probability scale available to study participants.
-We use the Brier score as the payoff function: $S(\omega,d)=1-(\omega-d)^{2}$, with the binary payoff-related state: $\omega\in\{0,1\}=\{\text{genuine},\text{fake}\}$.
-The scale of Brier score is $[0,1]$ and a random guess ($d\sim\text{Bernoulli}(0.5)$) achieves $0.75$ payoff under Brier score.
-We choose the Brier score instead of the norm-1 accuaracy used by Groh et al.
-(2022) because Brier score is a proper scoring rule where truthfully reporting the belief maximizes the payoff [^3] We prefer a proper scoring rule so that the rational decision-maker’s strategy is to reveal their true belief, ensuring that the signal’s information value accurately reflects its role in forming beliefs.
-If the goal is merely to evaluate the outcome performance of real decision-makers, a non-proper scoring rule may still suffice..
+They first reviewed the video and reported an initial decision ($D^{\text{Human}}$) without access to the AI model.
+Then, in a second round, they were provided with the recommendation ($D^{\text{AI}}$) of a multitask cascaded convolutional neural network (Zhang et al., 2016), with estimated $65\%$ accuracy on holdout data, and chose whether to change their initial decision.
+This produced a final decision ($D^{\text{Human-AI}}$).
+Both decisions were elicited as a percentage indicating how confident the participant was that the video was a deepfake, measured in $1\%$ increments: $d\in\{0\%,1\%,\ldots,100\%\}$.
+We round the predictions from the AI model to the same 100-scale probability scale available to study participants.
+We use the Brier score as the payoff function, with the binary payoff-related state: $\omega\in\{0,1\}=\{\text{genuine},\text{fake}\}$.
+This choice differs from the mean absolute error used by Groh et al.
+(2022), but again we use the quadratic score because it is a proper scoring rule where truthfully reporting the belief maximizes the score.
 We identify a set of features that were implicitly available to all three agents (human, AI, and human-AI).
-Because the video signal is high dimensional, we make use of seven seven video-level features using manually coded labels by Groh et al.
+Because the video signal is high dimensional, we make use of seven video-level features using manually coded labels by Groh et al.
 (2022): graininess, blurriness, darkness, presence of a flickering face, presence of two people, presence of a floating distraction, and the presence of an individual with dark skin, all of which are labeled as binary indictors.
 These are the basic signals in our framework.
-We estimate the data-generating process $\pi$ using the realizations of signals, state, first round human decisions, AI predictions, and second round human-AI decisions.
+We estimate the data-generating process $\pi$ using the realizations of signals, state, first-round human decisions, AI predictions, and second-round human-AI decisions.
+The no-information bound is the same as Section 4 while the full-information bound is defined as the expected payoff of a rational DM who has access to all video-level features and three agents’ decisions.
 5.2 Results
-How much decision-relevant information do each agent’s decisions offer?
-We first compare the information value of the AI predictions to the decision problem to that of the human decisions in the first round (without AI assistance).
-To contextualize the value of each, we first construct a scale ranging from the expected payoff of rational DM with no information, i.e., $\mathrm{R}^{\pi,S}(\emptyset)$, to that of the rational agent who has access to all information, i.e., $\mathrm{R}^{\pi,S}(\Sigma_{1}\cup\ldots\cup\Sigma_{n}\cup D^{\text{Human}}\cup D^{\text{Human-AI}}\cup D^{\text{AI}})$.
-The lower-bound represented by the rational DM with no information is $0.75$ in Brier score, which is equivalent to the payoff achieved by a random guess drawn from $\text{Bernoulli}(0.5)$.
-Using this scale, Figure 2(a) shows that AI predictions provide about $65\%$ of the total possible information value over the no information baseline, while human decisions only provide about $15\%$.
-Hence, human decisions are only weakly informative for the problem.
+How much decision-relevant information do agents’ decisions offer?
+We first compare the information value of the AI predictions to that of the human decisions in the first round (without AI assistance).
+Figure 2(a) shows that AI predictions provide about $65\%$ of the total possible information value over the no-information baseline, while human decisions only provide about $15\%$.
+Because the no information baseline, $0.75$, is equivalent to a random guess drawn from $\text{Bernoulli}(0.5)$, human decisions are only weakly informative for the problem.
 We next consider the human-AI decisions.
-Given that the AI predictions contained a significant portion of the total possible information value, we might hope that when participants are provided with AI predictions, their performance comes close to the full information baseline.
-However, the information value contained in human-AI decisions is also only take a small proportion of the total possible information value ($30\%$).
-This aligns with the findings by Guo et al.
-(2024) that humans tend to rely on the AI but are bad at distinguishing when AI predictions are correct.
-How much additional decision-relevant information do the available features offer over each agent’s decisions?
-To understand what information might improve human decisions, we assess the $ACIV$s of different signals over different agents.
-This describes the additional information value in the signal after conditioning on the exisiting information in the agents’ decisions.
-As shown on the fifth row in Figure 2, the presence of a flickering face offers larger $ACIV$ over human decisions than over AI predictions, meaning that human decisions could improve by a greater amount if they were to incorporate this information.
-Meanwhile, as shown on the fourth row in Figure 2, the presence of a individual with dark skin offers larger $ACIV$ over AI predictions than over human decisions.
+Given that the AI predictions contain a significant portion of the total possible information value, we might hope that when participants have access to the AI predictions, their performance will be close to the full information baseline.
+However, the information value of the human-AI decisions only achieves a small proportion of the total possible information value ($30\%$).
+This is consistent with the findings of Guo et al.
+(2024) that humans are bad at distinguishing when AI predictions are correct.
+How much additional decision-relevant information do the available features offer over agents’ decisions?
+To understand what information might improve human decisions, we assess the $\operatorname{ACIV}$s of different signals over different agents.
+This describes the additional information value in the signal after conditioning on the existing information in the agents’ decisions.
+As shown on the fifth row in Figure 2, the presence of a flickering face offers larger $\operatorname{ACIV}$ over human decisions than over AI predictions, meaning that human decisions could improve by a greater amount if they were to incorporate this information.
+Meanwhile, as shown on the fourth row in Figure 2, the presence of an individual with dark skin offers larger $\operatorname{ACIV}$ over AI predictions than over human decisions, suggesting that humans make greater use of this information.
 This suggests that the AI and human rely on differing information to make their initial predictions, where the AI relies more on information associated with the presence of a flickering face while human participants rely more on information associated with the presence of an individual with dark skin.
-By comparing the $ACIV$s of different signals over human decisions and human-AI decisions, we also find that simply displaying AI predictions to human did not help human-AI to exploit the observed signals in their decisions.
-As shown in Figure 2, with the assistance of AI, the human-AI teams’ decisions have similar $ACIV$ compared to human decisions (note that the $ACIV$ over human decisions shows significant difference on same signal from the $ACIV$ over AI predictions), except a little improvement on the presence of a flickering face.
-This finding further comfirms the hypothsis on human simply relying on AI predictions without processing the information contained in them.
+By comparing the $\operatorname{ACIV}$s of different signals over human decisions and human-AI decisions, we also find that simply displaying AI predictions to humans did not lead to the AI-assisted humans exploiting the observed signals in their decisions.
+As shown in Figure 2, with the assistance of AI, the $\operatorname{ACIV}$s of all signals over the human-AI teams’ decisions do not change much compared to the $\operatorname{ACIV}$s over human decisions, with the exception of a slight improvement in the presence of a flickering face.
+This finding further confirms the hypothesis that humans are simply relying on AI predictions without processing the information contained in them.
 6 Experiment III: Information-based Local Explanation on Recidivism Prediction
-We apply our framework to a recidivism prediction task, where the decision-maker decides whether to release a defendent. We apply the framework to evaluate and augment saliency-based explanations.
-Figure 3: (A) Summary of SHAP explanation and (B) summary of ILIV-SHAP explanation.
-(C) The waterfall plot of an example of SHAP on instance 4 and (D) the waterfall plot of an example of ILIV-SHAP on instance 4.
-See more examples of ILIV-SHAP in Appendix C.
+We apply our framework to a recidivism prediction task, where the decision-maker decides whether to release a defendant. We demonstrate using the framework to augment SHAP.
+Figure 3: (A) Summary of SHAP explanation and (B) summary of ILIV-SHAP explanation, where each point represents an instance from the test set.
+(C-F) Waterfall plots of applying SHAP and ILIV-SHAP on instance 4 and instance 18.
+The Waterfall plot starts from the prior expected prediction (or ILIV of the prior expected prediction), and then adds features one at a time until reaching the current model output (or ILIV of the current model output).
+See further comparisons between SHAP and ILIV-SHAP in Appendix D.
 6.1 Data and Model
-We use the dataset from COMPAS dataset (Angwin et al., 2022), which contains 11,758 defendents with associated features capturing demographics, information about their current charges, and their prior criminal history.
+We use the COMPAS dataset (Angwin et al., 2022), which contains 11,758 defendants with associated features capturing demographics, information about their current charges, and their prior criminal history.
 We merge the dataset with the experimental data from Lin et al.
-(2020) to represent the human decisions, which contains recidivism decisions on a subset of instances from Angwin et al.
-(2022)’s experiment on laypeople [^4] Here the experimental decisions avoid the problem of performative prediction (Perdomo et al., 2020).
+(2020) to obtain human decisions, in the form of recidivism decisions on a subset of instances from Angwin et al.
+(2022)’s experiment on laypeople.
 Human decisions were elicited on a 30-point probability scale.
@@
 (2020).
-We use Brier score as the payoff function, with the binary payoff-related state $\omega$ indicating whether the defendent gets rearrested within two years.
-We use the features of defendent contained in COMPAS dedaset as the signals in our demonstration: demographic features (age, sex, and race), information about the current charge (type of offense and whether it is a misdemeanor or a felony), prior criminal history (e.g., past number of arrest and charges for several offense categories) and the predicted score from COMPAS system [^5] https://doc.wi.gov/Pages/AboutDOC/COMPAS.aspx.
-We use the hold-out test set to estimate the data-generating process, which defines the joint distribution of state, signals, human decisions and AI predictions.
-Constructing instance-level signals.
-Denote the XGBoost model’s predictions as a random variable $Y^{pred}=f(X)$, where $X$ denotes the random variable of data features and $f(\cdot)$ denotes the model’s predictive function.
-We construct $Z_{f(x)}$ as $\mathbb{1}[Y^{pred}=f(x)]$ for every $x$ and use it to calculate the instance-level agent-complementary information value ($ILIV$) of $Y^{pred}$ on instances $x$ (Definition 3.3).
-We generate the information-based explanations by ILIV-SHAP in this demonstration.
+We use the Brier score as the payoff function, which is commonly used in previous studies on recidivism (Green and Chen, 2019a; Lin et al., 2020; Fogliato et al., 2021).
+The binary payoff-related state $\omega$ indicates whether the defendant gets rearrested within two years.
+We use the features of the defendant contained in the COMPAS dataset as the signals in our demonstration: demographic features (age, sex, and race), information about the current charge (type of offense and whether it is a misdemeanor or a felony), prior criminal history (e.g., past number of arrests and charges for several offense categories), and the predicted score from the COMPAS system [^4] https://doc.wi.gov/Pages/AboutDOC/COMPAS.aspx.
+We take the empirical distribution of the hold-out test set as the data-generating process, which defines the joint distribution of state, signals, human decisions, and AI predictions.
 6.2 Results
-How well does SHAP explain complementary information offered by the AI prediction over what humans already know?
-Figure 3 (A) and (B) compares the distribution of feature-attribution scores from SHAP to those from our ILIV-SHAP.
-We observe multiple discrepancies in feature importance across the methods.
-For instance, in Figure 3(A), the age feature negatively correlates with the AI’s prediction, indicating that younger defendants tend to receive higher predictions.
-However, Figure 3(B) ILIV-SHAP shows no such association between age and the AI’s information value over human decisions.
-This implies that, although age influences the AI’s prediction, on average it does not provide additional information that humans lack.
-Conversely, some features that correlate strongly with the $ILIV$ of the AI prediction are relatively unimportant for the AI’s raw prediction.
-For example, in Figure 3(A), low decile scores predicted by the COMPAS (colored in blue) have a relatively small impact on the AI’s numerical output compared with other features.
-In contrast, Figure 3(B) reveals that some low decile scores yield large positive $ILIV$, suggesting that this feature might offer valuable information for human decision-makers despite not significantly changing the AI’s raw prediction.
-How can an information-based explanation enhance understanding of AI predictions?
-We argue that an information-based explanation (via ILIV-SHAP) can serve as a supplement to saliency-based explanations, providing users a sense of whether the individual prediction can help them make decisions.
-First, an information-based explanation conveys whether the prediction correlates with the payoff-related state.
-For example, Figure 3(C) shows that the model prediction $f(x)$ deviates $+0.383$ from the prior prediction $\text{\bf E}\![{f(X)}]$ on instance 4.
-However, the information-based explanation of the same instance in Figure 3(D) shows that the $ILIV$ of the AI prediction over human decisions does not change much from the $ILIV$ of the prior prediction (which is very low at $0.005$).
-This suggests to the user that focusing on the AI prediction on this instance is not necessary to good decision-making.
-Second, the information-based explanation conveys to the user whether the AI gets useful information beyond their own information from a certain feature.
-In Figure 3(A), the defendent’s prior count record (priors_count=3) changes the AI prediction the most ($+0.186$).
-However, Figure 3(D) shows that priors_count = 3 contributes marginally to the $ILIV$ ($+0.00004$).
-Therefore, ILIV-SHAP conveys that even if priors_count=3 changes the AI prediction significantly, the AI does not necessarily better predict the payoff-related state as a result of it.
+SHAP explanations do not communicate the complementary information of AI predictions.
+Specifically, the changes in information value of AI predictions caused by marginalizing out features, i.e., ILIV-SHAP scores, are not predictable from the changes in AI predictions, i.e., SHAP scores, and the feature values.
+Figure 3 (A) and (B) compare the distribution of feature importance scores from SHAP to those from ILIV-SHAP.
+The scores of prior_counts (the number of prior criminal charges), age, and decile_score (the score from the COMPAS system), which are three features with the largest scores in both SHAP and ILIV-SHAP, show different correlations with the feature values in SHAP and ILIV-SHAP.
+For instance, in Figure 3(A), most of the low values (blue dots) of prior_counts have a negative SHAP score, and most of the high values (red dots) have a positive SHAP score.
+This indicates that when the value of prior_counts is low, marginalizing out prior_counts from the features leads to a higher expected model prediction than not marginalizing out, and vice versa.
+However, this correlation does not hold for ILIV-SHAP.
+In Figure 3(B), the high values of prior_counts appear on both the left side (below zero) and the right side (above zero), indicating having low ILIV-SHAP scores and high ILIV-SHAP scores at the same time.
+This mismatch between the distributions of SHAP scores and ILIV-SHAP scores implies that the impact of features on the information value cannot be directly inferred (at least monotonically) from their impact on model output and their values.
+ILIV-SHAP can augment SHAP on communicating the complementary information of AI predictions.
+We argue that ILIV-SHAP can supplement saliency-based explanations to provide users with a sense of how the AI predictions can help improve their decisions.
+With ILIV-SHAP, users can check whether the prediction contains decision-relevant information by looking at the total $\operatorname{ILIV}$ of the AI prediction, which indicates the maximum improvement in expected payoff that could be achieved by incorporating this prediction.
+Figure 3(D) and (F) show two different cases, where the AI predictions contain substantial human-complementary information value in (D) (shown by the distance between the tail and the head of the Waterfall plot) but very little information in (F).
+ILIV-SHAP also tells the user whether an AI model’s prediction encodes useful complementary information from a particular feature.
+In Figure 3(E), the defendant’s prior count record (prior_counts=1) ranks second among features that impact the AI prediction ($-0.139$).
+Upon seeing this, users might be interested in investigating how the AI uses prior_counts=1.
+However, Figure 3(F) shows that prior_counts=1 contributes marginally to the decision-relevant information in AI predictions (about $+0.007$).
+Thus, with ILIV-SHAP, users can see that even if prior_counts=1 changes the AI prediction significantly, the AI does not predict more accurately as a result of it.
 7 Discussion and Limitations
 We propose a decision-theoretic framework for assigning value to information in human-AI decision-making.
-Our methods quantify the additional information value of any signals over an agent’s decisions.
-The three demonstrations show how quantified information value can be used in model selection, empricial evaluation, and explanation design.
-These demonstrations are just a few of many possible uses cases.
-For example, alternative explanation approaches could be compared via their information value to human decisions.
-Information value analysis could drive elicitation of human private signals or decision rules to further improve the pairing.
-New explanation strategies could integrate visualized information value with conventional depictions of feature importance.
-Information value cannot definitively establish that particular signals were used by a human.
+Our methods quantify the additional information value of any signals over an agent’s decisions, and can be applied to any combination of artificial or human agent judgments.
+The three demonstrations we provide show how quantified information value can drive model selection, empirical evaluation, and explanation design.
+These are just a few of many possible use cases.
+For example, explanation approaches could be compared in terms of their information value to human decisions.
+Information value analysis could also drive elicitation of human private signals upon identifying that human decisions contain AI-complementary information.
+New explanation strategies based on ILIV-SHAP could more directly integrate visualized information value with conventional depictions of feature importance.
+It is important to note that when using ACIV to understand complementary information over human decisions, it cannot be interpreted as definitively establishing that particular signals were used by the human.
 It is always possible that the human has other private signals offering equivalent information to a feature being analyzed.
-Our framework cannot account for private unobservable signals even though they might have strong correlation with the payoff-related state and agent deicions.
-However, quantifying the value of observed information is a tool toward learning about information that may exist beyond a problem definition.
+This raises a question of whether it is definitive about use if it is an AI as the agent instead of a human.
+However, quantifying the value of observed information is an important first step toward learning about information that may exist beyond an explicit problem definition.
 Our framework quantifies the best-attainable performance improvement from integrating signals in decisions.
-This does not necessarily mean that empirically those signals will lead to agents better performing.
-However, the motivating idea is that Bayesian decision theory provides a theoretical basis that can be adapted to support informative comparisons to human behavior.
-For example, if we suspect a human decision-maker uses AI predictions and their own predictions strictly monotonically, we could constrain the Bayesian decision-maker to only make monotonic decisions with AI prediction and their own predictions.
+This does not necessarily mean that highlighting or otherwise emphasizing those signals for humans will lead to them performing better.
+However, the basis of our framework in Bayesian decision theory means that it provides a theoretical basis that can be adapted to support comparisons to human behavior to drive learning (Hullman and Gelman, 2021).
+For example, if we suspect a human decision-maker uses AI predictions and their own predictions strictly monotonically, we could constrain the rational decision-maker to only make monotonic decisions with AI prediction and their own predictions.
 Appendix A The Combinatorial Nature of the Value of Signals
@@
 The Shapley value calculates the average of the marginal contribution of a basic signal $\mathbf{\Sigma}_{i}$ in every combination of signals.
-$\phi(\Sigma_{i})=\frac{1}{n}\sum_{V\subseteq\{V_{1},\ldots,V_{n}\}/\{\Sigma_{i}\}}{(n-1)\choose|V|}^{-1}(ACIV(V\cup\{\Sigma_{i}\};D^{b})-ACIV(V;D^{b}))$ (1)
+$\phi(\Sigma_{i})=\frac{1}{n}\sum_{V\subseteq\{V_{1},\ldots,V_{n}\}/\{\Sigma_{i}\}}{(n-1)\choose|V|}^{-1}(\operatorname{ACIV}(V\cup\{\Sigma_{i}\};D^{b})-ACIV(V;D^{b}))$ (10)
 The Shapley value suggests how much information value of the basic signal is unexploited by the human decision-maker on average in all combinations.
-The following algorithm provides a polynomial-time approximation of the Shapley value of $ACIV$. Under the assumption of submodularity, it orders the signals the same as the Shapley value.
-Algorithm 1 Algorithm 1
+The following algorithm provides a polynomial-time approximation of the Shapley value of $\operatorname{ACIV}$.
+Under the assumption of submodularity, it orders the signals the same as the Shapley value.
+Algorithm 1 Greedy algorithm for marginal gain of $\operatorname{ACIV}$
 1:  $V^{*}=\{D^{b}\}$
@@
 3:  for $i=1$ to $n$ do
-4:     $\phi^{\prime}_{j}=ACIV(\Sigma_{j};V^{*})\ for\ each\ j$
+4:     $\phi^{\prime}_{j}=\operatorname{ACIV}(\Sigma_{j};V^{*})\ for\ each\ j$
 5:     $j^{*}=\arg\max_{j\ s.t.\ \Sigma_{j}\notin V^{*}}\phi^{\prime}_{j}$
@@
 10:  $output\ \phi_{j^{*}}$
-Appendix B Robustness analysis in Experiment I
+Appendix B Embeddings for High-dimensional Signals in Decision Problem
+Suppose the data-generating process is estimated from a set of observations $\{(\omega_{i},v_{i})\}_{i=1}^{T}$, i.e.,
+$\hat{\pi}(\omega,v)=\frac{1}{T}\sum_{i=1}^{T}\mathbb{1}_{\omega=\omega_{i}}\cdot\mathbb{1}_{v=v_{i}}$ (11)
+In this case, the sample size $T$ that is required to get a good estimate of DGP is the exponential of the number of dimensions of the signal $v$.
+The high-dimensional signals, such as images and text, that are usual in human-AI decision problems are not computationally feasible to be quantified under our framework.
+We propose a systematic procedure to pick the optimal reduction model to represent the high-dimensional signal space in a low-dimensional embedding space.
+The optimal reduction model and its generated embeddings should fulfill two properties.
+First, the generated signals should be as informative as possible for the decision task, such that they maintain the maximum information value contained in the original signals.
+Second, the derived DGP that is implied by the embedding signals should be feasible to be estimated from the observed dataset.
+We quantify the above objectives by a framework based on Train/Test split.
+Algorithm 2 Embedding high-dimensional signals in decision problems
+1:  Input: Observed dataset $\{(\omega_{i},v_{i})\}_{i=1}^{T}$, Candidate model generator $C$.
+2:  $R^{*}=0$
+3:  Split the dataset into a training set and a testing set.
+4:  repeat
+5:     Generate a candidate model $c$ from $C$.
+6:     Generate embedding $v^{c}_{i}=c(v_{i})$ on each data point of both sets.
+7:     Estimate the DGP $\hat{\pi}$ implied by the new signals $v^{c}_{i}$ by Equation 11.
+8:     Calculate rational decision rule:   $d^{r}(v^{c})=\arg\max_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\pi(\omega|v^{c})}\![{S(d,\omega)}]$.
+9:     Calculate training performance $R_{training}=\text{\bf E}_{(v^{c},\omega)\sim training\_set}\![{S(d^{r}(v^{c}),\omega)}]$.
+10:     Calculate testing performance $R_{testing}=\text{\bf E}_{(v^{c},\omega)\sim testing\_set}\![{S(d^{r}(v^{c}),\omega)}]$.
+11:     if $R_{training}-R_{testing}>\tau$ then
+12:        go to step 4.
+13:     end if
+14:     if $R^{*}<R_{training}$ then
+15:        $c^{*}=c$.
+16:        $R^{*}=R_{training}$.
+17:     end if
+18:  until no more new candidate model in $C$ or $R^{*}$ converages
+19:  Output: $c^{*}$
+The threshold $\tau$ represents the tolerance in the exchangeability between the signal-state tuples in the training set and the testing set.
+When the candidate model space is infinite, the optimal reduction model can still be found if the candidate space is a lattice that is defined by the informativeness of their generated embeddings, e.g., a class of embedding models with different parameter p in the dropout layer.
+Appendix C Robustness Analysis in Experiment I
 Figure 4: Robust analysis for experiment I on all V-shaped payoff functions.
 The kink $\mu$ is shown on the x-axis.
-Each subplot displays the difference between the $ACIV$ on the row model over human decisions and the $ACIV$ on the column model over human decisions.
+Each subplot displays the difference between the $\operatorname{ACIV}$ on the row model over human decisions and the $\operatorname{ACIV}$ on the column model over human decisions.
+A positive value (colored in blue) at $\mu$ indicates the model on the row contains more informative than the model on the column under the evaluation of V-shaped scoring rule with kink $\mu$.
 The subplots are symmetric along the diagonal, e.g., $(1,2)$ subplot and $(2,1)$ subplot display the same distribution with opposite signs.
-Appendix C More examples of ILIV-SHAP
+Appendix D More examples of ILIV-SHAP
 (a) SHAP on instance 0

### v4

@@
 Abstract
-Multiple agents—including humans and AI models—are often paired on decision tasks with the expectation of achieving complementary performance, where the combined performance of both agents outperforms either one alone.
-However, knowing how to improve the performance of a human-AI team is often difficult without knowing more about what particular information and strategies each agent employs.
-We provide a decision-theoretic framework for characterizing the value of information—and consequently, opportunities for agents to better exploit available information–in AI-assisted decision workflows.
-We demonstrate the use of the framework for model selection, empirical evaluation of human-AI performance, and explanation design.
-We propose a novel information-based explanation technique that adapts SHAP, a saliency-based explanation to explain information value in decision making.
+Multiple agents—including humans and AI models—are increasingly combined to make decisions with the expectation of achieving complementary performance, where the decisions they make together outperform those made individually.
+However, knowing how to improve the performance of collaborating agents is often difficult without knowing more about what particular information and strategies each agent employs.
+With a focus on human-AI pairings, we contribute a decision-theoretic framework for characterizing the value of information—and consequently, opportunities for agents to better exploit available information–in AI-assisted decision workflows.
+We present a novel explanation technique (ILIV-SHAP) that adapts SHAP explanations to highlight human-complementing information.
+We validate the effectiveness of the framework and ILIV-SHAP through a study of human-AI decision-making.
+We show that our measure of complementary information can be used to identify which AI model will best complement human decisions.
+We also find that presenting ILIV-SHAP with AI predictions leads to reliably greater reductions in error over non-AI assisted decisions more than vanilla SHAP.
+[^1] The code to calculate the main quantities in our framework and reproduce the experimental results is available at https://osf.io/p2qzy/?view_only=bf39de5d96f047f69e45ffd42689ebf9.
 1 Introduction
-As the performance of artificial intelligence (AI) for domain-specific predictions improves, workflows in which human and AI models are combined to make decisions in medicine, finance, and law, among other domains.
-Often, statistical models can be developed that exceed the accuracy of human experts on average (Ægisdóttir et al., 2006; Grove et al., 2000; Meehl, 1954).
-However, whenever humans have access to additional information over the AI, there is potential to achieve complementary performance by pairing the two, i.e., better performance than either the human or AI alone.
-For example, a physician may have access to additional information that may not be captured in tabular electronic health records or other structured data (Alur et al., 2024).
+As the performance of artificial intelligence (AI) models continues to improve across domains, workflows in which human experts and AI models are paired for decision-making are sought in medicine, finance, and law, among others.
+Statistical models can often exceed the accuracy of human experts on average (Ægisdóttir et al., 2006; Grove et al., 2000; Meehl, 1954).
+However, whenever humans have access to additional information over an AI model, there is potential to achieve complementary performance by pairing the two, i.e., better performance than either the human or AI alone.
+For example, a physician may have access to information that is not captured in structured health records (Alur et al., 2024).
 Evidence of complementary performance between humans and AI nonetheless remains limited, with many studies showing that human-AI teams underperform an AI alone (Buçinca et al., 2020; Bussone et al., 2015; Green and Chen, 2019b; Jacobs et al., 2021; Lai and Tan, 2019; Vaccaro and Waldo, 2019; Kononenko, 2001).
-A solid understanding of such results is limited by the fact that most analyses of human-AI decision-making focus on ranking the performance of human-AI teams or each individually using measures based on post hoc decision accuracy (Passi and Vorvoreanu, 2022).
-This approach is problematic for several reasons.
-First, it does not account for the best achievable performance based on the information available at the time of the decision (Kleinberg et al., 2015; Guo et al., 2024; Rambachan, 2024).
-Second, it cannot provide insight into the potential for available information to improve decision performance, making it difficult to design interventions that improve the team’s performance.
-In contrast, identifying information complementarities that contribute to the maximum achievable decision performance of a human and AI model—such as when one of the agents has access to information not contained in the other’s judgments, or has not fully integrated information available in the environment into their judgments—provides actionable information for improving the decision pipeline.
-For example, the degree to which a model provides information that complements human judgments can guide AI model selection.
-Evidence that human experts possess decision-relevant information over AI predictions motivates further data collection to improve the AI model.
-Evidence that model predictions contain decision-relevant information that humans do not exploit can guide the design of explanations highlighting complementary information.
+A rigorous understanding of such results is limited by the fact that most analyses of human-AI decision-making focus on ranking the performance of human-AI teams or agents individually using measures based on post-hoc decision accuracy (Passi and Vorvoreanu, 2022).
+However, such approaches do not account for the best achievable performance given the information available at the time of the decision (Kleinberg et al., 2015; Guo et al., 2024; Rambachan, 2024).
+Moreover, existing approaches cannot provide insight on the potential for available information in the environment to improve decision performance.
+Consequently, it is difficult to design interventions that improve the team’s performance.
+If we could identify information complementarities that can be exploited to maximize the decision performance of a human and AI model—such as when one of the agents has access to information not contained in the other’s judgments, or has not fully integrated contextually-available information in their judgments—we can intervene to improve decision-making.
+For example, if we can identify how much complementary information over human judgments an AI model provides, we can use this knowledge to guide model selection or motivate further data collection to improve the model.
+Conversely, finding evidence that model predictions contain decision-relevant information that humans do not exploit can motivate the design of explanations highlighting complementary information.
 We contribute a decision-theoretic framework for characterizing the value of information available within an AI-assisted decision workflow.
 In our framework, information is considered valuable to a decision-maker to the extent that it is possible, in theory, to incorporate it into their decisions to improve performance.
-Specifically, our approach analyzes the expected marginal payoff gain from best case (Bayes rational) use of additional information over best case use of the information already encoded in agent decisions for a given decision problem.
-The intuition behind this approach is that any information that is used by the agents will eventually reveal itself through variation in their decisions.
-We identify the value of the information in agent (human, AI, or human-AI) decisions by offering them as a signal to a Bayesian rational decision-maker.
+Specifically, our approach analyzes the expected marginal payoff gain from best case (Bayes rational) use of additional information over best case use of the information already encoded in agent decisions.
 While we focus on combinations of humans and AI agents, the methods we propose can be applied to any combination of agents.
@@
 The first—global human-complementary information value—calculates the value of a new piece of information to an agent over all of its possible realizations among all instances.
-The second—instance-level human-complementary information value—identifies opportunities for decision-makers to better use instance-level information.
-We demonstrate the framework by applying the metrics to three decision-making tasks where AI models serve as human decision-making assistants [^1] Code to replicate our experiments is available at https://osf.io/p2qzy/?view_only=ec06600d06cd4e59bb6051f992e54c08: chest X-ray diagnosis (Rajpurkar et al., 2018; Johnson et al., 2019), deepfake detection (Dolhansky et al., 2020; Groh et al., 2022), and recidivism prediction (Angwin et al., 2022; Dressel and Farid, 2018).
-First, we demonstrate the framework’s utility in model selection by evaluating how well different AI models complement human decision-makers, showing how even among models with similar accuracy, some models strictly offer more complementary information than others across decision problems.
-Next, we use the framework to empirically evaluate how providing AI assistance (alongside instance-level features) helps humans exploit available information for decision-making.
-Lastly, we demonstrate use of the framework to design explanations by extending SHAP (Lundberg and Lee, 2017) to highlight the portion of an AI’s explanation that complements human information.
+The second—instance-level human-complementary information value—identifies opportunities for decision-makers to better use instance-level information such as specific AI predictions.
+We derive a new explanation technique (ILIV-SHAP) that reveals how data features influence the value of complementary information for an individual prediction.
+To evaluate these tools, we contribute the results of a crowdsourced between-subjects experiment in which humans make decisions with and without AI models with varying human-complementary information and with access to different explanation approaches.
+We find that an AI model with more human-complementary information leads to greater improvements in human-AI team performance over the human-alone baseline.
+We also find that adding an ILIV-SHAP explanation to a traditional SHAP explanation leads to greater improvements in human-AI performance over the human-alone baseline than only the SHAP explanation or no explanation.
+We demonstrate further uses of the framework for model selection, empirical evaluation of human-AI performance, and explanation design on existing datasets spanning medical diagnosis, human deepfake detection, and recidivism prediction (see Appendices E, D and F).
+We also include an observational study on the dataset from Vodrahalli et al.
+(2022) which identified the AI model with more human-complementary information helps human-AI teams achieve greater improvements in performance over the human-alone baseline in Appendix C.
 2 Related work
@@
 (2023) propose a hypothesis testing framework to evaluate the added value of human expertise over AI forecasts.
-Our work shares the motivation of incorporating human expertise, but targets a slightly broader scope by quantifying the information value for all available signals and agent decisions in a human–AI decision pipeline.
-3 Methodology
+Our work shares the motivation of incorporating human expertise, but targets a slightly broader scope by quantifying the information value for all available signals and agent decisions in a specific scenario of human–AI decision tasks.
+3 Framework
 Our framework takes as input a decision problem associated with an information model and outputs the value of information of any available signals to any agent, conditioning on the existing information in their decisions within a Bayesian decision theoretic framework.
+The intuition behind our use of statistical decision theory is that while human decision-makers may deviate from Bayesian rational decision-making, understanding the possible value of a signal requires defining the best attainable decision performance with that signal.
 Our framework provides two separate functions to quantify the value of information: one globally across the data-generating process, and one locally in a realization drawn from the data-generating process.
@@
 Alternatively, they can be defined algorithmically, by strategically applying dimensionality reduction to generate low-dimensional embeddings.
-We introduce an algorithm for identifying an “optimal” reduction model in decision problems in Appendix B.
-We demonstrate these two methods in Section 4 and Section 5 respectively.
+We introduce an algorithm for identifying a set of optimal “coarsened” signals given a set of observed samples in Appendix B.
 Information value.
 Our framework quantifies the value of information in a signal $V$ as the expected payoff improvement of an idealized agent who has access to $V$ in addition to some baseline information set.
-We suppose a rational Bayesian DM who knows the prior probability of the state and conditional distribution of signals (i.e., the data-generating process), observes a signal realization, updates their prior to arrive at posterior beliefs, and then chooses a decision to maximize their expected payoff given their posterior belief.
+This corresponds to a rational Bayesian DM who knows the prior probability of the state and conditional distribution of signals (i.e., the data-generating process), observes a signal realization, updates their prior to arrive at posterior beliefs, and then chooses a decision to maximize their expected payoff given their posterior belief.
 Formally, given a decision task with payoff function $S$ and an information model $\pi$, the rational DM’s expected payoff given a (set of) signal(s) $V$ is
@@
 This baseline defines the maximum expected payoff that can be achieved with no information.
-Bayesian decision theory quantifies the information value of $V$ by the payoff improvement of $V$ over the payoff obtained without information.
+Bayesian decision theory quantifies the information value of $V$ as the payoff improvement of $V$ over the payoff obtained without information.
 Definition 3.1.
@@
 3.1 Agent-Complementary Information Value
-With the above definitions, it is possible to measure the additional value that new signals can provide over the information already captured by an agent’s decisions.
+With the above definitions, it is possible to measure the best case additional value that new signals can provide over the information already captured by an agent’s decisions.
 Here, agent may refer to a human, an AI system, or a human–AI team.
@@
 In the reverse scenario, if human decisions serve as $V$ and AI predictions are $D^{b}$, we can measure how much humans can contribute over the information captured in the AI predictions.
-We demonstrate uses of $\operatorname{ACIV}$ in Section 4 and Section 5.
+We demonstrate uses of $\operatorname{ACIV}$ in Appendix D and Appendix E.
 3.2 Instance-level Agent-complementary Information Value
@@
 To calculate instance-level information value, we rely on the performance of the Bayesian rational agent on the conditional distribution $\pi(\sigma_{1},\ldots,\sigma_{n},\omega,d^{b}|v)$ instead of the whole distribution $\pi(\sigma_{1},\ldots,\sigma_{n},\omega,d^{b})$.
-Formally, given a decision task with payoff function $S$ and information model $\pi$, the expected payoff of the rational DM given signal $V=v$ on instances where $V=v$ is
-$\mathrm{r}^{v,\pi,S}(v)=\text{\bf E}_{\omega\sim\pi(\omega|v)}\![{S(d^{r}(v),\omega)}]$ (7)
-where $d^{r}(v)$ is the Bayesian optimal decision on receiving $v$ as defined in Equation 3.
+Formally, given a decision task with payoff function $S$ and information model $\pi$, the expected payoff of the rational DM given signal $V=v^{\prime}$ on instances with the signal realization $V=v$ is
+$\mathrm{r}^{v,\pi,S}(v^{\prime})=\text{\bf E}_{\omega\sim\pi(\omega|v)}\![{S(d^{r}(v^{\prime}),\omega)}]$ (7)
+where $d^{r}(v)$ is the Bayesian optimal decision on receiving $v^{\prime}$ as defined in Equation 3.
 If we consider the agent decisions in addition to the realization $v$, the rational DM’s expected payoff on instances where $V=v$ can be written as
-$\mathrm{r}^{v,\pi,S}(v;D^{b})=\text{\bf E}_{(d^{b},\omega)\sim\pi(d^{b},\omega|v)}\![{S(d^{r}(v\cup d^{b}),\omega)}]$ (8)
+$\mathrm{r}^{v,\pi,S}(v^{\prime};D^{b})=\text{\bf E}_{(d^{b},\omega)\sim\pi(d^{b},\omega|v)}\![{S(d^{r}(v^{\prime}\cup d^{b}),\omega)}]$ (8)
+Note that we allow $v^{\prime}\neq v$, i.e.
+the signal $v^{\prime}$ observed by the rational DM being different from the actual realization $v$.
+When $v^{\prime}\neq v$, the rational DM is misinformed, thus $\mathrm{r}^{v,\pi,S}(v^{\prime})\leq\mathrm{r}^{v,\pi,S}(v)$.
+$\operatorname{ILIV}$ in Definition 3.3 quantifies this improvement in payoff when the rational DM observes the actual $v$.
 Definition 3.3.
-Given a decision task with payoff function $S$ and an information model $\pi$, we define the instance-level agent-complementary information value ($\operatorname{ILIV}$) of signal $V=v$ on instances where $V=v$ as:
-$\operatorname{ILIV}^{v,\pi,S}(v;D^{b})=\mathrm{r}^{v,\pi,S}(v;D^{b})-\mathrm{r}^{v,\pi,S}(\emptyset;D^{b})$ (9)
-where $\mathrm{r}^{v,\pi,S}(\emptyset;D^{b})$ represents the rational DM’s expected payoff on instances where $V=v$ with only the exisiting information encoded in agent decisions.
-Taking the expectation of $\operatorname{ILIV}$ over $V$ recovers the global agent-complementary information value ($\operatorname{ACIV}$), i.e.,
-$\operatorname{ACIV}^{\pi,S}(V;D^{b})=\text{\bf E}_{v\sim\pi(v)}\![{\operatorname{ILIV}^{v,\pi,S}(v;D^{b})}]$
+Given a decision task with payoff function $S$ and an information model $\pi$, we define the instance-level agent-complementary information value ($\operatorname{ILIV}$) of signal $V=v^{\prime}$ on instances where $V=v$ as:
+$\operatorname{ILIV}^{v,\pi,S}(v^{\prime};D^{b})=\mathrm{r}^{v,\pi,S}(v^{\prime};D^{b})-\mathrm{r}^{v,\pi,S}(\emptyset;D^{b}).$ (9)
+$\mathrm{r}^{v,\pi,S}(\emptyset;D^{b})$ represents the expected rational payoff on instances with $V=v$, where the rational DM only knows the information encoded in agent decisions $D^{b}$ without knowing any realizations of $V$.
+$\operatorname{ILIV}$ maximizes when the signal does not misinform, i.e., $\operatorname{ILIV}^{v,\pi,S}(v;D^{b})\geq\operatorname{ILIV}^{v,\pi,S}(v^{\prime};D^{b}),\text{for }v^{\prime}\in\mathbf{V}$.
 3.2.1 ILIV-SHAP Information-based Explanation
@@
 The scores $\phi_{i}(f,\mathbf{x})$ output by SHAP construct an explanation model for a model output, which quantifies the expected counterfactual change in the model output caused by the feature $x_{i}$.
-ILIV-SHAP extends SHAP to give an explanation model on how features impact the decision-relevant information value of an individual model output.
+ILIV-SHAP extends SHAP to give an explanation model of how features impact the decision-relevant information value of an individual model output.
 Definition 3.4 (ILIV-SHAP).
@@
 where $\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(g_{f}(\mathbf{x}^{\prime});D^{b})$ denotes a counterfactual evaluation of $\operatorname{ILIV}$, which quantifies the expected payoff gain from additionally knowing $g_{f}(\mathbf{x}^{\prime})$ on the instances where $f(\mathbf{x})$ is realized.
-This counterfactual version of $\operatorname{ILIV}$ is guaranteed to achieve the maximum at $\mathbf{x}^{\prime}$ where $f(\mathbf{x})=g_{f}(\mathbf{x}^{\prime})$, i.e., the features missing from $\mathbf{x}^{\prime}$ have no impact on $f(\mathbf{x})$.
+This counterfactual version of $\operatorname{ILIV}$ is guaranteed to achieve the maximum when $g_{f}(\mathbf{x}^{\prime})$ informs $f(\mathbf{x})$, i.e., $f(\mathbf{x})=g_{f}(\mathbf{x}^{\prime})$ – the features missing from $\mathbf{x}^{\prime}$ have no impact on $f(\mathbf{x})$.
 The explanation model offered by ILIV-SHAP is grounded in the following two properties.
@@
 Formally, for any $f$ and any $\mathbf{x}$, $\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(f(\mathbf{x});D^{b})=\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(\text{\bf E}\![{f(X)}];D^{b})+\sum_{i=1}^{m}\phi^{\operatorname{ILIV}}_{i}(f,\mathbf{x})$.
-We demonstrate use of ILIV-SHAP in Section 6.
-3.3 Robustness Analysis of Information Order
-Our approach assumes the specification of a decision problem on which agents’ decisions and use of information are evaluated.
-However, ambiguity around the appropriate decision problem, and in particular, the appropriate scoring rule, is not uncommon in human-AI decision settings.
-Ambiguity can arise as a result of challenges in eliciting utility functions and/or variance in these functions across decision-makers or groups of instances; for example, doctors penalize certain false negative results differently when diagnosing younger versus older patients (Mclaughlin and Spiess, 2023).
-Blackwell’s comparison of signals (Blackwell et al., 1951) is an appropriate tool for addressing ambiguity about the payoff function, as it defines a (set of) signal $V_{1}$ as more informative than $V_{2}$ if $V_{1}$ has a higher information value on all possible decision problems.
+We demonstrate use of ILIV-SHAP in Appendix F.
+3.3 Robust Analysis of Information Order
+Our approach assumes a decision problem as input and evaluates agents’ decisions and use of information on this problem.
+However, evaluators may face ambiguity around the appropriate decision problem specification, and in particular, the appropriate scoring rule.
+In particular, ambiguity can arise in payoff functions; doctors, for example, penalize false negative results differently when diagnosing younger versus older patients (Mclaughlin and Spiess, 2023).
+Blackwell’s comparison of signals (Blackwell et al., 1951) is an ideal tool for addressing ambiguity about the payoff function, as it defines a signal $V_{1}$ as more informative than $V_{2}$ if $V_{1}$ has a higher information value on all possible decision problems.
 We identify this partial order by decomposing the space of decision problems via a basis of proper scoring rules [^2] For rational DMs, any decision problem can be represented by an equivalent proper scoring rule in Equation 1, such that the partial order defined via proper scoring rules also applies to the corresponding decisin tasks.
@@
 then $V_{1}$ is Blackwell more informative than $V_{2}$.
-Extending this to agent-complementary information value, we say that $V_{1}$ offers a higher complementary value than $V_{2}$ under the Blackwell order if
+Extending this to ACIV, $V_{1}$ offers a higher complementary value than $V_{2}$ under the Blackwell order if
 $\operatorname{ACIV}^{\pi,S_{\mu}}(V_{1};D^{b})\geq\operatorname{ACIV}^{\pi,S_{\mu}}(V_{2};D^{b}),\forall\mu\in(0,1)$
-This definition allows us to rank signals (or sets of signals) without needing to commit to a specific payoff function. We present a use case in Section 4.
-4 Experiment I: Model Comparison on Chest Radiograph Diagnosis
-We apply our framework to a well-known cardiac dysfunction diagnosis task (Rajpurkar et al., 2018; Tang et al., 2020; Shreekumar, 2025).
+This definition allows us to rank signals (or sets of signals) without needing to commit to a specific payoff function. We present a use case in Appendix D.
+4 Demonstrations
+We demonstrate uses of our framework on three real-world decision-making tasks: medical diagnosis, deepfake video detection, and recidivism prediction.
+First, we demonstrate the framework’s utility in model selection by evaluating how well different cardiac dysfunction prediction models trained on chest X-rays complement radiologists’ predictions (Rajpurkar et al., 2018; Johnson et al., 2019), showing how even among models with similar accuracy,some models strictly offer more complementary information than others over the radiologists.
+Next, we show how the framework can be used to evaluate how well humans exploit available features in a deepfake detection task (Dolhansky et al., 2020; Groh et al., 2022), how this compares to how an AI models exploits the information, and how providing AI assistance affects the human use.
+Suprisingly, even though AI assistance helps human-AI teams achieve higher decision accuracy, the $\operatorname{ACIV}$s of the video-level features did not change much compared to the $\operatorname{ACIV}$s over human-only predictions, indicating that simply displaying AI’s prediction did not help humans exploit the available information.
+Lastly, we demonstrate how ILIV-SHAP explanations surface unique insight into information value that humans lack relative to conventional SHAP for a recidivism prediction task (Angwin et al., 2022; Dressel and Farid, 2018).
+Detailed descriptions and results of all demonstrations are in Appendices E, D and F.
+5 Experiment
+We use a preregistered between-subjects online experiment to answer two questions:
+  1. 1.
+    Can ACIV identify which AI model will result in the best human-AI decision-making?
+  2. 2.
+    Does ILIV-SHAP improved human-AI decision-making over SHAP alone?
+Our experiment assigns participants to one of two AI models with varying ex-ante ACIV and one of three explanation conditions: ILIV-SHAP and SHAP, SHAP, and no explanation.
+We hypothesize that the higher ACIV model will lead to better human-AI decisions and presenting ILIV-SHAP explanations for this model will lead to better human-AI decisions than the other explanation types.
+Metric AI1 AI2
+MAPE 14.30% 14.51%
+R2 0.81 0.81
+$\operatorname{ACIV}$
+(in MSE)
+$6.5\times 10^{8}$ $3.7\times 10^{8}$
+$\operatorname{ACIV}$
+(in MAPE)
+4.61% 2.00%
+Figure 1: Left: AI performance and ACIV. Right: Posterior predicted reduction in the human-AI team’s absolute percentage error (APE) with 95% confidence intervals.
+Task and data.
+We study an AI-assisted house price prediction, following other recent crowdsourced studies (e.g., Chiang and Yin (2021); Hemmer et al.
+(2022); Poursabzi-Sangdeh et al.
+(2021); Holstein et al.
+(2023)) .
+This task does not require any domain knowledge, making it suitable for a broad participant population.
+Additionally, the task is complex enough to benefit from AI assistance.
+We use the Ames, Iowa Housing Dataset (De Cock, 2011).
+Participants and procedure.
+We recruited 421 US-based participants via the crowdsourcing platform Prolific.
+Each participant was randomly assigned to 12 houses in the Ames Iowa Housing Dataset and performed a sequence of 24 trials (where two decisions were elicited for each house, one without AI assistance and one with), split evenly across a human-alone prediction phase and a human-AI prediction phase.
+In the first 12 trials, participants were asked to predict the house price of the assigned house using six features of the house, including year built, the living area above ground in square feet, the size of garage in car capacity, the number of fireplaces, year remodeled, and ratings of the house’s material and finish.
+We intentionally reduced the interpretability of the year remodeled and material and finish by witholding what variables they represented and rescaling their values.
+In the second 12 trials, participants were asked to revise their guesses in the first round after seeing the AI’s prediction and explanation.
+We incentivized participants with a base reward of $3.00$ USD and a bonus based on the mean squared error (MSE) between their guesses and the true house price in the 24 trials: $\textit{bonus}=3.00-\frac{\textit{MSE}}{3\times 10^{9}}$.
+AI models and explanations.
+Participants were assigned to one of six experimental conditions resulting from crossing the AI models (2 levels, where AI1 has high $\operatorname{ACIV}$ and AI2 has low $\operatorname{ACIV}$) with the explanation type (3 levels: ILIV-SHAP + SHAP, SHAP, and no explanation).
+We designed AI1 and AI2 so as to vary their potential to complement humans: AI1 was trained with the six features and the true house price, with noise added to the six features.
+On the other hand, AI2 was trained with the four human-interpretable features and the true house price, such that there would be less potential for the predictions to add unique information over human judgments.
+We used the same training data for both AI models and ensured that the two AI models achieved similar performance (Figure 1), so that model accuracy differences would not confound our results.
+In the SHAP explanation, we used the SHAP values of the six features to explain the AI’s prediction.
+To generate the ILIV-SHAP explanations, we used independent human decisions we collected on Prolific for the dataset using the same set-up prior to running the official study.
+We present ILIV-SHAP explanations of the six features that ranked and highlighted the features where the human-complementary information the AI’s prediction gets from the feature is higher than a threshold (see Appendix I for explanation examples).
+Evaluation metrics.
+We report the $\operatorname{ACIV}$ of the AI models in the mean squared error (MSE) [^3] We included MSE since it is a proper scoring rule while MAPE is not.
+and the mean absolute percentage error (MAPE).
+To measure the human unassisted and human-AI team performance, we fit a preregistered Bayesian hierarchical regression model [^4] https://aspredicted.org/jfm6-rrs3.pdf to the percentage error (PE) between the human’s prediction and the true house price, $\textit{PE}=(d-\omega)/\omega$ using R’s brms package (Bürkner, 2017).
+[…]
+[…]
+[…]
+$\nu$ $\sim\text{Gamma}(2,0.1)$
+[…]
+Results.
+Figure 1 confirms the expected ranking of AI models by $\operatorname{ACIV}$.
+AI1, which had access to the two human-uninterpretable features, achieves higher $\operatorname{ACIV}$ than AI2, which only had access to the four human-interpretable features.
+Note the similar predictive performance of the two AI models.
+Posterior predictions of human-AI team performance according to our regression model also align with our expectations based on the complementary information each model provides (Figure 1 far right).
+Collapsing across explanation conditions, AI1 reduces APE by more than AI2 ($6.27\%[5.91\%,6.63\%]$ vs $5.65\%[5.26\%,6.05\%]$ respectively).
+Access to the ILIV-SHAP + SHAP explanation results in a greater reduction in APE for AI1 than AI2 ($7.11\%[6.57\%,7.65\%]$ vs $5.83\%[5.17\%,6.49\%]$ respectively).
+Figure 1 also illustrates the effect of the different explanation conditions on the human-AI team performance.
+When the AI has sufficient complementing information (i.e., AI1), the ILIV-SHAP + SHAP explanation more effectively reduces APE for the human-AI decisions than the SHAP explanation and baseline (no explanation) ($7.11\%[6.57\%,7.65\%]$ for ILIV-SHAP, $5.20\%[4.53\%,5.87\%]$ for SHAP, and $6.50\%[5.86\%,7.14\%]$ for no explanation, all for model AI1).
+6 Discussion and Limitations
+Our decision-theoretic framework quantifies the information value of signals available in a human-AI decision setting over the information value of agent decisions.
+The basis of our framework in Bayesian decision theory means that it provides a theoretical basis that can be adapted to support comparisons to human behavior to drive learning (Hullman and Gelman, 2021), even if human agents deviate considerably from Bayesian rational decision-making.
+We experimentally demonstrate the power of the framework for analyzing and improving human-AI decision-making.
+Our proposed ILIV-SHAP explanation improves performance over an existing state-of-the-art explanation strategy.
+Our work provides a promising methodological framework for an emerging research agenda around optimally combining agents’ information for decisions (e.g., Alur et al.
+(2023, 2024)).
+While we focused on human-AI decisions, the framework an be applied to any combination of AI or human agent judgments.
+One direction for follow-up work is to develop tests that complement the framework by helping evaluators confirm the information that agents are using.
+Note that low ACIV of a signal over human decisions is not necessarily sufficient for concluding that particular signals were used by the human.
+It is often unknown whether a human has other private signals offering equivalent information to a feature being analyzed.
+Appendix A The Combinatorial Nature of the Value of Signals
+When decision-makers are provided with multiple signals, the signals have the combinatorial property by nature.
+Acknowledgd by recent works in decision theory and game theory [Chen and Waggoner, 2016], one signal may have no information value by itself, but it can be complementary to other signals to provide information value.
+For example, two signals $\Sigma_{1}$ and $\Sigma_{2}$ are uniformly random bits and the state $\omega=\Sigma_{1}\oplus\Sigma_{2}$, the XOR of $\Sigma_{1}$ and $\Sigma_{2}$.
+In this case, neither of the signals offers information value on its own, but knowing both leads to the maximum payoff.
+Though we did not explicitly observe the complementation between signals in our survey of human-AI decision-making tasks (see the results of deepfake detection in supplementary materials), we want to note that our framework can be extended to consider this complementation between signals.
+We use the Shapley value [Shapley, 1953] to interpret the contribution to the $\operatorname{ACIV}$ of each basic signal.
+$\phi$ is the average of the marginal contribution of a signal in every combination with other signals.
+$\phi(V)=\frac{1}{n}\sum_{V^{\prime}\subseteq\{\Sigma_{1},\ldots,\Sigma_{n}\}/V}{(n-\lvert V\rvert)\choose\lvert V^{\prime}\rvert}^{-1}(\operatorname{ACIV}^{\pi,S}(V^{\prime}\cup V;D^{b})-\operatorname{ACIV}^{\pi,S}(V^{\prime};D^{b}))$ (10)
+The following algorithm provides a polynomial-time approximation of the Shapley value of $\operatorname{ACIV}$.
+Under the assumption of submodularity, it orders the signals the same as the Shapley value.
+Algorithm 1 Greedy algorithm for marginal gain of $\operatorname{ACIV}$
+1:  $V^{*}=\{D^{b}\}$
+2:  $\Phi^{*}=\{\}$
+3:  for $i=1$ to $n$ do
+4:     $\phi^{\prime}_{j}=\operatorname{ACIV}(\Sigma_{j};V^{*})\ for\ each\ j$
+5:     $j^{*}=\arg\max_{j\ s.t.\ \Sigma_{j}\notin V^{*}}\phi^{\prime}_{j}$
+6:     $\phi_{j^{*}}=\max_{j\ s.t.\ \Sigma_{j}\notin V^{*}}\phi^{\prime}_{j}$
+7:     $add\ \Sigma_{j^{*}}\ to\ V^{*}$
+8:     $add\ \phi_{j^{*}}\ to\ \Phi^{*}$
+9:  end for
+10:  $output\ \phi_{j^{*}}$
+Appendix B Estimating the Data-Generating Process from Observations
+Suppose the data-generating process is estimated from the empirical distribution of the observed realizations of the payoff state value and the signal, $D=\{(\omega_{i},v_{i})\}_{i=1}^{T}$, i.e.,
+$\hat{\pi}_{D}(\omega,v)=\frac{1}{T}\sum_{i=1}^{T}\mathbb{1}_{\omega=\omega_{i}}\cdot\mathbb{1}_{v=v_{i}}$ (11)
+In applying our framework, the rational DM’s decision rule $d^{r}$ should not overfit to the observed dataset.
+This threatens the usefulness of analysis using the DGP.
+To avoid this situation, we can either collect more observations or “coarsen” the signal space, where “coarsen” means reducing the signal space by aggregating similar signals into a single signal.
+We propose a systematic procedure for the latter since it may be too costly or otherwise not possible to collect more observations.
+Our objective is to find the coarsened signal structure that has the largest signal space and avoids the rational decision rule from overlapping.
+Denote the maximum tolerance factor for overfitting as $\epsilon$.
+The rational DM’s decision rule $d^{r}_{\hat{\pi}_{D^{\prime}}}$ that is estimated from a set of observations $D^{\prime}$ should perform similarly on $D^{\prime}$ and on the evaluation of a holdout set $D\setminus D^{\prime}$, i.e.,
+$\text{\bf E}_{(\omega,v)\sim\hat{\pi}_{D^{\prime}}}\![{S(d^{r}_{\hat{\pi}_{D^{\prime}}}(v),\omega)}]-\text{\bf E}_{(\omega,v)\sim\hat{\pi}_{D\setminus D^{\prime}}}\![{S(d^{r}_{\hat{\pi}_{D^{\prime}}}(v),\omega)}]\leq\epsilon*\text{\bf E}_{(\omega,v)\sim\hat{\pi}_{D^{\prime}}}\![{S(d^{r}_{\hat{\pi}_{D^{\prime}}}(v),\omega)}]$ (12)
+Algorithm 2 Estimating the data-generating process from a set of observations
+1:  Input: Observed dataset $C=\{(\omega_{i},v_{i})\}_{i=1}^{T}$, Overfitting tolerance $\epsilon$, Ratio of the holdout set $\delta$.
+2:  randomly split the dataset into a training set $C^{\prime}$ and a holdout set $C\setminus C^{\prime}$ with
+$\lvert C^{\prime}\rvert=(1-\delta)*T$
+3:  $\hat{\pi}_{C^{\prime}}(\omega,v)=\frac{1}{T}\sum_{\hat{\omega},\hat{v}\in C^{\prime}}\mathbb{1}_{\omega=\hat{\omega}}\cdot\mathbb{1}_{v=\hat{v}}$
+4:  $R_{training}=\text{\bf E}_{(v,\omega)\sim C^{\prime}}\![{S(d^{r}_{\hat{\pi}_{C^{\prime}}}(v),\omega)}]$.
+5:  $R_{testing}=\text{\bf E}_{(v,\omega)\sim C\setminus C^{\prime}}\![{S(d^{r}_{\hat{\pi}_{C^{\prime}}}(v),\omega)}]$.
+6:  $bin\_size=\text{the number of unique }v$
+7:  if $R_{training}-R_{testing}>\epsilon*R_{training}$ then
+8:     $bin\_size=bin\_size/2$
+9:     $\{v^{\prime}_{i}\}_{i=1}^{T}=\text{coarsen }\{v_{i}\}_{i=1}^{T}\text{ with binsize }bin\_size$.
+10:     replace $\{v_{i}\}_{i=1}^{T}$ in $C$ with $\{v^{\prime}_{i}\}_{i=1}^{T}$.
+11:     go to step 3.
+12:  end if
+13:  coarsened signal space $\mathbf{V}^{\prime}=\text{unique}(\{v^{\prime}_{i}\}_{i=1}^{T})$
+14:  Output: $\hat{\pi}_{\pi^{C}}$, $\mathbf{V}^{\prime}$
+Denote the optimal coarsened signal space that satisfies Equation 12 as
+[…]
+The above algorithm finds the coarsened signal space $\mathbf{V}^{\prime}$ that has the 2-factor approximation of the optimal coarsened signal space $\mathbf{V}^{*}$, $\lvert\mathbf{V}^{\prime}\rvert\geq\frac{1}{2}\lvert\mathbf{V}^{*}\rvert$, in $O(log(\lvert\mathbf{V}^{*}\rvert))$ time.
+B.1 DGP with high-dimensional signals
+The above algorithm’s time complexity is linear in the number of dimensions of the signal space.
+However, in practice, high-dimensional signals, such as images and text that appear in human-AI decision problems can still be computationally infeasible to quantify with the above algorithm.
+To address this issue, we propose a two-step approach to estimate the DGP from a set of observations.
+First, we use a reduction model to reduce the dimensionality of the signal space.
+Second, we use the estimated DGP using the reduced signal space in the above algorithm.
+The optimal reduction model and its generated embeddings should fulfill two properties.
+First, the generated signals should be as informative as possible for the decision task, such that they maintain the maximum information value contained in the original signals.
+Second, the derived DGP that is implied by the embedding signals should be feasible to estimate from the observed dataset.
+We quantify the above objectives by a similar algorithm based on Train/Test split.
+Appendix C Observational Study on the Effect of ACIV on Human-AI Team Performance
+In this section, we present the results of an observational study on the effect of $\operatorname{ACIV}$ on human-AI team performance using the dataset from Vodrahalli et al. [2022].
+Data description.
+We experiment with the publicly available Human-AI Interactions dataset [Vodrahalli et al., 2022].
+The dataset comprises 34,783 unique predictions from 1,088 different human participants on four different binary prediction tasks (“Art”, “Sarcasm”, “Cities” and “Census”).
+In each of the tasks, human participants provide confidence values about their predictions before ($d^{H}$) and after ($d^{H+AI}$) receiving AI advice from a classifier in the form of the classifier’s confidence values ($d^{AI}$).
+$d^{H}$, $d^{H+AI}$ and $d^{AI}$ are in the range of $[0,1]$.
+Results.
+We present the $\operatorname{ACIV}$ of the AI confidence values ($d^{AI}$) over human decisions ($d^{H}$) measured by the payoff function as the mean squared error (MSE) in Figure 2, and show the reduction of the mean squared error (MSE) of the human-AI team over the human-alone baseline in Table 1.
+We find that between these four tasks, the $\operatorname{ACIV}$ of the AI over human decisions predicts the improvement of the human-AI team performance over the human-alone baseline.
+Figure 2: The $\operatorname{ACIV}$ of the AI confidence values over human decisions for different tasks. The error bars are 95% confidence intervals.
+Table 1: The $\operatorname{ACIV}$ of the AI confidence values over human decisions and the reduction of mean squared error (MSE) of the human-AI team over the human-alone baseline for different tasks.
+Task Human-complementary Info Human MSE Human+AI MSE Reduction in MSE
+Art 0.1772 [0.1747, 0.1798] 0.243 0.1847 0.0583
+Cities 0.0919 [0.0860, 0.0978] 0.1955 0.1615 0.034
+Sarcasm 0.1625 [0.1587, 0.1662] 0.2137 0.1767 0.037
+Census 0.1147 [0.1077, 0.1217] 0.1997 0.1861 0.0136
+Appendix D Demonstration I: Model Comparison on Chest Radiograph Diagnosis
+We apply our framework to a well-known cardiac dysfunction diagnosis task [Rajpurkar et al., 2018, Tang et al., 2020, Shreekumar, 2025].
 We demonstrate how our framework can be used in model evaluation for analyzing how much complementary information value a set of possible AI models offers to the radiology reports written by experts.
@@
 We list the AI predictions alone to show the AI-complementary information value offered by human decisions.
-4.1 Data and Model
-We use data from the MIMIC dataset (Goldberger et al., 2000), which contains anonymized electronic health records from Beth Israel Deaconess Medical Center (BIDMC), a large teaching hospital in Boston, Massachusetts, affiliated with Harvard Medical School.
-Specifically, we utilize chest x-ray images and radiology reports from the MIMIC-CXR database (Johnson et al., 2019) merged with patient and visit information from the broader MIMIC-IV database (Johnson et al., 2023).
+D.1 Data and Model
+We use data from the MIMIC dataset [Goldberger et al., 2000], which contains anonymized electronic health records from Beth Israel Deaconess Medical Center (BIDMC), a large teaching hospital in Boston, Massachusetts, affiliated with Harvard Medical School.
+Specifically, we utilize chest x-ray images and radiology reports from the MIMIC-CXR database [Johnson et al., 2019] merged with patient and visit information from the broader MIMIC-IV database [Johnson et al., 2023].
 The payoff-related state, cardiac dysfunction $\omega\in\{0,1\}$, is coded based on two common tests, the NT-proBNP and the troponin, using the age-specific cutoffs from Mueller et al.
-(2019) and Heidenreich et al.
-(2022).
-We label the radiology reports by a rule-based tool (Irvin et al., 2019) and use the labels as the human decisions (without AI assistance) in the diagnosis task to solve the problem of computational feasibility with high-dimensional textual reports.
+[2019] and Heidenreich et al.
+[2022].
+We label the radiology reports by a rule-based tool [Irvin et al., 2019] and use the labels as the human decisions (without AI assistance) in the diagnosis task to solve the problem of computational feasibility with high-dimensional textual reports.
 The labels are represented by the symptoms as positive, negative, or uncertain, i.e., $d\in\{+,?,-\}$.
-We fine-tuned five deep-learning models on the cardiac dysfunction diagnosis task, VisionTransformer (Alexey, 2020), SwinTransformer (Liu et al., 2021), ResNet (He et al., 2016), Inception-v3 (Szegedy et al., 2016), and DenseNet (Huang et al., 2017).
+We fine-tuned five deep-learning models on the cardiac dysfunction diagnosis task, VisionTransformer [Alexey, 2020], SwinTransformer [Liu et al., 2021], ResNet [He et al., 2016], Inception-v3 [Szegedy et al., 2016], and DenseNet [Huang et al., 2017].
 Our training set contains 12,228 images, and the validation set contains 6,115 images.
@@
 We use the five vision models’ predictions as embeddings of the original images.
-See more detail on the algorithm of choosing the optimal embeddings in Appendix B.
-4.2 Results
+See more detail on the algorithm of choosing the optimal embeddings in LABEL:app:high-dimensional.
+D.2 Results
 Can the AI models complement human judgment?
@@
 The VisionTransformer achieves a higher information value on all V-shaped scoring rules, implying a higher information value on all decision problems.
-5 Experiment II: Behavioral Analysis on Deepfake Detection
-We apply our framework to analyze a deepfake video detection task (Dolhansky et al., 2020), where participants are asked to judge whether a video was created by generative AI, including with the assistance of an AI model.
+Appendix E Demonstration II: Behavioral Analysis on Deepfake Detection
+We apply our framework to analyze a deepfake video detection task [Dolhansky et al., 2020], where participants are asked to judge whether a video was created by generative AI, including with the assistance of an AI model.
 Figure 4: Information value calculated under our framework in the information model defined by the experiment of Groh et al.
-(2022).
+[2022].
 Basic signals include the seven video level features and three types of agent decisions.
@@
 All the payoffs are calculated by $\mathrm{R}^{\pi,S}(\cdot)$, where $\cdot$ is the signal on the y-axis.
-5.1 Data and Model
+E.1 Data and Model
 We define the information model on the experiment data of Groh et al.
-(2022).
+[2022].
 Non-expert participants (n=5,524) were recruited through Prolific and asked to examine the videos.
@@
 They first reviewed the video and reported an initial decision ($D^{\text{Human}}$) without access to the AI model.
-Then, in a second round, they were provided with the recommendation ($D^{\text{AI}}$) of a multitask cascaded convolutional neural network (Zhang et al., 2016), with estimated $65\%$ accuracy on holdout data, and chose whether to change their initial decision.
+Then, in a second round, they were provided with the recommendation ($D^{\text{AI}}$) of a multitask cascaded convolutional neural network [Zhang et al., 2016], with estimated $65\%$ accuracy on holdout data, and chose whether to change their initial decision.
 This produced a final decision ($D^{\text{Human-AI}}$).
@@
 This choice differs from the mean absolute error used by Groh et al.
-(2022), but again we use the quadratic score because it is a proper scoring rule where truthfully reporting the belief maximizes the score.
+[2022], but again we use the quadratic score because it is a proper scoring rule where truthfully reporting the belief maximizes the score.
 We identify a set of features that were implicitly available to all three agents (human, AI, and human-AI).
 Because the video signal is high dimensional, we make use of seven video-level features using manually coded labels by Groh et al.
-(2022): graininess, blurriness, darkness, presence of a flickering face, presence of two people, presence of a floating distraction, and the presence of an individual with dark skin, all of which are labeled as binary indictors.
+[2022]: graininess, blurriness, darkness, presence of a flickering face, presence of two people, presence of a floating distraction, and the presence of an individual with dark skin, all of which are labeled as binary indictors.
 These are the basic signals in our framework.
 We estimate the data-generating process $\pi$ using the realizations of signals, state, first-round human decisions, AI predictions, and second-round human-AI decisions.
-The no-information bound is the same as Section 4 while the full-information bound is defined as the expected payoff of a rational DM who has access to all video-level features and three agents’ decisions.
-5.2 Results
+The no-information bound is the same as Appendix D while the full-information bound is defined as the expected payoff of a rational DM who has access to all video-level features and three agents’ decisions.
+E.2 Results
 How much decision-relevant information do agents’ decisions offer?
@@
 This is consistent with the findings of Guo et al.
-(2024) that humans are bad at distinguishing when AI predictions are correct.
+[2024] that humans are bad at distinguishing when AI predictions are correct.
 How much additional decision-relevant information do the available features offer over agents’ decisions?
@@
 This finding further confirms the hypothesis that humans are simply relying on AI predictions without processing the information contained in them.
-6 Experiment III: Information-based Local Explanation on Recidivism Prediction
+Appendix F Demonstration III: Information-based Local Explanation on Recidivism Prediction
 We apply our framework to a recidivism prediction task, where the decision-maker decides whether to release a defendant. We demonstrate using the framework to augment SHAP.
@@
 The Waterfall plot starts from the prior expected prediction (or ILIV of the prior expected prediction), and then adds features one at a time until reaching the current model output (or ILIV of the current model output).
-See further comparisons between SHAP and ILIV-SHAP in Appendix D.
-6.1 Data and Model
-We use the COMPAS dataset (Angwin et al., 2022), which contains 11,758 defendants with associated features capturing demographics, information about their current charges, and their prior criminal history.
+See further comparisons between SHAP and ILIV-SHAP in Appendix H.
+F.1 Data and Model
+We use the COMPAS dataset [Angwin et al., 2022], which contains 11,758 defendants with associated features capturing demographics, information about their current charges, and their prior criminal history.
 We merge the dataset with the experimental data from Lin et al.
-(2020) to obtain human decisions, in the form of recidivism decisions on a subset of instances from Angwin et al.
-(2022)’s experiment on laypeople.
+[2020] to obtain human decisions, in the form of recidivism decisions on a subset of instances from Angwin et al.
+[2022]’s experiment on laypeople.
 Human decisions were elicited on a 30-point probability scale.
 Merging the two datasets produces 9,269 instances (defendants).
-For the AI model, we trained an extreme gradient boosting (XGBoost) model (Chen and Guestrin, 2016) on a training set with 6,488 instances, achieving an AUC of $0.84$ on a hold-out test set with 2,781 instances.
+For the AI model, we trained an extreme gradient boosting (XGBoost) model [Chen and Guestrin, 2016] on a training set with 6,488 instances, achieving an AUC of $0.84$ on a hold-out test set with 2,781 instances.
 We round the model predictions to the same 30-point probability scale available to study participants in Lin et al.
-(2020).
-We use the Brier score as the payoff function, which is commonly used in previous studies on recidivism (Green and Chen, 2019a; Lin et al., 2020; Fogliato et al., 2021).
+[2020].
+We use the Brier score as the payoff function, which is commonly used in previous studies on recidivism [Green and Chen, 2019a, Lin et al., 2020, Fogliato et al., 2021].
 The binary payoff-related state $\omega$ indicates whether the defendant gets rearrested within two years.
@@
 We take the empirical distribution of the hold-out test set as the data-generating process, which defines the joint distribution of state, signals, human decisions, and AI predictions.
-6.2 Results
+F.2 Results
 SHAP explanations do not communicate the complementary information of AI predictions.
@@
 Thus, with ILIV-SHAP, users can see that even if prior_counts=1 changes the AI prediction significantly, the AI does not predict more accurately as a result of it.
-7 Discussion and Limitations
-We propose a decision-theoretic framework for assigning value to information in human-AI decision-making.
-Our methods quantify the additional information value of any signals over an agent’s decisions, and can be applied to any combination of artificial or human agent judgments.
-The three demonstrations we provide show how quantified information value can drive model selection, empirical evaluation, and explanation design.
-These are just a few of many possible use cases.
-For example, explanation approaches could be compared in terms of their information value to human decisions.
-Information value analysis could also drive elicitation of human private signals upon identifying that human decisions contain AI-complementary information.
-New explanation strategies based on ILIV-SHAP could more directly integrate visualized information value with conventional depictions of feature importance.
-It is important to note that when using ACIV to understand complementary information over human decisions, it cannot be interpreted as definitively establishing that particular signals were used by the human.
-It is always possible that the human has other private signals offering equivalent information to a feature being analyzed.
-This raises a question of whether it is definitive about use if it is an AI as the agent instead of a human.
-However, quantifying the value of observed information is an important first step toward learning about information that may exist beyond an explicit problem definition.
-Our framework quantifies the best-attainable performance improvement from integrating signals in decisions.
-This does not necessarily mean that highlighting or otherwise emphasizing those signals for humans will lead to them performing better.
-However, the basis of our framework in Bayesian decision theory means that it provides a theoretical basis that can be adapted to support comparisons to human behavior to drive learning (Hullman and Gelman, 2021).
-For example, if we suspect a human decision-maker uses AI predictions and their own predictions strictly monotonically, we could constrain the rational decision-maker to only make monotonic decisions with AI prediction and their own predictions.
-Appendix A The Combinatorial Nature of the Value of Signals
-We model the information value of a single signal over the existing information in agent decisions.
-When decision-makers are provided with multiple signals, they might use them in combination.
-Therefore, our definition of information value in Definition 3.2 may overlook the signals’ value in combination with other signals.
-Signals can be complemented [Chen and Waggoner, 2016], i.e, they contain no information value by themselves but a considerable value when combined with other signals.
-For example, two signals $\Sigma_{1}$ and $\Sigma_{2}$ ight be uniformly random bits and the state $\omega=\Sigma_{1}\oplus\Sigma_{2}$, the XOR of $\Sigma_{1}$ and $\Sigma_{2}$.
-In this case, neither of the signals offers information value on its own, but knowing both leads to the maximum payoff.
-To consider this complementation between signals, we use the Shapley value $\phi$ [Shapley, 1953] to interpret the contribution to information gain of each basic signal.
-The Shapley value calculates the average of the marginal contribution of a basic signal $\mathbf{\Sigma}_{i}$ in every combination of signals.
-$\phi(\Sigma_{i})=\frac{1}{n}\sum_{V\subseteq\{V_{1},\ldots,V_{n}\}/\{\Sigma_{i}\}}{(n-1)\choose|V|}^{-1}(\operatorname{ACIV}(V\cup\{\Sigma_{i}\};D^{b})-ACIV(V;D^{b}))$ (10)
-The Shapley value suggests how much information value of the basic signal is unexploited by the human decision-maker on average in all combinations.
-The following algorithm provides a polynomial-time approximation of the Shapley value of $\operatorname{ACIV}$.
-Under the assumption of submodularity, it orders the signals the same as the Shapley value.
-Algorithm 1 Greedy algorithm for marginal gain of $\operatorname{ACIV}$
-1:  $V^{*}=\{D^{b}\}$
-2:  $\Phi^{*}=\{\}$
-3:  for $i=1$ to $n$ do
-4:     $\phi^{\prime}_{j}=\operatorname{ACIV}(\Sigma_{j};V^{*})\ for\ each\ j$
-5:     $j^{*}=\arg\max_{j\ s.t.\ \Sigma_{j}\notin V^{*}}\phi^{\prime}_{j}$
-6:     $\phi_{j^{*}}=\max_{j\ s.t.\ \Sigma_{j}\notin V^{*}}\phi^{\prime}_{j}$
-7:     $add\ \Sigma_{j^{*}}\ to\ V^{*}$
-8:     $add\ \phi_{j^{*}}\ to\ \Phi^{*}$
-9:  end for
-10:  $output\ \phi_{j^{*}}$
-Appendix B Embeddings for High-dimensional Signals in Decision Problem
-Suppose the data-generating process is estimated from a set of observations $\{(\omega_{i},v_{i})\}_{i=1}^{T}$, i.e.,
-$\hat{\pi}(\omega,v)=\frac{1}{T}\sum_{i=1}^{T}\mathbb{1}_{\omega=\omega_{i}}\cdot\mathbb{1}_{v=v_{i}}$ (11)
-In this case, the sample size $T$ that is required to get a good estimate of DGP is the exponential of the number of dimensions of the signal $v$.
-The high-dimensional signals, such as images and text, that are usual in human-AI decision problems are not computationally feasible to be quantified under our framework.
-We propose a systematic procedure to pick the optimal reduction model to represent the high-dimensional signal space in a low-dimensional embedding space.
-The optimal reduction model and its generated embeddings should fulfill two properties.
-First, the generated signals should be as informative as possible for the decision task, such that they maintain the maximum information value contained in the original signals.
-Second, the derived DGP that is implied by the embedding signals should be feasible to be estimated from the observed dataset.
-We quantify the above objectives by a framework based on Train/Test split.
-Algorithm 2 Embedding high-dimensional signals in decision problems
-1:  Input: Observed dataset $\{(\omega_{i},v_{i})\}_{i=1}^{T}$, Candidate model generator $C$.
-2:  $R^{*}=0$
-3:  Split the dataset into a training set and a testing set.
-4:  repeat
-5:     Generate a candidate model $c$ from $C$.
-6:     Generate embedding $v^{c}_{i}=c(v_{i})$ on each data point of both sets.
-7:     Estimate the DGP $\hat{\pi}$ implied by the new signals $v^{c}_{i}$ by Equation 11.
-8:     Calculate rational decision rule:   $d^{r}(v^{c})=\arg\max_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\pi(\omega|v^{c})}\![{S(d,\omega)}]$.
-9:     Calculate training performance $R_{training}=\text{\bf E}_{(v^{c},\omega)\sim training\_set}\![{S(d^{r}(v^{c}),\omega)}]$.
-10:     Calculate testing performance $R_{testing}=\text{\bf E}_{(v^{c},\omega)\sim testing\_set}\![{S(d^{r}(v^{c}),\omega)}]$.
-11:     if $R_{training}-R_{testing}>\tau$ then
-12:        go to step 4.
-13:     end if
-14:     if $R^{*}<R_{training}$ then
-15:        $c^{*}=c$.
-16:        $R^{*}=R_{training}$.
-17:     end if
-18:  until no more new candidate model in $C$ or $R^{*}$ converages
-19:  Output: $c^{*}$
-The threshold $\tau$ represents the tolerance in the exchangeability between the signal-state tuples in the training set and the testing set.
-When the candidate model space is infinite, the optimal reduction model can still be found if the candidate space is a lattice that is defined by the informativeness of their generated embeddings, e.g., a class of embedding models with different parameter p in the dropout layer.
-Appendix C Robustness Analysis in Experiment I
+Appendix G Robustness Analysis in Demonstration I
 Figure 6: Robust analysis for experiment I on all V-shaped payoff functions.
@@
 The subplots are symmetric along the diagonal, e.g., $(1,2)$ subplot and $(2,1)$ subplot display the same distribution with opposite signs.
-Appendix D More examples of ILIV-SHAP
+Appendix H More examples of ILIV-SHAP
 (a) SHAP on instance 0
@@
 (b) ILIV-SHAP on instance 2
+Appendix I Screenshots of the Experiment
+Figure 10: Screenshot of the human-alone trials in the first round of the experiment.
+Figure 11: Screenshot of the no explanation condition in the second round of the experiment.
+Figure 12: Screenshot of the SHAP condition of AI1 in the second round of the experiment.
+Figure 13: Screenshot of the SHAP condition in the second round of the experiment.
+Figure 14: Screenshot of the ILIV-SHAP + SHAP condition of AI1 in the second round of the experiment.
+Figure 15: Screenshot of the ILIV-SHAP + SHAP condition of AI2 in the second round of the experiment.

### v5

@@
 Abstract
-Multiple agents—including humans and AI models—are increasingly combined to make decisions with the expectation of achieving complementary performance, where the decisions they make together outperform those made individually.
-However, knowing how to improve the performance of collaborating agents is often difficult without knowing more about what particular information and strategies each agent employs.
-With a focus on human-AI pairings, we contribute a decision-theoretic framework for characterizing the value of information—and consequently, opportunities for agents to better exploit available information–in AI-assisted decision workflows.
+Multiple agents are increasingly combined to make decisions with the expectation of achieving complementary performance, where the decisions they make together outperform those made individually.
+However, knowing how to improve the performance of collaborating agents requires knowing what information and strategies each agent employs.
+With a focus on human-AI pairings, we contribute a decision-theoretic framework for characterizing the value of information.
+By defining complementary information, our approach identifies opportunities for agents to better exploit available information in AI-assisted decision workflows.
 We present a novel explanation technique (ILIV-SHAP) that adapts SHAP explanations to highlight human-complementing information.
-We validate the effectiveness of the framework and ILIV-SHAP through a study of human-AI decision-making.
-We show that our measure of complementary information can be used to identify which AI model will best complement human decisions.
-We also find that presenting ILIV-SHAP with AI predictions leads to reliably greater reductions in error over non-AI assisted decisions more than vanilla SHAP.
-[^1] The code to calculate the main quantities in our framework and reproduce the experimental results is available at https://osf.io/p2qzy/?view_only=bf39de5d96f047f69e45ffd42689ebf9.
+We validate the effectiveness of ACIV and ILIV-SHAP through a study of human-AI decision-making, and demonstrate the framework on examples from chest X-ray diagnosis and deepfake detection.
+We find that presenting ILIV-SHAP with AI predictions leads to reliably greater reductions in error over non-AI assisted decisions more than vanilla SHAP [^1] The code to calculate the main quantities in our framework and reproduce the experimental results is available at https://osf.io/p2qzy/?view_only=bf39de5d96f047f69e45ffd42689ebf9..
 1 Introduction
@@
 For example, a physician may have access to information that is not captured in structured health records (Alur et al., 2024).
-Evidence of complementary performance between humans and AI nonetheless remains limited, with many studies showing that human-AI teams underperform an AI alone (Buçinca et al., 2020; Bussone et al., 2015; Green and Chen, 2019b; Jacobs et al., 2021; Lai and Tan, 2019; Vaccaro and Waldo, 2019; Kononenko, 2001).
+Evidence of complementary performance between humans and AI nonetheless remains limited, with many studies showing that human-AI teams underperform the AI alone (Buçinca et al., 2020; Bussone et al., 2015; Green and Chen, 2019; Jacobs et al., 2021; Lai and Tan, 2019; Vaccaro and Waldo, 2019; Kononenko, 2001).
 A rigorous understanding of such results is limited by the fact that most analyses of human-AI decision-making focus on ranking the performance of human-AI teams or agents individually using measures based on post-hoc decision accuracy (Passi and Vorvoreanu, 2022).
 However, such approaches do not account for the best achievable performance given the information available at the time of the decision (Kleinberg et al., 2015; Guo et al., 2024; Rambachan, 2024).
-Moreover, existing approaches cannot provide insight on the potential for available information in the environment to improve decision performance.
+Moreover, existing approaches cannot provide insight on the potential for available information in the environment to improve performance.
 Consequently, it is difficult to design interventions that improve the team’s performance.
-If we could identify information complementarities that can be exploited to maximize the decision performance of a human and AI model—such as when one of the agents has access to information not contained in the other’s judgments, or has not fully integrated contextually-available information in their judgments—we can intervene to improve decision-making.
-For example, if we can identify how much complementary information over human judgments an AI model provides, we can use this knowledge to guide model selection or motivate further data collection to improve the model.
-Conversely, finding evidence that model predictions contain decision-relevant information that humans do not exploit can motivate the design of explanations highlighting complementary information.
-We contribute a decision-theoretic framework for characterizing the value of information available within an AI-assisted decision workflow.
+Imagine one could identify information complementarities that can be exploited, such as when one of the agents has access to information not contained in the other’s judgments, or has not fully integrated contextually-available information in their judgments.
+This would motivate interventions to improve decision-making.
+For example, if we can identify how much complementary information AI predictions provide over human judgments, we can use this knowledge to guide model selection or motivate further data collection to improve the model.
+Conversely, finding evidence that model predictions contain decision-relevant information that humans do not exploit can motivate the design of explanations communicating complementary information.
+We contribute a decision-theoretic framework for characterizing the value of information available in an AI-assisted decision workflow.
 In our framework, information is considered valuable to a decision-maker to the extent that it is possible, in theory, to incorporate it into their decisions to improve performance.
 Specifically, our approach analyzes the expected marginal payoff gain from best case (Bayes rational) use of additional information over best case use of the information already encoded in agent decisions.
-While we focus on combinations of humans and AI agents, the methods we propose can be applied to any combination of agents.
+The rational framework allows us to upperbound the expected payoff that is achievable by any strategies in the same experiment, and identifies sub-optimality in agent use of information by comparing to rational behavior.
+The upper bound our framework estimates holds regardless of the human’s decision-making process, which may deviate from rationality, given a specified decision problem.
+Further, our methods can be used even when the decision problem definition is ambiguous, by using a robustness analysis over all possible proper scoring rules to identify the upper bound of performance in the worst case.
 We introduce two metrics for evaluating information value in human-AI collaboration.
@@
 The second—instance-level human-complementary information value—identifies opportunities for decision-makers to better use instance-level information such as specific AI predictions.
-We derive a new explanation technique (ILIV-SHAP) that reveals how data features influence the value of complementary information for an individual prediction.
-To evaluate these tools, we contribute the results of a crowdsourced between-subjects experiment in which humans make decisions with and without AI models with varying human-complementary information and with access to different explanation approaches.
+Applying the second metric, we derive a new explanation technique (ILIV-SHAP) that reveals how data features influence the value of complementary information for an individual prediction.
+To evaluate these tools, we contribute the results of a crowdsourced between-subjects experiment in which humans make decisions with and without AI models with varying human-complementary information and different explanation approaches.
 We find that an AI model with more human-complementary information leads to greater improvements in human-AI team performance over the human-alone baseline.
 We also find that adding an ILIV-SHAP explanation to a traditional SHAP explanation leads to greater improvements in human-AI performance over the human-alone baseline than only the SHAP explanation or no explanation.
-We demonstrate further uses of the framework for model selection, empirical evaluation of human-AI performance, and explanation design on existing datasets spanning medical diagnosis, human deepfake detection, and recidivism prediction (see Appendices E, D and F).
-We also include an observational study on the dataset from Vodrahalli et al.
-(2022) which identified the AI model with more human-complementary information helps human-AI teams achieve greater improvements in performance over the human-alone baseline in Appendix C.
+We demonstrate further uses of the framework in real-world decision-making tasks, including chest X-ray diagnosis (Rajpurkar et al., 2018; Johnson et al., 2019) and deepfake detection (Dolhansky et al., 2020; Groh et al., 2022) [^2] We also include an observational study on the dataset from Vodrahalli et al.
+(2022) which identified the AI model with more human-complementary information helps human-AI teams achieve greater improvements in performance over the human-alone baseline.
+See Appendix C..
 2 Related work
@@
 In response, a growing body of work seeks to evaluate and enhance complementarity in human–AI systems (Bansal et al., 2021b, 2019, a; Wilder et al., 2021; Rastogi et al., 2023; Mozannar et al., 2024b).
-The present work differs from much of this prior work by approaching human-AI complementarity from the perspective of information value and use, including asking whether the human and AI decisions provide additional information that is not used by the other.
+The present work differs from much of these prior works by approaching human-AI complementarity from the perspective of information value and use, including asking whether the human and AI decisions provide additional information that is not used by the other.
 Evaluation of human decision-making with machine learning.
@@
 Kleinberg et al.
-(2015) proposed that evaluations of human-AI collaboration should be based on the information that is available at the time of decisions.
+(2015) proposed that evaluations of human-AI collaboration should be based on the information that is available at the time of the decision.
 According to this view, our work defines Bayesian best-attainable-performance benchmarks similar to several prior works (Guo et al., 2024; Wu et al., 2023; Agrawal et al., 2020; Fudenberg et al., 2022).
@@
 Human information in machine learning.
-Some approaches focus on automating the decision pipeline by explicitly incorporating human expertise in developing machine learning models, such as by learning to defer (Mozannar et al., 2024a; Madras et al., 2018; Raghu et al., 2019; Keswani et al., 2022, 2021; Okati et al., 2021).
+Some approaches focus on automating the decision pipeline by explicitly incorporating human expertise in developing machine learning models, such as by learning to defer (Mozannar et al., 2024a; Madras et al., 2018; Raghu et al., 2019; Keswani et al., 2022, 2021; Okati et al., 2021; Chen et al., 2022).
 Corvelo Benz and Rodriguez (2023) propose multicalibration over human and AI model confidence information to guarantee the existence of an optimal monotonic decision rule.
@@
 (2023) propose a hypothesis testing framework to evaluate the added value of human expertise over AI forecasts.
-Our work shares the motivation of incorporating human expertise, but targets a slightly broader scope by quantifying the information value for all available signals and agent decisions in a specific scenario of human–AI decision tasks.
+Our work shares the motivation of incorporating human expertise, but provides a general analytical framework for quantifying the information value of all available signals and agent decisions in human–AI decision tasks.
 3 Framework
-Our framework takes as input a decision problem associated with an information model and outputs the value of information of any available signals to any agent, conditioning on the existing information in their decisions within a Bayesian decision theoretic framework.
-The intuition behind our use of statistical decision theory is that while human decision-makers may deviate from Bayesian rational decision-making, understanding the possible value of a signal requires defining the best attainable decision performance with that signal.
-Our framework provides two separate functions to quantify the value of information: one globally across the data-generating process, and one locally in a realization drawn from the data-generating process.
-We also introduce a robust analysis approach to information order, which enables us to compare the agent-complementary information in signals for all possible decision problems.
+Our framework takes as input a decision problem associated with an information model, including decisions from one or more agents.
+It outputs the value of information of available signals to the agents, conditioning on the existing information in their decisions.
+Understanding the possible value of a signal to a decision-maker requires defining the best attainable decision performance with that signal.
+We therefore adopt a Bayesian decision theoretic framework.
+Our framework provides two separate functions to quantify the value of information: one globally across the data-generating process, and one locally in a realization.
 Decision Problem.
@@
   • A payoff function $S:\mathbf{D}\times\mathbf{\Omega}\to\mathbb{R}$, used to assess the quality of a decision given a realization of the state.
-For example, $S(d=0,\omega=0)=0,S(d=0,\omega=1)=-100,S(d=1,\omega=0)=-50,S(d=1,\omega=1)=0$, which punishes the DM for selecting an action that does not match the weather.
-In decision problems corresponding to prediction tasks, the decision space is a probabilistic belief over the state space, i.e., $\mathbf{D}=\Delta(\mathbf{\Omega})$.
-For such problems, a payoff function is said to be a proper scoring rule if the optimal action is to predict the true distribution, i.e., $p=\arg\max_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim p}\![{S(d,\omega)}]$.
-For any decision problem with payoff function $S:\mathbf{D}\times\mathbf{\Omega}\to\mathbb{R}$, there is an equivalent proper scoring rule $\hat{S}:\Delta(\mathbf{\Omega})\times\mathbf{\Omega}\to\mathbb{R}$ defined by choosing the optimal decision under the reported belief.
-Formally,
-$\hat{S}(p,\omega)=S(\arg\max_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim p}\![{S(d,\omega)}],\omega).$ (1)
-Equation 1 shows a reduction from the payoff function $S$ to the proper scoring rule $\hat{S}$, i.e., any decision $p$ under $\hat{S}$ represents a decision $d$ in $S$ that best-responds to the distribution $p$.
-Therefore, the best-attainable performance defined in the proper scoring rule $\hat{S}$ is equivalent to the best-attainable performance defined in any payoff function $S$ that can be reduced to $\hat{S}$.
-Throughout our framework, we prefer proper scoring rules over non-proper scoring rules, since the best-attainable performance defined in the former implies the best-attainable performance in the latter.
+For example, $S(d=0,\omega=0)=0,S(d=0,\omega=1)=-100,S(d=1,\omega=0)=-50,S(d=1,\omega=1)=0$, which punishes the DM for wrongly taking or not taking the umbrella.
 Information Model.
-We cast the information available to a DM as a signal defined within an information model.
-We use the definition of an information model in Blackwell et al.
-(1951).
-The information model can be represented by a data-generating model with a set of signals.
+We cast the information available to a DM,including any available agent decisions, as a set of signals defined within an information model.
+Following the definition of an information model in Blackwell et al.
+(1951), the information model can be represented by a data-generating model with a set of signals.
   • Signals.
 There are $n$ “basic signals” represented as random variables $\Sigma_{1},\ldots,\Sigma_{n}$, from the signal spaces $\mathbf{\Sigma}_{1},\ldots,\mathbf{\Sigma}_{n}$.
-Basic signals represent information available to a decision-maker, e.g., $\mathbf{\Sigma}_{1}=\{\text{cloudy},\text{not cloudy}\}$, $\mathbf{\Sigma}_{2}\in\{0,\ldots,100\}$ for temperature in Celsius, etc.
+Basic signals represent information available to a decision-maker as they decide, e.g., $\mathbf{\Sigma}_{1}\in\Delta\mathbf{\Omega}$ for a probabilistic prediction of raining, $\mathbf{\Sigma}_{2}=\{\text{cloudy},\text{not cloudy}\}$, $\mathbf{\Sigma}_{3}\in\{0,\ldots,100\}$ for temperature in Celsius, etc.
 The decision-maker observes a signal, which is a subset of the basic signals, $V\subseteq 2^{\{\Sigma_{1},\dots,\Sigma_{n}\}}$.
-Specifically, we use $V=\{\Sigma_{j_{1}},\ldots,\Sigma_{j_{k}}\}$ for a signal having $k$ basic signals and denote the signal space as $\mathbf{V}=\mathbf{\Sigma}_{j_{1}}\times\ldots\times\mathbf{\Sigma}_{j_{k}}$.
-For example, a signal representing a combination of two basic signals $V=\{\Sigma_{1},\Sigma_{2}\}$ observed by the decision-maker might consist of cloudiness $\Sigma_{1}$ and the temperature $\Sigma_{2}$ of the day.
-Given a signal composed of $m$ basic signals, we write the realization of $V$ as $v=(\sigma_{j_{1}},\dots,\sigma_{j_{k}})$, where the realizations $\sigma_{j_{i}}\in\mathbf{\Sigma}_{j_{i}}$ are sorted by the index of the basic signals $j_{i}\in[n]$.
-The union $V$ of two signals $V_{1},V_{2}$ takes the set union, i.e., $V=V_{1}\cup V_{2}$.
+The combination of two signals $V_{1}$ and $V_{2}$ takes the set union $V=V_{1}\cup V_{2}$.
+For example, in the standard human-AI decision workflow where a human makes an independent judgment before consulting the AI, the set of all basic signals is $V=\{x,D^{\text{H}},D^{\text{AI}},e(D^{\text{AI}},x)\}$–the features of the instance, the human’s initial decisions, the AI’s predictions, and any explanation of the AI’s predictions.
   • Data-generating process.
-A data-generating process is a joint distribution $\pi\in\Delta(\mathbf{\Sigma}_{1}\times\ldots\times\mathbf{\Sigma}_{n}\times\mathbf{\Omega})$ over the basic signals and the payoff-relevant state.
+Decision benchmarks are defined relative to a specific data-generating process, a joint distribution $\pi\in\Delta(\mathbf{\Sigma}_{1}\times\ldots\times\mathbf{\Sigma}_{n}\times\mathbf{\Omega})$ over the basic signals and the payoff-relevant state.
 $\pi$ can be viewed as the combination of two distributions: the prior distribution of the state $\Pr[\omega]$ and the signal-generating distribution $\Pr[v|\omega]$ defining the conditional distribution of signals.
-The DM may only observe a subset $V$ of the $n$ basic signals.
-We can define the Bayesian posterior belief upon receiving a signal $V=v$ from the data-generating model as
+To represent the best-attainable performance of observing a subset $V$ of the $n$ basic signals, we use the Bayesian posterior belief upon receiving a signal $V=v$ as
     $\pi(\omega|v):=\Pr[\omega|v]=\frac{\pi(v,\omega)}{\pi(\omega)}$
-    where we slightly abuse notation to write $\pi(v,\omega)$ as the marginal probability of the signal realized to be $v$ and the state being $\omega$ with expectation over other signals.
-The choice of basic signals directly impacts how many observed samples are required to get a good estimate of the data-generating process.
-When high-dimensional signals such as images or text are used, it may not be computationally feasible to estimate the data-generating process.
-In such cases, one can preprocess the high-dimensional signals to get low-dimensional representations which (ideally) capture any important structure.
-These lower-dimensional signals can be defined by humans, such as when concepts are identified and then used to label the high-dimensional signals (e.g., images or parts of images).
-Alternatively, they can be defined algorithmically, by strategically applying dimensionality reduction to generate low-dimensional embeddings.
-We introduce an algorithm for identifying a set of optimal “coarsened” signals given a set of observed samples in Appendix B.
+    where $\pi(v,\omega)$ denotes the marginal probability of the signal realized to be $v$ and the state being $\omega$ with expectation over other signals, and $\pi(\omega)$ denotes the prior $\Pr[\omega]$.
+Throughout the paper, we use capital letters to denote a random variable of signal, e.g., $V$, and use little letetrs to denote a realization of signal, e.g., $v$.
 Information value.
@@
 Formally, given a decision task with payoff function $S$ and an information model $\pi$, the rational DM’s expected payoff given a (set of) signal(s) $V$ is
-$\mathrm{R}^{\pi,S}(V)=\text{\bf E}_{(v,\omega)\sim\pi}\![{S(d^{r}(v),\omega)}]$ (2)
+$\mathrm{R}(V):=\text{\bf E}_{(v,\omega)\sim\pi}\![{S(d^{r}(v),\omega)}]$ (1)
 where $d^{r}(\cdot):\mathbf{V}\rightarrow\mathbf{D}$ denotes the decision rule adopted by the rational DM.
 $d^{r}(v)=\arg\max_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\pi(\omega|v)}\![{S(d,\omega)}]$ (2)
-We use $\emptyset$ to represent a null signal, such that $\mathrm{R}^{\pi,S}(\emptyset)$ is the expected payoff of a Bayesian rational DM who only uses the prior distribution to make decisions.
-In this case, the Bayesian rational DM will take the best fixed action under the prior, and their expected payoff is:
-$\mathrm{R}^{\pi,S}(\emptyset)=\max_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\pi}\![{S(d,\omega)}]$ (4)
-This baseline defines the maximum expected payoff that can be achieved with no information.
-Bayesian decision theory quantifies the information value of $V$ as the payoff improvement of $V$ over the payoff obtained without information.
+We further characterize the maximum expected payoff that can be achieved with no information.
+This can be used as the baseline to quantify the information value of $V$ as the payoff improvement of $V$ over.
+We use $\emptyset$ to represent a null signal and $\mathrm{R}(\emptyset)$ represents the expected payoff of a Bayesian rational DM who only uses the prior distribution to make decisions.
+In this case, the Bayesian rational DM takes the best fixed action under the prior, and their expected payoff is:
+$\mathrm{R}(\emptyset):=\max_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\pi}\![{S(d,\omega)}]$ (3)
 Definition 3.1.
 Given a decision task with payoff function $S$ and an information model $\pi$, the information value of $V$ is defined as
-$\operatorname{IV}^{\pi,S}(V)=\mathrm{R}^{\pi,S}(V)-\mathrm{R}^{\pi,S}(\emptyset)$ (5)
-We adopt the same idea to define the agent-complementary information value in our framework.
+$\operatorname{IV}(V):=\mathrm{R}(V)-\mathrm{R}(\emptyset)$ (4)
+We adopt the same idea below to define the agent-complementary information value.
 3.1 Agent-Complementary Information Value
@@
 We recover the information value in agent decisions by offering the decisions as a signal to the Bayesian rational DM.
-We model the agent decisions as a random variable $D^{b}$ from the action space $\mathbf{D}$, which follows a joint distribution $\pi\in\Delta(\mathbf{\Sigma}_{1}\times\ldots\times\mathbf{\Sigma}_{n}\times\mathbf{\Omega}\times\mathbf{D})$ with the state and signals.
-The expected payoff of a Bayesian rational DM who knows $\pi$ is given by the function:
-$\mathrm{R}^{\pi,S}(D^{b})=\text{\bf E}_{(d^{b},\omega)\sim\pi}\![{S(d^{r}(d^{b}),\omega)}]$
-We seek to identify signals $V$ that can potentially improve agent decisions by analyzing the information value in the combined signal $D^{b}\cup V$ over the information value in $D^{b}$, which we define as the agent-complementary information value.
 Definition 3.2.
 Given a decision task with payoff function $S$ and an information model $\pi$, we define the agent-complementary information value ($\operatorname{ACIV}$) of $V$ on agent decisions $D^{b}$ as
-$\operatorname{ACIV}^{\pi,S}(V;D^{b})=\mathrm{R}^{\pi,S}(D^{b}\cup V)-\mathrm{R}^{\pi,S}(D^{b})$ (6)
+$\operatorname{ACIV}(V;D^{b}):=\mathrm{R}(D^{b}\cup V)-\mathrm{R}(D^{b})$ (5)
 If the $\operatorname{ACIV}$ of a signal $V$ is small relative to the baseline (3), this means either that the information value of $V$ to the decision problem is low (e.g., it is not correlated with $\omega$), or that the agent has already exploited the information in $V$ (e.g., the agent relies on $V$ or equivalent information to make their decisions such that their decisions correlate with $\omega$ in the same way as $V$ correlates with $\omega$).
@@
 Furthermore, $\operatorname{ACIV}$ can reveal complementary information between different types of agents.
-For instance, if we view AI predictions as $V$ and treat human decisions as the agent signal $D^{b}$, a large $\operatorname{ACIV}$ indicates that AI predictions add considerable value beyond what humans alone achieve.
+For instance, if we view AI predictions as $V$ and treat available human decisions as the agent signal $D^{b}$, a large $\operatorname{ACIV}$ indicates that AI predictions add considerable value beyond what humans alone achieve.
 In the reverse scenario, if human decisions serve as $V$ and AI predictions are $D^{b}$, we can measure how much humans can contribute over the information captured in the AI predictions.
-We demonstrate uses of $\operatorname{ACIV}$ in Appendix D and Appendix E.
+Of course, we are unlikely to observe identical realizations of the signals in continuous-valued or high-dimensional data, such as images or text.
+A natural relaxation is to consider the realizations that are sufficiently “similar” for the performance of the decision-maker.
+Algorithm 1 learns the posterior distribution of the payoff-relevant state from the signal and then uses that prediction as the probability distribution of the payoff-relevant state to choose the optimal action.
+This approach can be readily extended to the case where the signal is more complex, such as human decisions in the form of freeform text (e.g., radiology reports) or explanations of the AI predictions in the form of images (e.g., saliency map).
+Algorithm 1 A method to calculate the ACIV of $V$.
+1:Input: Observed realizations $\{v_{i},d^{b}_{i},\omega_{i}\}_{i=1}^{n}$, a regression algorithm $\mathcal{A}$, payoff function $S$, decision space $\mathbf{D}$, and state space $\mathbf{\Omega}$.
+2:Output: $\operatorname{ACIV}(V;D^{b})$
+[…]
+[…]
+5:for $i=1$ to $n$ do
+6:  $\hat{p}_{i}\leftarrow\hat{a}(v_{i},d^{b}_{i})$
+7:  $\hat{d}^{r}_{i}\leftarrow\operatorname{argmax}_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\hat{p}_{i}}\![{S(d,\omega)}]$
+8:  $\hat{p}^{b}_{i}\leftarrow\hat{a}^{b}(d^{b}_{i})$
+9:  $\hat{d}^{rb}_{i}\leftarrow\operatorname{argmax}_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\hat{p}^{b}_{i}}\![{S(d,\omega)}]$
+10:  $s_{i}\leftarrow S(\hat{d}^{r}_{i},\omega_{i})-S(\hat{d}^{rb}_{i},\omega_{i})$
+11:end for
+12:return $\frac{1}{n}\sum_{i=1}^{n}s_{i}$
 3.2 Instance-level Agent-complementary Information Value
-$\operatorname{ACIV}$ quantifies the value of the decision-relevant information in a signal $V$ across all possible realizations defined by the data-generating model.
+$\operatorname{ACIV}$ quantifies the value of the decision-relevant information in a signal $V$ across the distribution of all possible realizations defined by the data-generating model.
 To provide analogous instance-level quantification of information value, we define Instance-Level agent-complementary Information Value ($\operatorname{ILIV}$) to quantify the value of the decision-relevant information encoded by a single realization of the signal.
 This finer-grained view makes it possible to analyze how much an agent can benefit in theory from better incorporating instance-level information in their decision.
-Given a realization of signal $v=\{\sigma_{j_{1}},\ldots,\sigma_{j_{k}}\}$, we want to know the maximum expected payoff gain from gaining access to $v$ on the instances where $v$ is realized over the existing information encoded in agent decisions.
-Intuitively, this captures how much “room” there is for specific decisions to be improved through better use of the signal.
-To calculate instance-level information value, we rely on the performance of the Bayesian rational agent on the conditional distribution $\pi(\sigma_{1},\ldots,\sigma_{n},\omega,d^{b}|v)$ instead of the whole distribution $\pi(\sigma_{1},\ldots,\sigma_{n},\omega,d^{b})$.
+Given a realization of signal $v=\{\sigma_{j_{1}},\ldots,\sigma_{j_{k}}\}$, we want to know the maximum expected payoff gain from the access to $v$ on the instances where $v$ is realized over the existing information encoded in agent decisions.
+Intuitively, this captures how much “room” there is for a specific signal to be better used.
 Formally, given a decision task with payoff function $S$ and information model $\pi$, the expected payoff of the rational DM given signal $V=v^{\prime}$ on instances with the signal realization $V=v$ is
-$\mathrm{r}^{v,\pi,S}(v^{\prime})=\text{\bf E}_{\omega\sim\pi(\omega|v)}\![{S(d^{r}(v^{\prime}),\omega)}]$ (7)
-where $d^{r}(v)$ is the Bayesian optimal decision on receiving $v^{\prime}$ as defined in Equation 3.
+$\mathrm{r}^{v}(v^{\prime}):=\text{\bf E}_{\omega\sim\pi(\omega|v)}\![{S(d^{r}(v^{\prime}),\omega)}]$ (6)
+where $d^{r}(v^{\prime})$ is the Bayesian optimal decision on receiving $v^{\prime}$ as defined in Equation 2.
+Note that we allow $v^{\prime}\neq v$, i.e.
+the signal $v^{\prime}$ observed by the rational DM can be different from the actual realization $v$.
+The expected payoff of a rational DM who is misinformed is guaranteed to be lower than is correctly informed, i.e., $\mathrm{r}^{v}(v^{\prime})\leq\mathrm{r}^{v}(v)$ for any $v^{\prime}$.
+This notion of flexibility allows to consider the information value of a counterfactual realization of a signal, which we use in designing the explanation of ILIV-SHAP in the next section.
 If we consider the agent decisions in addition to the realization $v$, the rational DM’s expected payoff on instances where $V=v$ can be written as
-$\mathrm{r}^{v,\pi,S}(v^{\prime};D^{b})=\text{\bf E}_{(d^{b},\omega)\sim\pi(d^{b},\omega|v)}\![{S(d^{r}(v^{\prime}\cup d^{b}),\omega)}]$ (8)
-Note that we allow $v^{\prime}\neq v$, i.e.
-the signal $v^{\prime}$ observed by the rational DM being different from the actual realization $v$.
-When $v^{\prime}\neq v$, the rational DM is misinformed, thus $\mathrm{r}^{v,\pi,S}(v^{\prime})\leq\mathrm{r}^{v,\pi,S}(v)$.
-$\operatorname{ILIV}$ in Definition 3.3 quantifies this improvement in payoff when the rational DM observes the actual $v$.
+$\mathrm{r}^{v}(v^{\prime};D^{b}):=\text{\bf E}_{(d^{b},\omega)\sim\pi(d^{b},\omega|v)}\![{S(d^{r}(v^{\prime}\cup d^{b}),\omega)}]$ (7)
 Definition 3.3.
-Given a decision task with payoff function $S$ and an information model $\pi$, we define the instance-level agent-complementary information value ($\operatorname{ILIV}$) of signal $V=v^{\prime}$ on instances where $V=v$ as:
-$\operatorname{ILIV}^{v,\pi,S}(v^{\prime};D^{b})=\mathrm{r}^{v,\pi,S}(v^{\prime};D^{b})-\mathrm{r}^{v,\pi,S}(\emptyset;D^{b}).$ (9)
-$\mathrm{r}^{v,\pi,S}(\emptyset;D^{b})$ represents the expected rational payoff on instances with $V=v$, where the rational DM only knows the information encoded in agent decisions $D^{b}$ without knowing any realizations of $V$.
-$\operatorname{ILIV}$ maximizes when the signal does not misinform, i.e., $\operatorname{ILIV}^{v,\pi,S}(v;D^{b})\geq\operatorname{ILIV}^{v,\pi,S}(v^{\prime};D^{b}),\text{for }v^{\prime}\in\mathbf{V}$.
-3.2.1 ILIV-SHAP Information-based Explanation
-We apply $\operatorname{ILIV}$ to define an information-based explanation technique (ILIV-SHAP) that extends the canonical SHAP feature saliency explanation of a model’s prediction.
-ILIV-SHAP communicates how the data features lead to changes in the information value of AI predictions.
-While traditional SHAP summarizes the average contribution of each feature to a specific prediction over the baseline (prior) prediction, ILIV-SHAP summarizes the average contribution of each feature to the decision-relevant information value contained in the prediction.
-Suppose a model $f$ that takes as input $m$ features and outputs a real number.
+Given a decision task with payoff function $S$ and an information model $\pi$, we define the instance-level agent-complementary information value ($\operatorname{ILIV}$) of signal realization $V=v^{\prime}$ on instances where $V=v$ as:
+$\operatorname{ILIV}^{v}(v^{\prime};D^{b}):=\mathrm{r}^{v}(v^{\prime};D^{b})-\mathrm{r}^{v}(\emptyset;D^{b}).$ (8)
+where $\mathrm{r}^{v}(\emptyset;D^{b})$ represents the expected rational payoff on instances of $V=v$, where the rational DM only knows the information encoded in agent decisions without knowing any realizations of $V$.
+$\operatorname{ILIV}$ maximizes when the signal does not misinform, i.e., $\operatorname{ILIV}^{v}(v;D^{b})\geq\operatorname{ILIV}^{v}(v^{\prime};D^{b}),\text{for }v^{\prime}\in\mathbf{V}$.
+Algorithm 2 A method to calculate the ILIV of $v^{\prime}$ on the instances of $v$.
+1:Input: Observed realizations $\{v_{i},d^{b}_{i},\omega_{i}\}_{i=1}^{n}$, test counterfactual signal realization $v^{\prime}$, test signal realization $v$, a regression algorithm $\mathcal{A}$, payoff function $S$, decision space $\mathbf{D}$, and state space $\mathbf{\Omega}$.
+2:Output: $\operatorname{ILIV}^{v}(v^{\prime};D^{b})$
+[…]
+[…]
+[…]
+6:for $i=1$ to $k$ do
+7:  $\hat{p}_{i}\leftarrow\hat{a}(v^{\prime},d^{b}_{j_{i}})$
+8:  $\hat{d}^{r}_{i}\leftarrow\operatorname{argmax}_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\hat{p}_{i}}\![{S(d,\omega)}]$
+9:  $\hat{p}^{b}_{i}\leftarrow\hat{a}^{b}(d^{b}_{j_{i}})$
+10:  $\hat{d}^{rb}_{i}\leftarrow\operatorname{argmax}_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\hat{p}^{b}_{i}}\![{S(d,\omega)}]$
+11:  $s_{i}\leftarrow S(\hat{d}^{r}_{i},\omega_{i})-S(\hat{d}^{rb}_{i},\omega_{i})$
+12:end for
+13:return $\frac{1}{k}\sum_{i=1}^{k}s_{i}$
+4 Information-based Explanation
+We define an information-based explanation (ILIV-SHAP) to communicate where the AI prediction offers complementary information over the agent decisions.
+Traditional saliency-based explanations communicate the average contribution of each feature to a prediction over the baseline (prior) prediction, while ILIV-SHAP communicates the average contribution of each feature to the agent-complementary information value contained in the prediction.
+Specifically, suppose a model $f$ that, for example, takes as input $m$ features and outputs a real number.
 Given an instance $\mathbf{x}=(x_{1},\ldots,x_{m})$, the importance of one feature $x_{i}$ to the model output $f(\mathbf{x})$ is encoded by the expected difference of model outputs when $x_{i}$ is marginalized out.
-Specifically, this is quantified by $f(\mathbf{x})-\text{\bf E}\![{f(X)|X_{-i}=\mathbf{x}_{-i}}]$, where $X_{-i}$ denotes all features except $X_{i}$.
+This is quantified by $f(\mathbf{x})-\text{\bf E}\![{f(X)|X_{-i}=\mathbf{x}_{-i}}]$, where $X_{-i}$ denotes all features except $X_{i}$.
 Considering the interaction between features, SHAP (Lundberg and Lee, 2017) uses the Shapley value to quantify the importance scores averaged on different combinations of features:
-$\phi_{i}(f,\mathbf{x})=\sum_{\mathbf{x}^{\prime}\subseteq\mathbf{x}}\frac{|\mathbf{x}^{\prime}|!(m-|\mathbf{x}^{\prime}|-1)!}{m!}[g_{f}(\mathbf{x}^{\prime})-g_{f}(\mathbf{x}^{\prime}\backslash x_{i})]$
-where $g_{f}(\mathbf{x}^{\prime})$ denotes the expected model output conditioned on $\mathbf{x}^{\prime}$, i.e., $\text{\bf E}\![{f(X)|X^{\prime}=\mathbf{x}^{\prime}}]$ for any $\mathbf{x}^{\prime}\subseteq\mathbf{x}$.
-The scores $\phi_{i}(f,\mathbf{x})$ output by SHAP construct an explanation model for a model output, which quantifies the expected counterfactual change in the model output caused by the feature $x_{i}$.
-ILIV-SHAP extends SHAP to give an explanation model of how features impact the decision-relevant information value of an individual model output.
+[…]
+where $g_{f}(\mathbf{x}^{\prime})$ denotes the expected output conditioned on $\mathbf{x}^{\prime}$, i.e., $\text{\bf E}\![{f(X)|X^{\prime}=\mathbf{x}^{\prime}}]$ for any $\mathbf{x}^{\prime}\subseteq\mathbf{x}$ where $\mathbf{x}^{\prime}$ is the features that are not marginalized out.
 Definition 4.1 (ILIV-SHAP).
 Given a model $f$ and data features $\mathbf{x}=(x_{1},\ldots,x_{m})$, the importance score of the i-th feature by ILIV-SHAP is
-$\phi_{i}^{\operatorname{ILIV}}(f,\mathbf{x})=\sum_{\mathbf{x}^{\prime}\subseteq\mathbf{x}}\frac{|\mathbf{x}^{\prime}|!(m-|\mathbf{x}^{\prime}|-1)!}{m!}[\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(g_{f}(\mathbf{x}^{\prime});D^{b})-\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(g_{f}(\mathbf{x}^{\prime}\backslash x_{i});D^{b})]$
-where $\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(g_{f}(\mathbf{x}^{\prime});D^{b})$ denotes a counterfactual evaluation of $\operatorname{ILIV}$, which quantifies the expected payoff gain from additionally knowing $g_{f}(\mathbf{x}^{\prime})$ on the instances where $f(\mathbf{x})$ is realized.
-This counterfactual version of $\operatorname{ILIV}$ is guaranteed to achieve the maximum when $g_{f}(\mathbf{x}^{\prime})$ informs $f(\mathbf{x})$, i.e., $f(\mathbf{x})=g_{f}(\mathbf{x}^{\prime})$ – the features missing from $\mathbf{x}^{\prime}$ have no impact on $f(\mathbf{x})$.
-The explanation model offered by ILIV-SHAP is grounded in the following two properties.
-First, $\phi^{\operatorname{ILIV}}_{i}(f,\mathbf{x})$ are consistent with the extent to which feature $x_{i}$ contributes to the information value in $f(\mathbf{x})$.
-Specifically, for any two models $f$ and $f^{\prime}$ and any $\mathbf{x}^{\prime}\subseteq\mathbf{x}$, if $x_{i}$ contributes more to the information value in $f^{\prime}$ than the information value in $f$, i.e., $\operatorname{ILIV}^{f^{\prime}(\mathbf{x}),\pi,S}(g_{f^{\prime}}(\mathbf{x}^{\prime});D^{b})-\operatorname{ILIV}^{f^{\prime}(\mathbf{x}),\pi,S}(g_{f^{\prime}}(\mathbf{x}^{\prime}\backslash x_{i});D^{b})\geq\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(g_{f}(\mathbf{x}^{\prime});D^{b})-\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(g_{f}(\mathbf{x}^{\prime}\backslash x_{i});D^{b})$, then $\phi^{\operatorname{ILIV}}_{i}(f^{\prime},\mathbf{x})\geq\phi^{\operatorname{ILIV}}_{i}(f,\mathbf{x})$.
-This property of ILIV-SHAP follows the consistency property of SHAP (Lundberg and Lee, 2017).
-Second, summing up the scores $\phi^{\operatorname{ILIV}}_{i}(f,\mathbf{x})$ recovers the information value of model output, $\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(f(\mathbf{x});D^{b})$.
-Formally, for any $f$ and any $\mathbf{x}$, $\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(f(\mathbf{x});D^{b})=\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(\text{\bf E}\![{f(X)}];D^{b})+\sum_{i=1}^{m}\phi^{\operatorname{ILIV}}_{i}(f,\mathbf{x})$.
-We demonstrate use of ILIV-SHAP in Appendix F.
-3.3 Robust Analysis of Information Order
-Our approach assumes a decision problem as input and evaluates agents’ decisions and use of information on this problem.
-However, evaluators may face ambiguity around the appropriate decision problem specification, and in particular, the appropriate scoring rule.
-In particular, ambiguity can arise in payoff functions; doctors, for example, penalize false negative results differently when diagnosing younger versus older patients (Mclaughlin and Spiess, 2023).
-Blackwell’s comparison of signals (Blackwell et al., 1951) is an ideal tool for addressing ambiguity about the payoff function, as it defines a signal $V_{1}$ as more informative than $V_{2}$ if $V_{1}$ has a higher information value on all possible decision problems.
-We identify this partial order by decomposing the space of decision problems via a basis of proper scoring rules [^2] For rational DMs, any decision problem can be represented by an equivalent proper scoring rule in Equation 1, such that the partial order defined via proper scoring rules also applies to the corresponding decisin tasks.
-(Li et al., 2022; Kleinberg et al., 2023).
-Definition 3.5 (Blackwell Order of Information).
-A signal $V_{1}$ is Blackwell more informative than $V_{2}$ if $V_{1}$ achieves a higher best-attainable payoff on any decision problems:
-$\mathrm{R}^{\pi,S}(V_{1})\geq\mathrm{R}^{\pi,S}(V_{2}),\forall S$
-The Blackwell order is evaluated over all possible decision problems, which cannot be tested directly.
-Fortunately, we only need to test over all proper scoring rules since any decision problem can be represented by an equivalent proper scoring rule, and the space of proper scoring rules can be characterized by a set of V-shaped scoring rules.
-A V-shaped scoring rule is parameterized by the kink of the piecewise-linear utility function.
-Definition 3.6.
-(V-shaped scoring rule) A V-shaped scoring rule with kink $\mu\in(0,\frac{1}{2}]$ is defined as
-$S_{\mu}(d,\omega)=\left\{\begin{array}[]{cc}\frac{1}{2}-\frac{1}{2}\cdot\frac{\omega-\mu}{1-\mu}&\text{if }d\leq\mu\\ \frac{1}{2}+\frac{1}{2}\cdot\frac{\omega-\mu}{1-\mu}&\text{else},\end{array}\right.$
-When $\mu^{\prime}\in(\frac{1}{2},1)$, the V-shaped scoring rule can be symmetrically defined by $S_{\mu^{\prime}}=S_{1-\mu^{\prime}}(1-y,\omega)$.
-Intuitively, the kink $\mu$ represents the threshold belief where the decision-maker switches between two actions.
-Larger $\mu$ means that the decision-makers will prefer $d=1$ more.
-The closer $\mu$ is to $0.5$, the more indifferent the decision-maker is to $d=0$ or $d=1$.
-Proposition 3.7 shows that if $V_{1}$ achieves a higher information value on the basis of V-shaped proper scoring rules than $V_{2}$, then $V_{1}$ is Blackwell more informative than $V_{2}$.
-Proposition 3.7 follows from the fact that any best-responding payoff can be linearly decomposed into the payoff on V-shaped scoring rules.
-Proposition 3.7 (Hu and Wu 2024).
-If $\forall\mu\in(0,1)$
-$\mathrm{R}^{\pi,S_{\mu}}(V_{1})\geq\mathrm{R}^{\pi,S_{\mu}}(V_{2}),$
-then $V_{1}$ is Blackwell more informative than $V_{2}$.
-Extending this to ACIV, $V_{1}$ offers a higher complementary value than $V_{2}$ under the Blackwell order if
-$\operatorname{ACIV}^{\pi,S_{\mu}}(V_{1};D^{b})\geq\operatorname{ACIV}^{\pi,S_{\mu}}(V_{2};D^{b}),\forall\mu\in(0,1)$
-This definition allows us to rank signals (or sets of signals) without needing to commit to a specific payoff function. We present a use case in Appendix D.
-4 Demonstrations
-We demonstrate uses of our framework on three real-world decision-making tasks: medical diagnosis, deepfake video detection, and recidivism prediction.
-First, we demonstrate the framework’s utility in model selection by evaluating how well different cardiac dysfunction prediction models trained on chest X-rays complement radiologists’ predictions (Rajpurkar et al., 2018; Johnson et al., 2019), showing how even among models with similar accuracy,some models strictly offer more complementary information than others over the radiologists.
-Next, we show how the framework can be used to evaluate how well humans exploit available features in a deepfake detection task (Dolhansky et al., 2020; Groh et al., 2022), how this compares to how an AI models exploits the information, and how providing AI assistance affects the human use.
-Suprisingly, even though AI assistance helps human-AI teams achieve higher decision accuracy, the $\operatorname{ACIV}$s of the video-level features did not change much compared to the $\operatorname{ACIV}$s over human-only predictions, indicating that simply displaying AI’s prediction did not help humans exploit the available information.
-Lastly, we demonstrate how ILIV-SHAP explanations surface unique insight into information value that humans lack relative to conventional SHAP for a recidivism prediction task (Angwin et al., 2022; Dressel and Farid, 2018).
-Detailed descriptions and results of all demonstrations are in Appendices E, D and F.
+[…]
+where $\operatorname{ILIV}^{f(\mathbf{x})}(g_{f}(\mathbf{x}^{\prime});D^{b})$ denotes a counterfactual evaluation of $\operatorname{ILIV}$, which quantifies the expected payoff gain from additionally knowing $g_{f}(\mathbf{x}^{\prime})$ over $D^{b}$ on the instances where the actual prediction is $f(\mathbf{x})$.
+ILIV-SHAP shares the same properties as SHAP (Lundberg and Lee, 2017) by similarly constructing the importance scores.
+For example, ILIV-SHAP satisfies the critical efficiency axiom, i.e., the sum of the importance scores equals the information value of the model output, and the symmetry axiom, i.e., the importance scores are the same for any two features that contribute the same amount to the information value of the model output.
 5 Experiment
@@
   2. 2.
-    Does ILIV-SHAP improved human-AI decision-making over SHAP alone?
-Our experiment assigns participants to one of two AI models with varying ex-ante ACIV and one of three explanation conditions: ILIV-SHAP and SHAP, SHAP, and no explanation.
-We hypothesize that the higher ACIV model will lead to better human-AI decisions and presenting ILIV-SHAP explanations for this model will lead to better human-AI decisions than the other explanation types.
-Metric AI1 AI2
+    Does ILIV-SHAP improve human-AI decision-making over SHAP alone?
+Our experiment assigns participants to one of two AI models with varying ex-ante ACIV {AI1=high ACIV, AI2=low ACIV}, and one of three explanation conditions {ILIV-SHAP and SHAP, SHAP, No Explanation}.
+We hypothesize that the higher ACIV model will lead to better human-AI decisions, and that presenting ILIV-SHAP explanations for this model will lead to better human-AI decisions than the other explanation types.
+Metric AI1 (higher ACIV) AI2 (lower ACIV)
 MAPE 14.30% 14.51%
-R2 0.81 0.81
+R squared 0.81 0.81
 $\operatorname{ACIV}$
@@
 4.61% 2.00%
-Figure 1: Left: AI performance and ACIV. Right: Posterior predicted reduction in the human-AI team’s absolute percentage error (APE) with 95% confidence intervals.
+Figure 1: Left: AI performance and ACIV.
+Right: Reduction in the human-AI team’s absolute percentage error (APE) with 95% confidence intervals according to posterior predictions of our regression model.
 Task and data.
-We study an AI-assisted house price prediction, following other recent crowdsourced studies (e.g., Chiang and Yin (2021); Hemmer et al.
+We study AI-assisted house price prediction, following other recent crowdsourced studies (e.g., Chiang and Yin (2021); Hemmer et al.
 (2022); Poursabzi-Sangdeh et al.
@@
 We recruited 421 US-based participants via the crowdsourcing platform Prolific.
-Each participant was randomly assigned to 12 houses in the Ames Iowa Housing Dataset and performed a sequence of 24 trials (where two decisions were elicited for each house, one without AI assistance and one with), split evenly across a human-alone prediction phase and a human-AI prediction phase.
+Each participant was randomly assigned to 12 houses out of 733 in the Ames Iowa Housing Dataset.
+Each completed a sequence of 24 trials, where two decisions were elicited for each house: one without AI assistance and one with.
 In the first 12 trials, participants were asked to predict the house price of the assigned house using six features of the house, including year built, the living area above ground in square feet, the size of garage in car capacity, the number of fireplaces, year remodeled, and ratings of the house’s material and finish.
-We intentionally reduced the interpretability of the year remodeled and material and finish by witholding what variables they represented and rescaling their values.
+We intentionally reduced the interpretability of the year remodeled and material and finish by relabeling them as “Feature X” and “Feature Y” and rescaling their values.
+Therefore, the AI model that takes these two features as input is expected to have complementary information over human participants.
 In the second 12 trials, participants were asked to revise their guesses in the first round after seeing the AI’s prediction and explanation.
@@
 AI models and explanations.
-Participants were assigned to one of six experimental conditions resulting from crossing the AI models (2 levels, where AI1 has high $\operatorname{ACIV}$ and AI2 has low $\operatorname{ACIV}$) with the explanation type (3 levels: ILIV-SHAP + SHAP, SHAP, and no explanation).
-We designed AI1 and AI2 so as to vary their potential to complement humans: AI1 was trained with the six features and the true house price, with noise added to the six features.
-On the other hand, AI2 was trained with the four human-interpretable features and the true house price, such that there would be less potential for the predictions to add unique information over human judgments.
-We used the same training data for both AI models and ensured that the two AI models achieved similar performance (Figure 1), so that model accuracy differences would not confound our results.
+Participants were assigned to one of six experimental conditions resulting from crossing the two AI models with the three explanation types.
+We designed AI1 and AI2 to have varying potential to complement humans: AI1 was trained with the six features and the true house price, with noise added to the six features (which is meant to be bring down the accuracy of AI1 to be comparable with AI2).
+On the other hand, AI2 was trained with the four human-interpretable features and the true house price, such that there would be less potential for the predictions to add complementary information over human judgments.
+We used the same training data for both AI models and ensured that the two AI models achieved similar performance (Figure 1 left), so that model accuracy differences would not confound our results.
 In the SHAP explanation, we used the SHAP values of the six features to explain the AI’s prediction.
-To generate the ILIV-SHAP explanations, we used independent human decisions we collected on Prolific for the dataset using the same set-up prior to running the official study.
-We present ILIV-SHAP explanations of the six features that ranked and highlighted the features where the human-complementary information the AI’s prediction gets from the feature is higher than a threshold (see Appendix I for explanation examples).
+To generate the ILIV-SHAP explanations, we used independent human decisions we collected on Prolific for the dataset using the same task and features prior to running the official study.
+We present ILIV-SHAP explanations of the six features, which ranked and highlighted the features where the human-complementary information (ILIV) over the AI’s prediction exceeds a threshold (see Appendix H for explanation examples).
 Evaluation metrics.
-We report the $\operatorname{ACIV}$ of the AI models in the mean squared error (MSE) [^3] We included MSE since it is a proper scoring rule while MAPE is not.
-and the mean absolute percentage error (MAPE).
-To measure the human unassisted and human-AI team performance, we fit a preregistered Bayesian hierarchical regression model [^4] https://aspredicted.org/jfm6-rrs3.pdf to the percentage error (PE) between the human’s prediction and the true house price, $\textit{PE}=(d-\omega)/\omega$ using R’s brms package (Bürkner, 2017).
-[…]
-[…]
-[…]
+We report the $\operatorname{ACIV}$ of the AI models in mean squared error (MSE) [^3] We included MSE since it is a proper scoring rule while MAPE is not.
+and mean absolute percentage error (MAPE).
+To measure the human alone and human-AI team performance, we fit a preregistered Bayesian hierarchical regression model with weakly informed priors444link to preregistration redacted for peer review to the percentage error (PE) between the human’s prediction and the true house price, $\textit{PE}=(d-\omega)/\omega$, using R’s brms package (Bürkner, 2017).
+PE $\sim\text{student\_t}(\mu,\sigma,\nu)$
+$\mu$ $=\textit{AI}*\textit{explanation}+\textit{round}+(1|\textit{participant\_id})$
+$\log(\sigma)$ $=\textit{AI}*\textit{explanation}+\textit{round}+(1|\textit{participant\_id})$
 $\nu$ $\sim\text{Gamma}(2,0.1)$
-[…]
+where AI and explanation are indicators of the experimental conditions, round is an indicator of whether the trial is in the first or second round, and participant_id is a unique identifier for each participant.
+We evaluate human-AI team performance by the reduction in absolute percentage error (APE) over human-alone predictions in the first round, i.e., $\text{Reduction in APE}=\text{\bf E}\![{\lvert\textit{PE}\rvert\mid\textit{round}=1}]-\text{\bf E}\![{\lvert\textit{PE}\rvert\mid\textit{round}=2}]$, where the expectation is taken over the Bayesian posterior distribution of the above Bayesian model.
 Results.
@@
 Note the similar predictive performance of the two AI models.
-Posterior predictions of human-AI team performance according to our regression model also align with our expectations based on the complementary information each model provides (Figure 1 far right).
-Collapsing across explanation conditions, AI1 reduces APE by more than AI2 ($6.27\%[5.91\%,6.63\%]$ vs $5.65\%[5.26\%,6.05\%]$ respectively).
-Access to the ILIV-SHAP + SHAP explanation results in a greater reduction in APE for AI1 than AI2 ($7.11\%[6.57\%,7.65\%]$ vs $5.83\%[5.17\%,6.49\%]$ respectively).
-Figure 1 also illustrates the effect of the different explanation conditions on the human-AI team performance.
-When the AI has sufficient complementing information (i.e., AI1), the ILIV-SHAP + SHAP explanation more effectively reduces APE for the human-AI decisions than the SHAP explanation and baseline (no explanation) ($7.11\%[6.57\%,7.65\%]$ for ILIV-SHAP, $5.20\%[4.53\%,5.87\%]$ for SHAP, and $6.50\%[5.86\%,7.14\%]$ for no explanation, all for model AI1).
+The posterior predictions of human-AI team performance from our regression model also align with expectations based on the complementary information each model provides (Figure 1 far right).
+Collapsing across explanation conditions, AI1 (higher ACIV) reduces APE by more than AI2 ($6.27\%$$5.91\%,$$6.63\%]$ vs $5.65\%[5.26\%,6.05\%]$ respectively).
+Access to the ILIV-SHAP + SHAP explanation results in a greater reduction in APE for AI1 than AI2 ($7.11\%[6.57\%,7.65\%]$ vs $5.83\%[5.17\%,6.49\%]$ respectively) [^5] See Appendix D for the significant test results.
+Figure 1 also illustrates the effect of the different explanation conditions on human-AI team performance.
+When the AI has sufficient complementing information (AI1), the ILIV-SHAP + SHAP explanation more effectively reduces APE for the human-AI decisions than the SHAP explanation and baseline (no explanation) ($7.11\%[6.57\%,7.65\%]$ for ILIV-SHAP, $5.20\%[4.53\%,5.87\%]$ for SHAP, and $6.50\%[5.86\%,7.14\%]$ for no explanation).
+6 Demonstrations
+6.1 Chest X-ray Diagnosis
+We study a chest X-ray diagnosis task.
+As agent decisions, $D^{\text{AI}}$ and $D^{\text{H}}$, we consider five predictive models fine-tuned on the MIMIC-CXR database (Johnson et al., 2019) and radiologists’ textual reports recorded in MIMIC-CXR.
+We use five pretrained image models with the same choices in Irvin et al.
+(2019).
+We train the models on 12,228 radiographs, and validated on 6,115 randomly sampled radiographs.
+For the payoff-relevant state, $\omega\in\mathbf{\Omega}=\{0,1\}$, we used results from two types of blood tests (‘NT-proBNP’ and ‘troponin’ cut by age-specific thresholds).
+The decision task is formalized as a prediction problem with $\mathbf{D}=\Delta\mathbf{\Omega}$ and $S(d,\omega)=1-(d-\omega)^{2}$.
+Can the AI models complement human judgment?
+Figure 2(a) shows all the models offer complementary information value to the radiologists’ reports (green distributions improve over the purple distribution), and in the other direction, the radiologists’ reports offer complementary information value to the models (green distributions improve over the orange distributions).
+Which AI model offers the most decision-relevant information over human judgments?
+Figure 2(a) shows that VisionTransformer contains slightly higher information value than the other models, and Inception v3 contains slightly lower information value than the other models.
+We further assess the robustness of VisionTransformer’s superiority over the other AI models across many possible payoff functions to test if there is a Blackwell ordering of models in Appendix F and Figure 6.
+(a) Chest X-ray diagnosis results.
+(b) Deepfake detection results
+Figure 2: Demonstration results for both tasks.
+Distributions show information values.
+For a), we plot ACIV of radiologist decisions, different choices of AI models and signals that combine AI predictions and radiologist decisions.
+For b), we plot the information value of the combination of video-level features and agent decisions (including human decisions, AI predictions, and human-AI teams’ decisions).
+These two figures demonstrate that our framework can be used to a) reveal the complementary information between AI predictions and human decisions and b) analyze how humans use information in data features differently from AI.
+As shown in this demonstration, doctors may use our framework to learn how much complementary information value the AI models offer over their decisions, and which model offers the most.
+[^6] For space constraints, we refer to the full descriptions and results in Appendices F and E.
+6.2 Deepfake Detection
+We study a deepfake detection task (Dolhansky et al., 2020; Groh et al., 2022).
+We select the model in the Deepfake Detection Challenge (Zhang et al., 2016), with estimated $65\%$ accuracy on holdout data.
+We use the human decisions and the human-AI team’s decisions from Groh et al.
+(2022), who collected judgments on the videos from n=5,524 participants recruited on Prolific.
+We use the Brier score as the payoff function, with the binary payoff-related state: $\omega\in\{0,1\}=\{\text{genuine},\text{fake}\}$.
+This choice differs from Groh et al.
+(2022)’s use of mean absolute error, but again we prefer the quadratic score because it is a proper scoring rule where truthfully reporting beliefs maximizes the score.
+Groh et al.
+(2022) manually label seven video-level features, which we use as binary indicators in place of Algorithm 1 in light of the high dimensional signals: graininess, blurriness, darkness, presence of a flickering face, presence of two people, presence of a floating distraction, and the presence of an individual with dark skin.
+We show the ACIVs of the seven features over the decisions, the AI predictions, and the human-AI decisions in Figure 2(b).
+How much decision-relevant information do agents’ decisions offer?
+We first compare the information value of the AI predictions to that of the human decisions.
+Figure 2(b) shows that AI predictions provide about $65\%$ of the total possible information value over the no-information baseline, while human decisions only provide about $15\%$.
+We next consider the human-AI decisions.
+Given that the AI predictions contain a significant portion of the total possible information value, we might hope that when participants have access to the AI predictions, their performance will be close to the full information baseline.
+However, the information value of the human-AI decisions only achieves a small proportion of the total possible information value ($30\%$).
+How much additional decision-relevant information do the available features offer over agents’ decisions?
+To understand what information might improve human decisions, we assess the $\operatorname{ACIV}$s of different video-level features over different agents.
+As shown on the fifth row in Figure 2(b), the presence of a flickering face offers larger $\operatorname{ACIV}$ over human decisions than over AI predictions, meaning that human decisions could improve by a greater amount if they were to incorporate this information.
+Meanwhile, as shown on the fourth row in Figure 2(b), the presence of an individual with dark skin offers larger $\operatorname{ACIV}$ over AI predictions than over human decisions, suggesting that humans make greater use of this information.
+This suggests that the AI and human rely on differing information to make their initial predictions.
 7 Discussion and Limitations
 Our decision-theoretic framework quantifies the information value of signals available in a human-AI decision setting over the information value of agent decisions.
-The basis of our framework in Bayesian decision theory means that it provides a theoretical basis that can be adapted to support comparisons to human behavior to drive learning (Hullman and Gelman, 2021), even if human agents deviate considerably from Bayesian rational decision-making.
+Importantly, the basis of our framework in Bayesian decision theory does not require that actual (e.g., human) decision-makers achieve Bayesian rational decision-making.
+Rather, it provides a theoretical basis to support comparisons to human behavior to drive learning (see, e.g., (Guo et al., 2024; Hullman and Gelman, 2021; Wu et al., 2023).
+This theoretical basis is necessary to establish well-defined benchmarks.
 We experimentally demonstrate the power of the framework for analyzing and improving human-AI decision-making.
 Our proposed ILIV-SHAP explanation improves performance over an existing state-of-the-art explanation strategy.
+This suggests that valuing information value in terms of what it says about the payoff relevant state, not just the AI prediction, can improve the design of signals for human-AI decision-making.
+Hence our work offers theoretical support for attempts to design new explanations (e.g., Li et al.
+(2025)).
 Our work provides a promising methodological framework for an emerging research agenda around optimally combining agents’ information for decisions (e.g., Alur et al.
@@
 While we focused on human-AI decisions, the framework an be applied to any combination of AI or human agent judgments.
-One direction for follow-up work is to develop tests that complement the framework by helping evaluators confirm the information that agents are using.
-Note that low ACIV of a signal over human decisions is not necessarily sufficient for concluding that particular signals were used by the human.
-It is often unknown whether a human has other private signals offering equivalent information to a feature being analyzed.
+Though most of our definitions and analysis are focused on decision tasks with single well-defined payoff function, the framework can be readily extended to more complex decision tasks with an ambiguous set of payoff functions or totally unidentifiable payoff functions.
+We present a robustness analysis framework in Appendix B to calculate the Blackwell ordering of models over proper scoring rules.
+A demonstration of the robustness analysis is shown in Appendix F.
 Appendix A The Combinatorial Nature of the Value of Signals
@@
 Algorithm 3 Greedy algorithm for marginal gain of $\operatorname{ACIV}$
-1:  $V^{*}=\{D^{b}\}$
-2:  $\Phi^{*}=\{\}$
+[…]
+[…]
 3:for $i=1$ to $n$ do
@@
 9:end for
-10:  $output\ \phi_{j^{*}}$
-Appendix B Estimating the Data-Generating Process from Observations
-Suppose the data-generating process is estimated from the empirical distribution of the observed realizations of the payoff state value and the signal, $D=\{(\omega_{i},v_{i})\}_{i=1}^{T}$, i.e.,
-$\hat{\pi}_{D}(\omega,v)=\frac{1}{T}\sum_{i=1}^{T}\mathbb{1}_{\omega=\omega_{i}}\cdot\mathbb{1}_{v=v_{i}}$ (11)
-In applying our framework, the rational DM’s decision rule $d^{r}$ should not overfit to the observed dataset.
-This threatens the usefulness of analysis using the DGP.
-To avoid this situation, we can either collect more observations or “coarsen” the signal space, where “coarsen” means reducing the signal space by aggregating similar signals into a single signal.
-We propose a systematic procedure for the latter since it may be too costly or otherwise not possible to collect more observations.
-Our objective is to find the coarsened signal structure that has the largest signal space and avoids the rational decision rule from overlapping.
-Denote the maximum tolerance factor for overfitting as $\epsilon$.
-The rational DM’s decision rule $d^{r}_{\hat{\pi}_{D^{\prime}}}$ that is estimated from a set of observations $D^{\prime}$ should perform similarly on $D^{\prime}$ and on the evaluation of a holdout set $D\setminus D^{\prime}$, i.e.,
-$\text{\bf E}_{(\omega,v)\sim\hat{\pi}_{D^{\prime}}}\![{S(d^{r}_{\hat{\pi}_{D^{\prime}}}(v),\omega)}]-\text{\bf E}_{(\omega,v)\sim\hat{\pi}_{D\setminus D^{\prime}}}\![{S(d^{r}_{\hat{\pi}_{D^{\prime}}}(v),\omega)}]\leq\epsilon*\text{\bf E}_{(\omega,v)\sim\hat{\pi}_{D^{\prime}}}\![{S(d^{r}_{\hat{\pi}_{D^{\prime}}}(v),\omega)}]$ (12)
-Algorithm 2 Estimating the data-generating process from a set of observations
-1:  Input: Observed dataset $C=\{(\omega_{i},v_{i})\}_{i=1}^{T}$, Overfitting tolerance $\epsilon$, Ratio of the holdout set $\delta$.
-2:  randomly split the dataset into a training set $C^{\prime}$ and a holdout set $C\setminus C^{\prime}$ with
-$\lvert C^{\prime}\rvert=(1-\delta)*T$
-3:  $\hat{\pi}_{C^{\prime}}(\omega,v)=\frac{1}{T}\sum_{\hat{\omega},\hat{v}\in C^{\prime}}\mathbb{1}_{\omega=\hat{\omega}}\cdot\mathbb{1}_{v=\hat{v}}$
-4:  $R_{training}=\text{\bf E}_{(v,\omega)\sim C^{\prime}}\![{S(d^{r}_{\hat{\pi}_{C^{\prime}}}(v),\omega)}]$.
-5:  $R_{testing}=\text{\bf E}_{(v,\omega)\sim C\setminus C^{\prime}}\![{S(d^{r}_{\hat{\pi}_{C^{\prime}}}(v),\omega)}]$.
-6:  $bin\_size=\text{the number of unique }v$
-7:  if $R_{training}-R_{testing}>\epsilon*R_{training}$ then
-8:     $bin\_size=bin\_size/2$
-9:     $\{v^{\prime}_{i}\}_{i=1}^{T}=\text{coarsen }\{v_{i}\}_{i=1}^{T}\text{ with binsize }bin\_size$.
-10:     replace $\{v_{i}\}_{i=1}^{T}$ in $C$ with $\{v^{\prime}_{i}\}_{i=1}^{T}$.
-11:     go to step 3.
-12:  end if
-13:  coarsened signal space $\mathbf{V}^{\prime}=\text{unique}(\{v^{\prime}_{i}\}_{i=1}^{T})$
-14:  Output: $\hat{\pi}_{\pi^{C}}$, $\mathbf{V}^{\prime}$
-Denote the optimal coarsened signal space that satisfies Equation 12 as
 […]
-The above algorithm finds the coarsened signal space $\mathbf{V}^{\prime}$ that has the 2-factor approximation of the optimal coarsened signal space $\mathbf{V}^{*}$, $\lvert\mathbf{V}^{\prime}\rvert\geq\frac{1}{2}\lvert\mathbf{V}^{*}\rvert$, in $O(log(\lvert\mathbf{V}^{*}\rvert))$ time.
-B.1 DGP with high-dimensional signals
-The above algorithm’s time complexity is linear in the number of dimensions of the signal space.
-However, in practice, high-dimensional signals, such as images and text that appear in human-AI decision problems can still be computationally infeasible to quantify with the above algorithm.
-To address this issue, we propose a two-step approach to estimate the DGP from a set of observations.
-First, we use a reduction model to reduce the dimensionality of the signal space.
-Second, we use the estimated DGP using the reduced signal space in the above algorithm.
-The optimal reduction model and its generated embeddings should fulfill two properties.
-First, the generated signals should be as informative as possible for the decision task, such that they maintain the maximum information value contained in the original signals.
-Second, the derived DGP that is implied by the embedding signals should be feasible to estimate from the observed dataset.
-We quantify the above objectives by a similar algorithm based on Train/Test split.
+Appendix B Robust Analysis of Information Order
+Our approach assumes a decision problem as input and evaluates agents’ decisions and use of information on this problem.
+However, evaluators may face ambiguity around the appropriate decision problem specification, and in particular, the appropriate scoring rule.
+In particular, ambiguity can arise in payoff functions; doctors, for example, penalize false negative results differently when diagnosing younger versus older patients [Mclaughlin and Spiess, 2023].
+Blackwell’s comparison of signals [Blackwell et al., 1951] is an ideal tool for addressing ambiguity about the payoff function, as it defines a signal $V_{1}$ as more informative than $V_{2}$ if $V_{1}$ has a higher information value on all possible decision problems.
+We identify this partial order by decomposing the space of decision problems via a basis of proper scoring rules [Li et al., 2022, Kleinberg et al., 2023].
+Definition B.1 (Blackwell Order of Information).
+A signal $V_{1}$ is Blackwell more informative than $V_{2}$ if $V_{1}$ achieves a higher best-attainable payoff on any decision problems:
+$\mathrm{R}^{\pi,S}(V_{1})\geq\mathrm{R}^{\pi,S}(V_{2}),\forall S$
+where $\mathrm{R}^{\pi,S}(V)$ denotes the expected performance of the rational DM on payoff function $S$ and information structure $\pi$ when observing $V$.
+The Blackwell order is evaluated over all possible decision problems, which cannot be tested directly.
+Fortunately, we only need to test over all proper scoring rules since any decision problem can be represented by an equivalent proper scoring rule, and the space of proper scoring rules can be characterized by a set of V-shaped scoring rules.
+A V-shaped scoring rule is parameterized by the kink of the piecewise-linear utility function.
+Definition B.2.
+(V-shaped scoring rule) A V-shaped scoring rule with kink $\mu\in(0,\frac{1}{2}]$ is defined as
+$S_{\mu}(d,\omega)=\left\{\begin{array}[]{cc}\frac{1}{2}-\frac{1}{2}\cdot\frac{\omega-\mu}{1-\mu}&\text{if }d\leq\mu\\ \frac{1}{2}+\frac{1}{2}\cdot\frac{\omega-\mu}{1-\mu}&\text{else},\end{array}\right$.
+When $\mu^{\prime}\in(\frac{1}{2},1)$, the V-shaped scoring rule can be symmetrically defined by $S_{\mu^{\prime}}=S_{1-\mu^{\prime}}(1-y,\omega)$.
+Intuitively, the kink $\mu$ represents the threshold belief where the decision-maker switches between two actions.
+Larger $\mu$ means that the decision-makers will prefer $d=1$ more.
+The closer $\mu$ is to $0.5$, the more indifferent the decision-maker is to $d=0$ or $d=1$.
+Proposition B.3 shows that if $V_{1}$ achieves a higher information value on the basis of V-shaped proper scoring rules than $V_{2}$, then $V_{1}$ is Blackwell more informative than $V_{2}$.
+Proposition B.3 follows from the fact that any best-responding payoff can be linearly decomposed into the payoff on V-shaped scoring rules.
+Proposition B.3 (Hu and Wu 2024).
+If $\forall\mu\in(0,1)$
+$\mathrm{R}^{\pi,S_{\mu}}(V_{1})\geq\mathrm{R}^{\pi,S_{\mu}}(V_{2}),$
+then $V_{1}$ is Blackwell more informative than $V_{2}$.
+Extending this to ACIV, $V_{1}$ offers a higher complementary value than $V_{2}$ under the Blackwell order if
+$\operatorname{ACIV}^{\pi,S_{\mu}}(V_{1};D^{b})\geq\operatorname{ACIV}^{\pi,S_{\mu}}(V_{2};D^{b}),\forall\mu\in(0,1)$
+This definition allows us to rank signals (or sets of signals) without needing to commit to a specific payoff function. We present a use case in Appendix E.
 Appendix C Observational Study on the Effect of ACIV on Human-AI Team Performance
@@
 Census 0.1147 [0.1077, 0.1217] 0.1997 0.1861 0.0136
-Appendix D Demonstration I: Model Comparison on Chest Radiograph Diagnosis
+Appendix D Statistical Tests on Experiment Results
+We calculate the p-values of the significance tests with Welch’s t-test, with $\alpha=0.05$. The degrees of freedom ($\nu$) of Welch’s t-test is calculated with the following formula:
+$\nu=\frac{(\frac{s_{1}^{2}}{N_{1}}+\frac{s_{2}^{2}}{N_{2}})^{2}}{\frac{s_{1}^{2}}{N_{1}^{2}(N_{1}-1)}+\frac{s_{2}^{2}}{N_{2}^{2}(N_{2}-1)}}$ (10)
+where $s_{1}$ and $s_{2}$ are the standard deviations of the two samples, and $N_{1}$ and $N_{2}$ are the sizes of the two samples (see sample sizes and standard deviations in Table 3).
+We used Bonferroni correction as the multiple comparisons correction to control the overall type I error rate.
+The results are shown in Table 2.
+Table 2: 95% confidence intervals and significance test results (p-values) for the reduction of APE with different AI models and explanations in Figure 1.
+P(ILIV-SHAP + SHAP
+¿ SHAP)
+P(ILIV-SHAP + SHAP
+¿ No explanation)
+P(AI1 ¿ AI2)
+P(AI1 + ILIV-SHAP + SHAP
+¿ AI2 + ILIV-SHAP + SHAP)
+Reduction of APE with 95% confidence intervals
+No explanation SHAP ILIV-SHAP + SHAP
+AI1
+$5.95\%$
+$5.49\%,6.40\%]$
+$6.06\%$
+$5.65\%,6.47\%]$
+$7.14\%$
+$6.70\%,7.58\%]$
+0.003 / 0.001
+0.311 / ¡0.001
+AI2
+$6.24\%$
+$5.79\%,6.68\%]$
+$6.29\%$
+$5.80\%,6.78\%]$
+$5.47\%$
+$4.95\%,5.98\%]$
+0.138 / 0.159
+Table 3: Sample sizes and standard deviations of the reduction of APE for each experimental condition
+Model Explanation Participants Observations (N) Standard Deviation
+AI1 ILIV-SHAP + SHAP 70 840 0.0653
+SHAP 79 948 0.0642
+No explanation 69 828 0.0668
+Total (all explanations) 218 2,616 0.0656
+AI2 ILIV-SHAP + SHAP 67 804 0.0746
+SHAP 62 744 0.0679
+No explanation 74 888 0.0673
+Total (all explanations) 203 2,436 0.0700
+Total 421 5,052 0.0678
+Appendix E Demonstration I: Model Comparison on Chest Radiograph Diagnosis
 We apply our framework to a well-known cardiac dysfunction diagnosis task [Rajpurkar et al., 2018, Tang et al., 2020, Shreekumar, 2025].
@@
 We list the AI predictions alone to show the AI-complementary information value offered by human decisions.
-D.1 Data and Model
+E.1 Data and Model
 We use data from the MIMIC dataset [Goldberger et al., 2000], which contains anonymized electronic health records from Beth Israel Deaconess Medical Center (BIDMC), a large teaching hospital in Boston, Massachusetts, affiliated with Harvard Medical School.
@@
 The full-information bound is defined as the expected payoff of a rational DM who has access to all signals, human label from radiology report and predictions from five AI models.
-We do not consider the radiology images as a basic signal in this task because of the problem of computational feasibility with high-dimensional signals.
-We use the five vision models’ predictions as embeddings of the original images.
-See more detail on the algorithm of choosing the optimal embeddings in LABEL:app:high-dimensional.
-D.2 Results
+E.2 Results
 Can the AI models complement human judgment?
@@
 We assess the stability of VisionTransformer’s superiority over the other AI models across many possible losses to test if there is a Blackwell ordering of models.
-By Proposition 3.7, we test the payoff of models on all V-shaped scoring rules, shown in Figure 6.
+By Proposition B.3, we test the payoff of models on all V-shaped scoring rules, shown in Figure 6.
 Across all the V-shaped payoff functions, we find that VisionTransformer is Blackwell more informative and Inception v3 is Blackwell less informative than all other models.
 The VisionTransformer achieves a higher information value on all V-shaped scoring rules, implying a higher information value on all decision problems.
-Appendix E Demonstration II: Behavioral Analysis on Deepfake Detection
+Appendix F Demonstration II: Behavioral Analysis on Deepfake Detection
 We apply our framework to analyze a deepfake video detection task [Dolhansky et al., 2020], where participants are asked to judge whether a video was created by generative AI, including with the assistance of an AI model.
@@
 All the payoffs are calculated by $\mathrm{R}^{\pi,S}(\cdot)$, where $\cdot$ is the signal on the y-axis.
-E.1 Data and Model
+F.1 Data and Model
 We define the information model on the experiment data of Groh et al.
@@
 We estimate the data-generating process $\pi$ using the realizations of signals, state, first-round human decisions, AI predictions, and second-round human-AI decisions.
-The no-information bound is the same as Appendix D while the full-information bound is defined as the expected payoff of a rational DM who has access to all video-level features and three agents’ decisions.
-E.2 Results
+The no-information bound is the same as Appendix E while the full-information bound is defined as the expected payoff of a rational DM who has access to all video-level features and three agents’ decisions.
+F.2 Results
 How much decision-relevant information do agents’ decisions offer?
@@
 This finding further confirms the hypothesis that humans are simply relying on AI predictions without processing the information contained in them.
-Appendix F Demonstration III: Information-based Local Explanation on Recidivism Prediction
-We apply our framework to a recidivism prediction task, where the decision-maker decides whether to release a defendant. We demonstrate using the framework to augment SHAP.
-Figure 5: (A) Summary of SHAP explanation and (B) summary of ILIV-SHAP explanation, where each point represents an instance from the test set.
-(C-F) Waterfall plots of applying SHAP and ILIV-SHAP on instance 4 and instance 18.
-The Waterfall plot starts from the prior expected prediction (or ILIV of the prior expected prediction), and then adds features one at a time until reaching the current model output (or ILIV of the current model output).
-See further comparisons between SHAP and ILIV-SHAP in Appendix H.
-F.1 Data and Model
-We use the COMPAS dataset [Angwin et al., 2022], which contains 11,758 defendants with associated features capturing demographics, information about their current charges, and their prior criminal history.
-We merge the dataset with the experimental data from Lin et al.
-[2020] to obtain human decisions, in the form of recidivism decisions on a subset of instances from Angwin et al.
-[2022]’s experiment on laypeople.
-Human decisions were elicited on a 30-point probability scale.
-Merging the two datasets produces 9,269 instances (defendants).
-For the AI model, we trained an extreme gradient boosting (XGBoost) model [Chen and Guestrin, 2016] on a training set with 6,488 instances, achieving an AUC of $0.84$ on a hold-out test set with 2,781 instances.
-We round the model predictions to the same 30-point probability scale available to study participants in Lin et al.
-[2020].
-We use the Brier score as the payoff function, which is commonly used in previous studies on recidivism [Green and Chen, 2019a, Lin et al., 2020, Fogliato et al., 2021].
-The binary payoff-related state $\omega$ indicates whether the defendant gets rearrested within two years.
-We use the features of the defendant contained in the COMPAS dataset as the signals in our demonstration: demographic features (age, sex, and race), information about the current charge (type of offense and whether it is a misdemeanor or a felony), prior criminal history (e.g., past number of arrests and charges for several offense categories), and the predicted score from the COMPAS system [^6] https://doc.wi.gov/Pages/AboutDOC/COMPAS.aspx.
-We take the empirical distribution of the hold-out test set as the data-generating process, which defines the joint distribution of state, signals, human decisions, and AI predictions.
-F.2 Results
-SHAP explanations do not communicate the complementary information of AI predictions.
-Specifically, the changes in information value of AI predictions caused by marginalizing out features, i.e., ILIV-SHAP scores, are not predictable from the changes in AI predictions, i.e., SHAP scores, and the feature values.
-Figure 5 (A) and (B) compare the distribution of feature importance scores from SHAP to those from ILIV-SHAP.
-The scores of prior_counts (the number of prior criminal charges), age, and decile_score (the score from the COMPAS system), which are three features with the largest scores in both SHAP and ILIV-SHAP, show different correlations with the feature values in SHAP and ILIV-SHAP.
-For instance, in Figure 5(A), most of the low values (blue dots) of prior_counts have a negative SHAP score, and most of the high values (red dots) have a positive SHAP score.
-This indicates that when the value of prior_counts is low, marginalizing out prior_counts from the features leads to a higher expected model prediction than not marginalizing out, and vice versa.
-However, this correlation does not hold for ILIV-SHAP.
-In Figure 5(B), the high values of prior_counts appear on both the left side (below zero) and the right side (above zero), indicating having low ILIV-SHAP scores and high ILIV-SHAP scores at the same time.
-This mismatch between the distributions of SHAP scores and ILIV-SHAP scores implies that the impact of features on the information value cannot be directly inferred (at least monotonically) from their impact on model output and their values.
-ILIV-SHAP can augment SHAP on communicating the complementary information of AI predictions.
-We argue that ILIV-SHAP can supplement saliency-based explanations to provide users with a sense of how the AI predictions can help improve their decisions.
-With ILIV-SHAP, users can check whether the prediction contains decision-relevant information by looking at the total $\operatorname{ILIV}$ of the AI prediction, which indicates the maximum improvement in expected payoff that could be achieved by incorporating this prediction.
-Figure 5(D) and (F) show two different cases, where the AI predictions contain substantial human-complementary information value in (D) (shown by the distance between the tail and the head of the Waterfall plot) but very little information in (F).
-ILIV-SHAP also tells the user whether an AI model’s prediction encodes useful complementary information from a particular feature.
-In Figure 5(E), the defendant’s prior count record (prior_counts=1) ranks second among features that impact the AI prediction ($-0.139$).
-Upon seeing this, users might be interested in investigating how the AI uses prior_counts=1.
-However, Figure 5(F) shows that prior_counts=1 contributes marginally to the decision-relevant information in AI predictions (about $+0.007$).
-Thus, with ILIV-SHAP, users can see that even if prior_counts=1 changes the AI prediction significantly, the AI does not predict more accurately as a result of it.
 Appendix G Robustness Analysis in Demonstration I
@@
 The subplots are symmetric along the diagonal, e.g., $(1,2)$ subplot and $(2,1)$ subplot display the same distribution with opposite signs.
-Appendix H More examples of ILIV-SHAP
-(a) SHAP on instance 0
-(b) ILIV-SHAP on instance 0
-(a) SHAP on instance 1
-(b) ILIV-SHAP on instance 1
-(a) SHAP on instance 2
-(b) ILIV-SHAP on instance 2
-Appendix I Screenshots of the Experiment
+Appendix H Screenshots of the Experiment
 Figure 7: Screenshot of the human-alone trials in the first round of the experiment.
@@
 Figure 9: Screenshot of the SHAP condition of AI1 in the second round of the experiment.
-Figure 13: Screenshot of the SHAP condition in the second round of the experiment.
+Figure 10: Screenshot of the SHAP condition of AI2 in the second round of the experiment.
 Figure 11: Screenshot of the ILIV-SHAP + SHAP condition of AI1 in the second round of the experiment.

### v6

@@
-The Value of Information in Human-AI Decision-making
+Explaining and Improving Information Complementarities in Multi-Agent Decision-making
 Abstract
@@
 We present a novel explanation technique (ILIV-SHAP) that adapts SHAP explanations to highlight human-complementing information.
-We validate the effectiveness of ACIV and ILIV-SHAP through a study of human-AI decision-making, and demonstrate the framework on examples from chest X-ray diagnosis and deepfake detection.
+We validate the effectiveness of our framework and ILIV-SHAP through a study of human-AI decision-making, and demonstrate the framework on examples from chest X-ray diagnosis and deepfake detection.
 We find that presenting ILIV-SHAP with AI predictions leads to reliably greater reductions in error over non-AI assisted decisions more than vanilla SHAP [^1] The code to calculate the main quantities in our framework and reproduce the experimental results is available at https://osf.io/p2qzy/?view_only=bf39de5d96f047f69e45ffd42689ebf9..
@@
 For example, a physician may have access to information that is not captured in structured health records (Alur et al., 2024).
-Evidence of complementary performance between humans and AI nonetheless remains limited, with many studies showing that human-AI teams underperform the AI alone (Buçinca et al., 2020; Bussone et al., 2015; Green and Chen, 2019; Jacobs et al., 2021; Lai and Tan, 2019; Vaccaro and Waldo, 2019; Kononenko, 2001).
-A rigorous understanding of such results is limited by the fact that most analyses of human-AI decision-making focus on ranking the performance of human-AI teams or agents individually using measures based on post-hoc decision accuracy (Passi and Vorvoreanu, 2022).
-However, such approaches do not account for the best achievable performance given the information available at the time of the decision (Kleinberg et al., 2015; Guo et al., 2024; Rambachan, 2024).
-Moreover, existing approaches cannot provide insight on the potential for available information in the environment to improve performance.
-Consequently, it is difficult to design interventions that improve the team’s performance.
-Imagine one could identify information complementarities that can be exploited, such as when one of the agents has access to information not contained in the other’s judgments, or has not fully integrated contextually-available information in their judgments.
+Many empirical studies, however, have found that human-AI teams underperform the AI alone (Buçinca et al., 2020; Bussone et al., 2015; Green and Chen, 2019; Jacobs et al., 2021; Lai and Tan, 2019; Vaccaro and Waldo, 2019; Kononenko, 2001).
+Two sources of ambiguity complicate such results.
+One concerns the role of measurement: performance is often scored against post-hoc decision accuracy (Passi and Vorvoreanu, 2022) rather than accounting for the best achievable performance given information available at the time of the decision (Kleinberg et al., 2015; Guo et al., 2024; Rambachan, 2024).
+Additionally, it often remains unclear how agents differed in their information access or use, making it difficult to design interventions to improve these aspects.
+Imagine one could identify information complementarities that can be exploited, such as when one of the agents has access to information not contained in the other’s judgments, or has not fully integrated contextually-available information (e.g., instance features) in their judgments.
 This would motivate interventions to improve decision-making.
@@
 (2024) finds that, on average, human–AI teams perform worse than the better of the two agents alone.
-In response, a growing body of work seeks to evaluate and enhance complementarity in human–AI systems (Bansal et al., 2021b, 2019, a; Wilder et al., 2021; Rastogi et al., 2023; Mozannar et al., 2024b).
+In response, a growing body of work seeks to evaluate and enhance complementarity in human–AI systems (Bansal et al., 2021b; 2019; a; Wilder et al., 2021; Rastogi et al., 2023; Mozannar et al., 2024b).
 The present work differs from much of these prior works by approaching human-AI complementarity from the perspective of information value and use, including asking whether the human and AI decisions provide additional information that is not used by the other.
 Evaluation of human decision-making with machine learning.
-Our work contributes methods for evaluating the decisions of human-AI teams (Kleinberg et al., 2015, 2018; Lakkaraju et al., 2017; Mullainathan and Obermeyer, 2022; Rambachan, 2024; Guo et al., 2024; Ben-Michael et al., 2024; Shreekumar, 2025).
+Our work contributes methods for evaluating the decisions of human-AI teams (Kleinberg et al., 2015; 2018; Lakkaraju et al., 2017; Mullainathan and Obermeyer, 2022; Rambachan, 2024; Guo et al., 2024; Ben-Michael et al., 2024; Shreekumar, 2025).
 Kleinberg et al.
@@
 This benchmark provides a basis for identifying exploitable information within a decision problem.
-Human information in machine learning.
-Some approaches focus on automating the decision pipeline by explicitly incorporating human expertise in developing machine learning models, such as by learning to defer (Mozannar et al., 2024a; Madras et al., 2018; Raghu et al., 2019; Keswani et al., 2022, 2021; Okati et al., 2021; Chen et al., 2022).
+Complementarity by design.
+Some approaches focus on automating the decision pipeline by explicitly incorporating human expertise in developing machine learning models or human-AI collaboration pipeline, such as by learning to defer (Mozannar et al., 2024a; Madras et al., 2018; Raghu et al., 2019; Keswani et al., 2022; 2021; Okati et al., 2021; Chen et al., 2022).
 Corvelo Benz and Rodriguez (2023) propose multicalibration over human and AI model confidence information to guarantee the existence of an optimal monotonic decision rule.
+Other approaches exploit information asymmetry (i.e., cases where humans have additional contextual knowledge) and offer principled methods with provable guarantees to improve team decision performance (Straitouri et al., 2023; De Toni et al., 2024; Arnaiz-Rodriguez et al., 2025).
 Alur et al.
-(2023) propose a hypothesis testing framework to evaluate the added value of human expertise over AI forecasts.
-Our work shares the motivation of incorporating human expertise, but provides a general analytical framework for quantifying the information value of all available signals and agent decisions in human–AI decision tasks.
+(2024) propose a framework to incorporate human decisions into machine learning algorithms when the state is indistinguishable from the algorithm alone but can be discriminated by humans.
+Our work also concerns information asymmetry, but provides an interpretable analytical framework for quantifying the information value of all available signals and agent decisions in human–AI decision tasks, enabling the design of information-based interventions.
 3 Framework
@@
 We cast the information available to a DM,including any available agent decisions, as a set of signals defined within an information model.
-Following the definition of an information model in Blackwell et al.
-(1951), the information model can be represented by a data-generating model with a set of signals.
+Following the definition of an information model in Blackwell and others (1951), the information model can be represented by a data-generating model with a set of signals.
   • Signals.
 There are $n$ “basic signals” represented as random variables $\Sigma_{1},\ldots,\Sigma_{n}$, from the signal spaces $\mathbf{\Sigma}_{1},\ldots,\mathbf{\Sigma}_{n}$.
-Basic signals represent information available to a decision-maker as they decide, e.g., $\mathbf{\Sigma}_{1}\in\Delta\mathbf{\Omega}$ for a probabilistic prediction of raining, $\mathbf{\Sigma}_{2}=\{\text{cloudy},\text{not cloudy}\}$, $\mathbf{\Sigma}_{3}\in\{0,\ldots,100\}$ for temperature in Celsius, etc.
+Basic signals represent information available to a decision-maker as they decide, e.g., $\mathbf{\Sigma}_{1}=\Delta\mathbf{\Omega}$ for a probabilistic prediction of raining, $\mathbf{\Sigma}_{2}=\{\text{cloudy},\text{not cloudy}\}$, $\mathbf{\Sigma}_{3}=\{0,\ldots,100\}$ for temperature in Celsius, $\mathbf{\Sigma}_{4}=\mathbf{D}$ for human decisions $D^{H}$ on taking umbrella or not, etc.
 The decision-maker observes a signal, which is a subset of the basic signals, $V\subseteq 2^{\{\Sigma_{1},\dots,\Sigma_{n}\}}$.
 The combination of two signals $V_{1}$ and $V_{2}$ takes the set union $V=V_{1}\cup V_{2}$.
-For example, in the standard human-AI decision workflow where a human makes an independent judgment before consulting the AI, the set of all basic signals is $V=\{x,D^{\text{H}},D^{\text{AI}},e(D^{\text{AI}},x)\}$–the features of the instance, the human’s initial decisions, the AI’s predictions, and any explanation of the AI’s predictions.
+In the standard human-AI decision workflow where a human makes an independent judgment before consulting the AI, all basic signals are $\{x,D^{\text{H}},D^{\text{AI}}\}$–the features of the instance, the human’s initial decisions and the AI’s predictions.
   • Data-generating process.
@@
 $\operatorname{IV}(V):=\mathrm{R}(V)-\mathrm{R}(\emptyset)$ (4)
-We adopt the same idea below to define the agent-complementary information value.
+The full information value in a human-AI decision task is the information value of the set of all basic signals.
+For example, in the weather decision example, $\operatorname{IV}(\{\Sigma_{1},\Sigma_{2},\Sigma_{3},\Sigma_{4}\})$ represents the full information value of the probabilistic prediction of rain, the cloudiness, the temperature, and the human decisions.
+This defines the upper bound of the information value of any signal, including the agent-complementary information value and instance-level agent-complementary information value that we will define in the following sections.
 3.1 Agent-Complementary Information Value
@@
 The intuition behind our approach is that any information that is used by decision-makers should eventually reveal itself through variation in their decisions.
-We recover the information value in agent decisions by offering the decisions as a signal to the Bayesian rational DM.
+Definition˜3.2 captures how much complementary information value is offered by a signal over the agent decisions in expectation over the data-generating process.
+For example, in the weather decision task, this indicates how much the payoff from the decision-maker’s choice of whether to take the umbrella or not can be further improved by incorporating the temperature in their decision rule.
 Definition 3.2.
@@
 A natural relaxation is to consider the realizations that are sufficiently “similar” for the performance of the decision-maker.
-Algorithm 1 learns the posterior distribution of the payoff-relevant state from the signal and then uses that prediction as the probability distribution of the payoff-relevant state to choose the optimal action.
+Algorithm˜1 learns the posterior distribution of the payoff-relevant state from the signal and then uses that prediction as the probability distribution of the payoff-relevant state to choose the optimal action.
 This approach can be readily extended to the case where the signal is more complex, such as human decisions in the form of freeform text (e.g., radiology reports) or explanations of the AI predictions in the form of images (e.g., saliency map).
+[^3] Note that the algorithm $(\hat{a},\hat{a}^{b})$ should be cross-validated to avoid overfitting to the observed data and also be evaluated for calibration error since the rational DM treats it as the Bayesian posterior.
+We provide a sensitivity analysis in Appendix H where we compare the ACIV estimated under the rational belief estimator using linear regression (LR), gradient boosting methods (GBM), and neural networks (NN).
 Algorithm 1 A method to calculate the ACIV of $V$.
-1:Input: Observed realizations $\{v_{i},d^{b}_{i},\omega_{i}\}_{i=1}^{n}$, a regression algorithm $\mathcal{A}$, payoff function $S$, decision space $\mathbf{D}$, and state space $\mathbf{\Omega}$.
+1:Input: Observed realizations $\{v_{i},d^{b}_{i},\omega_{i}\}_{i=1}^{n}$, a predictive algorithm $\mathcal{A}$, payoff function $S$, decision space $\mathbf{D}$, and state space $\mathbf{\Omega}$.
 2:Output: $\operatorname{ACIV}(V;D^{b})$
@@
 6:  $\hat{p}_{i}\leftarrow\hat{a}(v_{i},d^{b}_{i})$
-7:  $\hat{d}^{r}_{i}\leftarrow\operatorname{argmax}_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\hat{p}_{i}}\![{S(d,\omega)}]$
+7:  $\hat{d}^{r}_{i}\leftarrow\operatorname*{arg\,max}_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\hat{p}_{i}}\![{S(d,\omega)}]$
 8:  $\hat{p}^{b}_{i}\leftarrow\hat{a}^{b}(d^{b}_{i})$
-9:  $\hat{d}^{rb}_{i}\leftarrow\operatorname{argmax}_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\hat{p}^{b}_{i}}\![{S(d,\omega)}]$
+9:  $\hat{d}^{rb}_{i}\leftarrow\operatorname*{arg\,max}_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\hat{p}^{b}_{i}}\![{S(d,\omega)}]$
 10:  $s_{i}\leftarrow S(\hat{d}^{r}_{i},\omega_{i})-S(\hat{d}^{rb}_{i},\omega_{i})$
@@
 This finer-grained view makes it possible to analyze how much an agent can benefit in theory from better incorporating instance-level information in their decision.
+For example, in the weather decision task, this indicates how much the payoff of the decision-maker can be improved by knowing the temperature is 21C.
 Given a realization of signal $v=\{\sigma_{j_{1}},\ldots,\sigma_{j_{k}}\}$, we want to know the maximum expected payoff gain from the access to $v$ on the instances where $v$ is realized over the existing information encoded in agent decisions.
@@
 $\mathrm{r}^{v}(v^{\prime}):=\text{\bf E}_{\omega\sim\pi(\omega|v)}\![{S(d^{r}(v^{\prime}),\omega)}]$ (6)
-where $d^{r}(v^{\prime})$ is the Bayesian optimal decision on receiving $v^{\prime}$ as defined in Equation 2.
+where $d^{r}(v^{\prime})$ is the Bayesian optimal decision on receiving $v^{\prime}$ as defined in Equation˜2.
 Note that we allow $v^{\prime}\neq v$, i.e.
@@
 The expected payoff of a rational DM who is misinformed is guaranteed to be lower than is correctly informed, i.e., $\mathrm{r}^{v}(v^{\prime})\leq\mathrm{r}^{v}(v)$ for any $v^{\prime}$.
-This notion of flexibility allows to consider the information value of a counterfactual realization of a signal, which we use in designing the explanation of ILIV-SHAP in the next section.
+This notion of flexibility allows to consider the information value of a counterfactual realization of a signal.
+For example, in the weather decision task, this notation is able to describe how much payoff can change when the decision maker is misinformed that the temperature is 18C when the temperature is actually 21C.
+We use this in designing the explanation of ILIV-SHAP in the next section.
 If we consider the agent decisions in addition to the realization $v$, the rational DM’s expected payoff on instances where $V=v$ can be written as
@@
 $\operatorname{ILIV}^{v}(v^{\prime};D^{b}):=\mathrm{r}^{v}(v^{\prime};D^{b})-\mathrm{r}^{v}(\emptyset;D^{b}).$ (8)
-where $\mathrm{r}^{v}(\emptyset;D^{b})$ represents the expected rational payoff on instances of $V=v$, where the rational DM only knows the information encoded in agent decisions without knowing any realizations of $V$.
+where $\mathrm{r}^{v}(\emptyset;D^{b})$ represents the expected rational payoff on instances of $V=v$, where the rational DM only knows the agent decisions without knowing any realizations of $V$.
 $\operatorname{ILIV}$ maximizes when the signal does not misinform, i.e., $\operatorname{ILIV}^{v}(v;D^{b})\geq\operatorname{ILIV}^{v}(v^{\prime};D^{b}),\text{for }v^{\prime}\in\mathbf{V}$.
 Algorithm 2 A method to calculate the ILIV of $v^{\prime}$ on the instances of $v$.
-1:Input: Observed realizations $\{v_{i},d^{b}_{i},\omega_{i}\}_{i=1}^{n}$, test counterfactual signal realization $v^{\prime}$, test signal realization $v$, a regression algorithm $\mathcal{A}$, payoff function $S$, decision space $\mathbf{D}$, and state space $\mathbf{\Omega}$.
+1:Input: Observed realizations $\{v_{i},d^{b}_{i},\omega_{i}\}_{i=1}^{n}$, test counterfactual signal realization $v^{\prime}$, test signal realization $v$, a predictive algorithm $\mathcal{A}$, payoff function $S$, decision space $\mathbf{D}$, and state space $\mathbf{\Omega}$.
 2:Output: $\operatorname{ILIV}^{v}(v^{\prime};D^{b})$
@@
 7:  $\hat{p}_{i}\leftarrow\hat{a}(v^{\prime},d^{b}_{j_{i}})$
-8:  $\hat{d}^{r}_{i}\leftarrow\operatorname{argmax}_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\hat{p}_{i}}\![{S(d,\omega)}]$
+8:  $\hat{d}^{r}_{i}\leftarrow\operatorname*{arg\,max}_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\hat{p}_{i}}\![{S(d,\omega)}]$
 9:  $\hat{p}^{b}_{i}\leftarrow\hat{a}^{b}(d^{b}_{j_{i}})$
-10:  $\hat{d}^{rb}_{i}\leftarrow\operatorname{argmax}_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\hat{p}^{b}_{i}}\![{S(d,\omega)}]$
+10:  $\hat{d}^{rb}_{i}\leftarrow\operatorname*{arg\,max}_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\hat{p}^{b}_{i}}\![{S(d,\omega)}]$
 11:  $s_{i}\leftarrow S(\hat{d}^{r}_{i},\omega_{i})-S(\hat{d}^{rb}_{i},\omega_{i})$
@@
 For example, ILIV-SHAP satisfies the critical efficiency axiom, i.e., the sum of the importance scores equals the information value of the model output, and the symmetry axiom, i.e., the importance scores are the same for any two features that contribute the same amount to the information value of the model output.
+Sample-based methods for approximating SHAP values [^4] We note that ILIV-SHAP does not resolve the foundational issues of SHAP (Huang and Marques-Silva, 2024), but just as a demonstration on how the information value such as ILIV can be used in explanations.
+can also be applied to ILIV-SHAP, such as permutation sampling (Strumbelj and Kononenko, 2010), Kernel SHAP (Lundberg and Lee, 2017), and Partition SHAP (Lundberg et al., 2018).
+Because $\operatorname{ILIV}$ is expected to be non-decreasing in the number of features included in $g_{f}(\mathbf{x}^{\prime})$—more features mean more information—sample-based methods can achieve better stability across the permutations on ILIV-SHAP than on SHAP.
 5 Experiment
@@
 We hypothesize that the higher ACIV model will lead to better human-AI decisions, and that presenting ILIV-SHAP explanations for this model will lead to better human-AI decisions than the other explanation types.
-Metric AI1 (higher ACIV) AI2 (lower ACIV)
-MAPE 14.30% 14.51%
-R squared 0.81 0.81
-$\operatorname{ACIV}$
-(in MSE)
-$6.5\times 10^{8}$ $3.7\times 10^{8}$
-$\operatorname{ACIV}$
-(in MAPE)
-4.61% 2.00%
-Figure 1: Left: AI performance and ACIV.
-Right: Reduction in the human-AI team’s absolute percentage error (APE) with 95% confidence intervals according to posterior predictions of our regression model.
 Task and data.
@@
 We incentivized participants with a base reward of $3.00$ USD and a bonus based on the mean squared error (MSE) between their guesses and the true house price in the 24 trials: $\textit{bonus}=3.00-\frac{\textit{MSE}}{3\times 10^{9}}$.
+(a) AI1 with ILIV-SHAP and SHAP explanation.
+(b) AI2 with ILIV-SHAP and SHAP explanation.
+Figure 1: The screenshot of the interface for ILIV-SHAP and SHAP explanations.
+The third column is the SHAP value and the highlights are based on the ILIV-SHAP value.
+Because AI2 does not have access to Feature X/Y, both their ILIV-SHAP and SHAP values are zero for AI2.
 AI models and explanations.
-Participants were assigned to one of six experimental conditions resulting from crossing the two AI models with the three explanation types.
-We designed AI1 and AI2 to have varying potential to complement humans: AI1 was trained with the six features and the true house price, with noise added to the six features (which is meant to be bring down the accuracy of AI1 to be comparable with AI2).
-On the other hand, AI2 was trained with the four human-interpretable features and the true house price, such that there would be less potential for the predictions to add complementary information over human judgments.
-We used the same training data for both AI models and ensured that the two AI models achieved similar performance (Figure 1 left), so that model accuracy differences would not confound our results.
-In the SHAP explanation, we used the SHAP values of the six features to explain the AI’s prediction.
-To generate the ILIV-SHAP explanations, we used independent human decisions we collected on Prolific for the dataset using the same task and features prior to running the official study.
-We present ILIV-SHAP explanations of the six features, which ranked and highlighted the features where the human-complementary information (ILIV) over the AI’s prediction exceeds a threshold (see Appendix H for explanation examples).
+[…]
 Evaluation metrics.
@@
 and mean absolute percentage error (MAPE).
-To measure the human alone and human-AI team performance, we fit a preregistered Bayesian hierarchical regression model with weakly informed priors444link to preregistration redacted for peer review to the percentage error (PE) between the human’s prediction and the true house price, $\textit{PE}=(d-\omega)/\omega$, using R’s brms package (Bürkner, 2017).
+To measure the human alone and human-AI team performance, we fit a preregistered Bayesian hierarchical regression model with weakly informed priors777link to preregistration redacted for peer review to the percentage error (PE) between the human’s prediction and the true house price, $\textit{PE}=(d-\omega)/\omega$, using R’s brms package (Bürkner, 2017).
 PE $\sim\text{student\_t}(\mu,\sigma,\nu)$
@@
 We evaluate human-AI team performance by the reduction in absolute percentage error (APE) over human-alone predictions in the first round, i.e., $\text{Reduction in APE}=\text{\bf E}\![{\lvert\textit{PE}\rvert\mid\textit{round}=1}]-\text{\bf E}\![{\lvert\textit{PE}\rvert\mid\textit{round}=2}]$, where the expectation is taken over the Bayesian posterior distribution of the above Bayesian model.
+Table 1: The two AI models that are used in the experiment. AI1 has access to all the features, while AI2 has access to only the human-interpretable features.
+Input features MAPE R squared $\operatorname{ACIV}$ (in MSE) $\operatorname{ACIV}$ (in MAPE)
+AI1 All features 14.30% 0.81 6.5 $\times$ 108 4.61%
+AI2 Human-interpretable features 14.51% 0.81 $3.7\times 10^{8}$ 2.00%
+Figure 2: Reduction in the human-AI team’s absolute percentage error (APE) with 95% confidence intervals according to posterior predictions of our regression model.
 Results.
-Figure 1 confirms the expected ranking of AI models by $\operatorname{ACIV}$.
+Table˜1 confirms the expected ranking of AI models by $\operatorname{ACIV}$.
 AI1, which had access to the two human-uninterpretable features, achieves higher $\operatorname{ACIV}$ than AI2, which only had access to the four human-interpretable features.
-Note the similar predictive performance of the two AI models.
-The posterior predictions of human-AI team performance from our regression model also align with expectations based on the complementary information each model provides (Figure 1 far right).
-Collapsing across explanation conditions, AI1 (higher ACIV) reduces APE by more than AI2 ($6.27\%$$5.91\%,$$6.63\%]$ vs $5.65\%[5.26\%,6.05\%]$ respectively).
-Access to the ILIV-SHAP + SHAP explanation results in a greater reduction in APE for AI1 than AI2 ($7.11\%[6.57\%,7.65\%]$ vs $5.83\%[5.17\%,6.49\%]$ respectively) [^5] See Appendix D for the significant test results.
-Figure 1 also illustrates the effect of the different explanation conditions on human-AI team performance.
-When the AI has sufficient complementing information (AI1), the ILIV-SHAP + SHAP explanation more effectively reduces APE for the human-AI decisions than the SHAP explanation and baseline (no explanation) ($7.11\%[6.57\%,7.65\%]$ for ILIV-SHAP, $5.20\%[4.53\%,5.87\%]$ for SHAP, and $6.50\%[5.86\%,7.14\%]$ for no explanation).
+Note the similar predictive performance of the two AIs.
+The posterior predictions of human-AI team performance from our regression model also align with expectations based on the complementary information each model provides (Figure˜2 far right).
+Collapsing across explanation conditions, AI1 (higher ACIV) reduces APE by more than AI2 ($6.24\%$$5.99\%,$$6.5\%]$ vs $5.96\%[5.68\%,6.24\%]$ respectively).
+Access to the ILIV-SHAP + SHAP explanation results in a greater reduction in APE for AI1 than AI2 ($6.94\%[6.50\%,7.38\%]$ vs $5.31\%[4.80\%,5.83\%]$ respectively) [^8] See Appendix D for the significant test results.
+Figure˜2 also illustrates the effect of the different explanation conditions on human-AI team performance.
+When the AI has sufficient complementing information (AI1), the ILIV-SHAP + SHAP explanation more effectively reduces APE for the human-AI decisions than the SHAP explanation and baseline (no explanation) ($6.94\%[6.50\%,7.38\%]$ for ILIV-SHAP, $5.88\%[5.47\%,6.28\%]$ for SHAP, and $5.96\%[5.50\%,6.42\%]$ for no explanation).
 6 Demonstrations
@@
 Can the AI models complement human judgment?
-Figure 2(a) shows all the models offer complementary information value to the radiologists’ reports (green distributions improve over the purple distribution), and in the other direction, the radiologists’ reports offer complementary information value to the models (green distributions improve over the orange distributions).
+Figure˜3(a) shows all the models offer complementary information value to the radiologists’ reports (green distributions improve over the purple distribution), and in the other direction, the radiologists’ reports offer complementary information value to the models (green distributions improve over the orange distributions).
 Which AI model offers the most decision-relevant information over human judgments?
-Figure 2(a) shows that VisionTransformer contains slightly higher information value than the other models, and Inception v3 contains slightly lower information value than the other models.
-We further assess the robustness of VisionTransformer’s superiority over the other AI models across many possible payoff functions to test if there is a Blackwell ordering of models in Appendix F and Figure 6.
+Figure˜3(a) shows that VisionTransformer contains slightly higher information value than the other models, and Inception v3 contains slightly lower information value than the other models.
+We further assess the robustness of VisionTransformer’s superiority over the other AI models across many possible payoff functions to test if there is a Blackwell ordering of models in Appendix˜G and Figure˜7.
 (a) Chest X-ray diagnosis results.
 (b) Deepfake detection results
-Figure 2: Demonstration results for both tasks.
-Distributions show information values.
+Figure 3: Distributions show information values.
 For a), we plot ACIV of radiologist decisions, different choices of AI models and signals that combine AI predictions and radiologist decisions.
 For b), we plot the information value of the combination of video-level features and agent decisions (including human decisions, AI predictions, and human-AI teams’ decisions).
-These two figures demonstrate that our framework can be used to a) reveal the complementary information between AI predictions and human decisions and b) analyze how humans use information in data features differently from AI.
+Full information represents all the information available to the human decision-makers–the radiology images for the chest X-ray diagnosis task and the seven video-level features, human decisions, and AI predictions for the deepfake detection task.
+See Appendix˜H for the sensitivity analysis of Algorithm˜1.
 As shown in this demonstration, doctors may use our framework to learn how much complementary information value the AI models offer over their decisions, and which model offers the most.
-[^6] For space constraints, we refer to the full descriptions and results in Appendices F and E.
 6.2 Deepfake Detection
@@
 Groh et al.
-(2022) manually label seven video-level features, which we use as binary indicators in place of Algorithm 1 in light of the high dimensional signals: graininess, blurriness, darkness, presence of a flickering face, presence of two people, presence of a floating distraction, and the presence of an individual with dark skin.
-We show the ACIVs of the seven features over the decisions, the AI predictions, and the human-AI decisions in Figure 2(b).
+(2022) manually label seven video-level features, which we use as binary indicators in place of Algorithm˜1 in light of the high dimensional signals: graininess, blurriness, darkness, presence of a flickering face, presence of two people, presence of a floating distraction, and the presence of an individual with dark skin [^9] The “dark-skin individual” label reflects a subjective visual attribute.
+It is used here only for robustness analysis—not for causal interpretation or demographic inference..
+We show the ACIVs of the seven features over the decisions, the AI predictions, and the human-AI decisions in Figure˜3(b).
 How much decision-relevant information do agents’ decisions offer?
 We first compare the information value of the AI predictions to that of the human decisions.
-Figure 2(b) shows that AI predictions provide about $65\%$ of the total possible information value over the no-information baseline, while human decisions only provide about $15\%$.
+Figure˜3(b) shows that AI predictions provide about $65\%$ of the total possible information value over the no-information baseline, while human decisions only provide about $15\%$.
 We next consider the human-AI decisions.
@@
 To understand what information might improve human decisions, we assess the $\operatorname{ACIV}$s of different video-level features over different agents.
-As shown on the fifth row in Figure 2(b), the presence of a flickering face offers larger $\operatorname{ACIV}$ over human decisions than over AI predictions, meaning that human decisions could improve by a greater amount if they were to incorporate this information.
-Meanwhile, as shown on the fourth row in Figure 2(b), the presence of an individual with dark skin offers larger $\operatorname{ACIV}$ over AI predictions than over human decisions, suggesting that humans make greater use of this information.
+As shown on the fifth row in Figure˜3(b), the presence of a flickering face offers larger $\operatorname{ACIV}$ over human decisions than over AI predictions, meaning that human decisions could improve by a greater amount if they were to incorporate this information.
+Meanwhile, as shown on the fourth row in Figure˜3(b), the presence of an individual with dark skin offers larger $\operatorname{ACIV}$ over AI predictions than over human decisions, suggesting that humans make greater use of this information.
 This suggests that the AI and human rely on differing information to make their initial predictions.
+[^10] For space constraints, we refer to the full descriptions and results in Appendices G and F.
 7 Discussion and Limitations
@@
 Our work provides a promising methodological framework for an emerging research agenda around optimally combining agents’ information for decisions (e.g., Alur et al.
-(2023, 2024)).
+(2023; 2024)).
 While we focused on human-AI decisions, the framework an be applied to any combination of AI or human agent judgments.
 Though most of our definitions and analysis are focused on decision tasks with single well-defined payoff function, the framework can be readily extended to more complex decision tasks with an ambiguous set of payoff functions or totally unidentifiable payoff functions.
-We present a robustness analysis framework in Appendix B to calculate the Blackwell ordering of models over proper scoring rules.
-A demonstration of the robustness analysis is shown in Appendix F.
+We present a robustness analysis framework in Appendix˜B to calculate the Blackwell ordering of models over proper scoring rules.
+A demonstration of the robustness analysis is shown in Appendix˜G.
+The experiment is designed to provide complementarity by construction in order to function as a proof of concept.
+We believe future work is needed to evaluate the impact of ILIV-SHAP in real deployment of human-AI decision-making workflows, such as where the human or AI might have additional private information over each other.
+8 Reproducibility Statement
+The main results of our paper is to provide a framework guiding the analysis of complementary information in human-AI decision-making.
+We provide a python library in supplementary martial to calculate the quantities defined in our framework.
+We provide the data and code (which are all put in separate Jupyter notebooks) to reproduce all the results in our paper, including the demonstrations and the empirical study in the main text, and the observational study and robustness analysis in the Appendix.
 Appendix A The Combinatorial Nature of the Value of Signals
 When decision-makers are provided with multiple signals, the signals have the combinatorial property by nature.
-Acknowledgd by recent works in decision theory and game theory [Chen and Waggoner, 2016], one signal may have no information value by itself, but it can be complementary to other signals to provide information value.
+Acknowledgd by recent works in decision theory and game theory (Chen and Waggoner, 2016), one signal may have no information value by itself, but it can be complementary to other signals to provide information value.
 For example, two signals $\Sigma_{1}$ and $\Sigma_{2}$ are uniformly random bits and the state $\omega=\Sigma_{1}\oplus\Sigma_{2}$, the XOR of $\Sigma_{1}$ and $\Sigma_{2}$.
@@
 Though we did not explicitly observe the complementation between signals in our survey of human-AI decision-making tasks (see the results of deepfake detection in supplementary materials), we want to note that our framework can be extended to consider this complementation between signals.
-We use the Shapley value [Shapley, 1953] to interpret the contribution to the $\operatorname{ACIV}$ of each basic signal.
+We use the Shapley value (Shapley, 1953) to interpret the contribution to the $\operatorname{ACIV}$ of each basic signal.
 $\phi$ is the average of the marginal contribution of a signal in every combination with other signals.
@@
 However, evaluators may face ambiguity around the appropriate decision problem specification, and in particular, the appropriate scoring rule.
-In particular, ambiguity can arise in payoff functions; doctors, for example, penalize false negative results differently when diagnosing younger versus older patients [Mclaughlin and Spiess, 2023].
-Blackwell’s comparison of signals [Blackwell et al., 1951] is an ideal tool for addressing ambiguity about the payoff function, as it defines a signal $V_{1}$ as more informative than $V_{2}$ if $V_{1}$ has a higher information value on all possible decision problems.
-We identify this partial order by decomposing the space of decision problems via a basis of proper scoring rules [Li et al., 2022, Kleinberg et al., 2023].
+In particular, ambiguity can arise in payoff functions; doctors, for example, penalize false negative results differently when diagnosing younger versus older patients (Mclaughlin and Spiess, 2023).
+Blackwell’s comparison of signals (Blackwell and others, 1951) is an ideal tool for addressing ambiguity about the payoff function, as it defines a signal $V_{1}$ as more informative than $V_{2}$ if $V_{1}$ has a higher information value on all possible decision problems.
+We identify this partial order by decomposing the space of decision problems via a basis of proper scoring rules (Li et al., 2022; Kleinberg et al., 2023).
 Definition B.1 (Blackwell Order of Information).
@@
 The closer $\mu$ is to $0.5$, the more indifferent the decision-maker is to $d=0$ or $d=1$.
-Proposition B.3 shows that if $V_{1}$ achieves a higher information value on the basis of V-shaped proper scoring rules than $V_{2}$, then $V_{1}$ is Blackwell more informative than $V_{2}$.
-Proposition B.3 follows from the fact that any best-responding payoff can be linearly decomposed into the payoff on V-shaped scoring rules.
+Proposition˜B.3 shows that if $V_{1}$ achieves a higher information value on the basis of V-shaped proper scoring rules than $V_{2}$, then $V_{1}$ is Blackwell more informative than $V_{2}$.
+Proposition˜B.3 follows from the fact that any best-responding payoff can be linearly decomposed into the payoff on V-shaped scoring rules.
 Proposition B.3 (Hu and Wu 2024).
@@
 $\operatorname{ACIV}^{\pi,S_{\mu}}(V_{1};D^{b})\geq\operatorname{ACIV}^{\pi,S_{\mu}}(V_{2};D^{b}),\forall\mu\in(0,1)$
-This definition allows us to rank signals (or sets of signals) without needing to commit to a specific payoff function. We present a use case in Appendix E.
+This definition allows us to rank signals (or sets of signals) without needing to commit to a specific payoff function. We present a use case in Appendix˜F.
 Appendix C Observational Study on the Effect of ACIV on Human-AI Team Performance
-In this section, we present the results of an observational study on the effect of $\operatorname{ACIV}$ on human-AI team performance using the dataset from Vodrahalli et al. [2022].
+In this section, we present the results of an observational study on the effect of $\operatorname{ACIV}$ on human-AI team performance using the dataset from Vodrahalli et al. (2022).
 Data description.
-We experiment with the publicly available Human-AI Interactions dataset [Vodrahalli et al., 2022].
+We experiment with the publicly available Human-AI Interactions dataset (Vodrahalli et al., 2022).
 The dataset comprises 34,783 unique predictions from 1,088 different human participants on four different binary prediction tasks (“Art”, “Sarcasm”, “Cities” and “Census”).
@@
 Results.
-We present the $\operatorname{ACIV}$ of the AI confidence values ($d^{AI}$) over human decisions ($d^{H}$) measured by the payoff function as the mean squared error (MSE) in Figure 3, and show the reduction of the mean squared error (MSE) of the human-AI team over the human-alone baseline in Table 1.
+We present the $\operatorname{ACIV}$ of the AI confidence values ($d^{AI}$) over human decisions ($d^{H}$) measured by the payoff function as the mean squared error (MSE) in Figure˜4, and show the reduction of the mean squared error (MSE) of the human-AI team over the human-alone baseline in Table˜2.
 We find that between these four tasks, the $\operatorname{ACIV}$ of the AI over human decisions predicts the improvement of the human-AI team performance over the human-alone baseline.
@@
 $\nu=\frac{(\frac{s_{1}^{2}}{N_{1}}+\frac{s_{2}^{2}}{N_{2}})^{2}}{\frac{s_{1}^{2}}{N_{1}^{2}(N_{1}-1)}+\frac{s_{2}^{2}}{N_{2}^{2}(N_{2}-1)}}$ (10)
-where $s_{1}$ and $s_{2}$ are the standard deviations of the two samples, and $N_{1}$ and $N_{2}$ are the sizes of the two samples (see sample sizes and standard deviations in Table 3).
+where $s_{1}$ and $s_{2}$ are the standard deviations of the two samples, and $N_{1}$ and $N_{2}$ are the sizes of the two samples (see sample sizes and standard deviations in Table˜5).
 We used Bonferroni correction as the multiple comparisons correction to control the overall type I error rate.
-The results are shown in Table 2.
-Table 2: 95% confidence intervals and significance test results (p-values) for the reduction of APE with different AI models and explanations in Figure 1.
-P(ILIV-SHAP + SHAP
-¿ SHAP)
-P(ILIV-SHAP + SHAP
-¿ No explanation)
-P(AI1 ¿ AI2)
-P(AI1 + ILIV-SHAP + SHAP
-¿ AI2 + ILIV-SHAP + SHAP)
-Reduction of APE with 95% confidence intervals
+The results are shown in Table˜3 and Table˜4.
+Table 3: 95% confidence intervals for the reduction of APE with different AI models and explanations in Figure˜2.
 No explanation SHAP ILIV-SHAP + SHAP
 AI1
-$5.95\%$
-$5.49\%,6.40\%]$
-$6.06\%$
-$5.65\%,6.47\%]$
-$7.14\%$
-$6.70\%,7.58\%]$
-0.003 / 0.001
-0.311 / ¡0.001
+$5.96\%$
+$5.50\%,6.42\%]$
+$5.88\%$
+$5.47\%,6.28\%]$
+$6.94\%$
+$6.50\%,7.38\%]$
 AI2
-$6.24\%$
-$5.79\%,6.68\%]$
-$6.29\%$
-$5.80\%,6.78\%]$
-$5.47\%$
-$4.95\%,5.98\%]$
-0.138 / 0.159
+$6.33\%$
+$5.90\%,6.76\%]$
+$6.22\%$
+$5.71\%,6.72\%]$
+$5.31\%$
+$4.80\%,5.83\%]$
+Table 4: P-values for significance tests of reduction in APE between experimental conditions (after Bonferroni correction).
+Null Hypothesis ($H_{0}$) P-value
+AI1 + ILIV-SHAP + SHAP < AI1 + SHAP 0.002
+AI1 + ILIV-SHAP + SHAP < AI1 + No explanation 0.010
+AI1 < AI2 0.555
+AI1 + ILIV-SHAP + SHAP < AI2 + ILIV-SHAP + SHAP <0.001
 Table 5: Sample sizes and standard deviations of the reduction of APE for each experimental condition
@@
 Total 421 5,052 0.0678
-Appendix E Demonstration I: Model Comparison on Chest Radiograph Diagnosis
-We apply our framework to a well-known cardiac dysfunction diagnosis task [Rajpurkar et al., 2018, Tang et al., 2020, Shreekumar, 2025].
+Appendix E Cognitive Loads and Anchoring Effects in Experiment
+We also examine the spent time on task and the degree of anchoring on the human versus AI decisions by looking at the distance between human-AI team decisions and AI or human-alone decisions.
+Time spent.
+Table˜6 shows that there is no significant increase in the duration of the experiment for ILIV-SHAP (23.3 [13.1, 32.1] minutes for AI1 with ILIV-SHAP and 22.2 [11.9, 28.0] minutes for AI2 with ILIV-SHAP, where the square brackets show the 25% and 75% quantiles).
+Anchoring effects.
+Table˜6 also shows that ILIV-SHAP does not increase anchoring on the AI or human alone decisions, while the presence of SHAP alone tends to make the participants anchor more on the AI model.
+With no explanation, the participants anchor more on their own decisions in the first round.
+Table 6: The number of participants (after filtering based on the criteria in the pre-registration), the mean duration of the experiment, and the anchoring on AI and human for each condition.
+Anchoring on AI is calculated as the mean of the absolute difference between the human-AI team’s prediction and the AI’s prediction normalized by the actual sale price, $|d^{\text{Human-AI}}-d^{\text{AI}}|/\omega$, and anchoring on human is calculated as the mean of the absolute difference between the human-AI team’s prediction and the human’s prediction normalized by the actual sale price, $|d^{\text{Human-AI}}-d^{\text{Human}}|/\omega$.
+The square brackets indicate the 25th and 75th percentiles.
+Condition # of participants Mean Duration (minutes) Anchoring on AI ($\downarrow$) Anchoring on Human ($\downarrow$)
+AI1 + ILIV-SHAP+SHAP 70 23.3 [13.1, 32.1] 0.303 [0.0277, 0.253] 0.404 [0.0746, 0.491]
+AI1 + SHAP 79 21.3 [12.0, 27.9] 0.211 [0.0201, 0.222] 0.427 [0.0721, 0.533]
+AI1 + No Explanation 69 22.4 [14.5, 28.4] 0.226 [0.0276, 0.258] 0.367 [0.0740, 0.483]
+AI2 + ILIV-SHAP+SHAP 67 22.2 [11.9, 28.0] 0.206 [0.0273, 0.260] 0.458 [0.0840, 0.613]
+AI2 + SHAP 62 23.1 [13.2, 27.0] 0.203 [0.0215, 0.224] 0.456 [0.0777, 0.602]
+AI2 + No Explanation 74 22.5 [12.7, 30.5] 0.259 [0.0265, 0.260] 0.434 [0.0611, 0.501]
+Appendix F Demonstration I: Model Comparison on Chest Radiograph Diagnosis
+We apply our framework to a well-known cardiac dysfunction diagnosis task (Rajpurkar et al., 2018; Tang et al., 2020; Shreekumar, 2025).
 We demonstrate how our framework can be used in model evaluation for analyzing how much complementary information value a set of possible AI models offers to the radiology reports written by experts.
@@
 We list the AI predictions alone to show the AI-complementary information value offered by human decisions.
-E.1 Data and Model
-We use data from the MIMIC dataset [Goldberger et al., 2000], which contains anonymized electronic health records from Beth Israel Deaconess Medical Center (BIDMC), a large teaching hospital in Boston, Massachusetts, affiliated with Harvard Medical School.
-Specifically, we utilize chest x-ray images and radiology reports from the MIMIC-CXR database [Johnson et al., 2019] merged with patient and visit information from the broader MIMIC-IV database [Johnson et al., 2023].
+F.1 Data and Model
+We use data from the MIMIC dataset (Goldberger et al., 2000), which contains anonymized electronic health records from Beth Israel Deaconess Medical Center (BIDMC), a large teaching hospital in Boston, Massachusetts, affiliated with Harvard Medical School.
+Specifically, we utilize chest x-ray images and radiology reports from the MIMIC-CXR database (Johnson et al., 2019) merged with patient and visit information from the broader MIMIC-IV database (Johnson et al., 2023).
 The payoff-related state, cardiac dysfunction $\omega\in\{0,1\}$, is coded based on two common tests, the NT-proBNP and the troponin, using the age-specific cutoffs from Mueller et al.
-[2019] and Heidenreich et al.
-[2022].
-We label the radiology reports by a rule-based tool [Irvin et al., 2019] and use the labels as the human decisions (without AI assistance) in the diagnosis task to solve the problem of computational feasibility with high-dimensional textual reports.
+(2019) and Heidenreich et al.
+(2022).
+We label the radiology reports by a rule-based tool (Irvin et al., 2019) and use the labels as the human decisions (without AI assistance) in the diagnosis task to solve the problem of computational feasibility with high-dimensional textual reports.
 The labels are represented by the symptoms as positive, negative, or uncertain, i.e., $d\in\{+,?,-\}$.
-We fine-tuned five deep-learning models on the cardiac dysfunction diagnosis task, VisionTransformer [Alexey, 2020], SwinTransformer [Liu et al., 2021], ResNet [He et al., 2016], Inception-v3 [Szegedy et al., 2016], and DenseNet [Huang et al., 2017].
+We fine-tuned five deep-learning models on the cardiac dysfunction diagnosis task, VisionTransformer (Alexey, 2020), SwinTransformer (Liu et al., 2021), ResNet (He et al., 2016), Inception-v3 (Szegedy et al., 2016), and DenseNet (Huang et al., 2017).
 Our training set contains 12,228 images, and the validation set contains 6,115 images.
@@
 The full-information bound is defined as the expected payoff of a rational DM who has access to all signals, human label from radiology report and predictions from five AI models.
-E.2 Results
+F.2 Results
 Can the AI models complement human judgment?
-We first analyze the agent-complementary information values in Figure 4, using the Brier score as the payoff function.
+We first analyze the agent-complementary information values in Figure˜5, using the Brier score as the payoff function.
 We find that all AI models provide complementary information value to the aforementioned human judgment.
-As shown in Figure 4 (comparison between $\mathrm{R}^{\pi,S}(D^{\text{Human}}\cup D^{\text{AI}})$ and $\mathrm{R}^{\pi,S}(D^{\text{Human}})$), all AI models capture at least $20\%$ of the total available information value (across all AI model and human decisions) that is not exploited by human decisions.
+As shown in Figure˜5 (comparison between $\mathrm{R}^{\pi,S}(D^{\text{Human}}\cup D^{\text{AI}})$ and $\mathrm{R}^{\pi,S}(D^{\text{Human}})$), all AI models capture at least $20\%$ of the total available information value (across all AI model and human decisions) that is not exploited by human decisions.
 This motivates deploying an AI to assist humans in this scenario.
-In the other direction, the human decisions also provide complementary information to all AI models, comparing $\mathrm{R}^{\pi,S}(D^{\text{Human}})$ with $\mathrm{R}^{\pi,S}(D^{\text{AI}})$ in Figure 4.
+In the other direction, the human decisions also provide complementary information to all AI models, comparing $\mathrm{R}^{\pi,S}(D^{\text{Human}})$ with $\mathrm{R}^{\pi,S}(D^{\text{AI}})$ in Figure˜5.
 This observation might inspire, for example, further investigation of the information the humans can access to that is not represented in AI training data.
 Which AI model offers the most decision-relevant information over human judgments?
-Figure 4 shows that VisionTransformer contains slightly higher information value than the other models, and Inception v3 contains slightly lower information value than the other models.
+Figure˜5 shows that VisionTransformer contains slightly higher information value than the other models, and Inception v3 contains slightly lower information value than the other models.
 We assess the stability of VisionTransformer’s superiority over the other AI models across many possible losses to test if there is a Blackwell ordering of models.
-By Proposition B.3, we test the payoff of models on all V-shaped scoring rules, shown in Figure 6.
+By Proposition˜B.3, we test the payoff of models on all V-shaped scoring rules, shown in Figure˜7.
 Across all the V-shaped payoff functions, we find that VisionTransformer is Blackwell more informative and Inception v3 is Blackwell less informative than all other models.
 The VisionTransformer achieves a higher information value on all V-shaped scoring rules, implying a higher information value on all decision problems.
-Appendix F Demonstration II: Behavioral Analysis on Deepfake Detection
-We apply our framework to analyze a deepfake video detection task [Dolhansky et al., 2020], where participants are asked to judge whether a video was created by generative AI, including with the assistance of an AI model.
+Appendix G Demonstration II: Behavioral Analysis on Deepfake Detection
+We apply our framework to analyze a deepfake video detection task (Dolhansky et al., 2020), where participants are asked to judge whether a video was created by generative AI, including with the assistance of an AI model.
 Figure 6: Information value calculated under our framework in the information model defined by the experiment of Groh et al.
-[2022].
+(2022).
 Basic signals include the seven video level features and three types of agent decisions.
@@
 All the payoffs are calculated by $\mathrm{R}^{\pi,S}(\cdot)$, where $\cdot$ is the signal on the y-axis.
-F.1 Data and Model
+G.1 Data and Model
 We define the information model on the experiment data of Groh et al.
-[2022].
+(2022).
 Non-expert participants (n=5,524) were recruited through Prolific and asked to examine the videos.
@@
 They first reviewed the video and reported an initial decision ($D^{\text{Human}}$) without access to the AI model.
-Then, in a second round, they were provided with the recommendation ($D^{\text{AI}}$) of a multitask cascaded convolutional neural network [Zhang et al., 2016], with estimated $65\%$ accuracy on holdout data, and chose whether to change their initial decision.
+Then, in a second round, they were provided with the recommendation ($D^{\text{AI}}$) of a multitask cascaded convolutional neural network (Zhang et al., 2016), with estimated $65\%$ accuracy on holdout data, and chose whether to change their initial decision.
 This produced a final decision ($D^{\text{Human-AI}}$).
@@
 This choice differs from the mean absolute error used by Groh et al.
-[2022], but again we use the quadratic score because it is a proper scoring rule where truthfully reporting the belief maximizes the score.
+(2022), but again we use the quadratic score because it is a proper scoring rule where truthfully reporting the belief maximizes the score.
 We identify a set of features that were implicitly available to all three agents (human, AI, and human-AI).
 Because the video signal is high dimensional, we make use of seven video-level features using manually coded labels by Groh et al.
-[2022]: graininess, blurriness, darkness, presence of a flickering face, presence of two people, presence of a floating distraction, and the presence of an individual with dark skin, all of which are labeled as binary indictors.
+(2022): graininess, blurriness, darkness, presence of a flickering face, presence of two people, presence of a floating distraction, and the presence of an individual with dark skin, all of which are labeled as binary indictors.
 These are the basic signals in our framework.
 We estimate the data-generating process $\pi$ using the realizations of signals, state, first-round human decisions, AI predictions, and second-round human-AI decisions.
-The no-information bound is the same as Appendix E while the full-information bound is defined as the expected payoff of a rational DM who has access to all video-level features and three agents’ decisions.
-F.2 Results
+The no-information bound is the same as Appendix˜F while the full-information bound is defined as the expected payoff of a rational DM who has access to all video-level features and three agents’ decisions.
+G.2 Results
 How much decision-relevant information do agents’ decisions offer?
 We first compare the information value of the AI predictions to that of the human decisions in the first round (without AI assistance).
-Figure 5(a) shows that AI predictions provide about $65\%$ of the total possible information value over the no-information baseline, while human decisions only provide about $15\%$.
+Figure˜6(a) shows that AI predictions provide about $65\%$ of the total possible information value over the no-information baseline, while human decisions only provide about $15\%$.
 Because the no information baseline, $0.75$, is equivalent to a random guess drawn from $\text{Bernoulli}(0.5)$, human decisions are only weakly informative for the problem.
@@
 This is consistent with the findings of Guo et al.
-[2024] that humans are bad at distinguishing when AI predictions are correct.
+(2024) that humans are bad at distinguishing when AI predictions are correct.
 How much additional decision-relevant information do the available features offer over agents’ decisions?
@@
 This describes the additional information value in the signal after conditioning on the existing information in the agents’ decisions.
-As shown on the fifth row in Figure 5, the presence of a flickering face offers larger $\operatorname{ACIV}$ over human decisions than over AI predictions, meaning that human decisions could improve by a greater amount if they were to incorporate this information.
-Meanwhile, as shown on the fourth row in Figure 5, the presence of an individual with dark skin offers larger $\operatorname{ACIV}$ over AI predictions than over human decisions, suggesting that humans make greater use of this information.
+As shown on the fifth row in Figure˜6, the presence of a flickering face offers larger $\operatorname{ACIV}$ over human decisions than over AI predictions, meaning that human decisions could improve by a greater amount if they were to incorporate this information.
+Meanwhile, as shown on the fourth row in Figure˜6, the presence of an individual with dark skin offers larger $\operatorname{ACIV}$ over AI predictions than over human decisions, suggesting that humans make greater use of this information.
 This suggests that the AI and human rely on differing information to make their initial predictions, where the AI relies more on information associated with the presence of a flickering face while human participants rely more on information associated with the presence of an individual with dark skin.
 By comparing the $\operatorname{ACIV}$s of different signals over human decisions and human-AI decisions, we also find that simply displaying AI predictions to humans did not lead to the AI-assisted humans exploiting the observed signals in their decisions.
-As shown in Figure 5, with the assistance of AI, the $\operatorname{ACIV}$s of all signals over the human-AI teams’ decisions do not change much compared to the $\operatorname{ACIV}$s over human decisions, with the exception of a slight improvement in the presence of a flickering face.
+As shown in Figure˜6, with the assistance of AI, the $\operatorname{ACIV}$s of all signals over the human-AI teams’ decisions do not change much compared to the $\operatorname{ACIV}$s over human decisions, with the exception of a slight improvement in the presence of a flickering face.
 This finding further confirms the hypothesis that humans are simply relying on AI predictions without processing the information contained in them.
-Appendix G Robustness Analysis in Demonstration I
+Appendix H Sensitivity Analysis of the Rational Belief Estimator
+In this section, we first present a theoretical analysis on the quality of the ACIV estimator by connecting the calibration error of the given $\mathcal{A}$.
+Then, we present an empirical analysis on the chest X-ray diagnosis demonstration with different algorithms (linear regression (LR), gradient boosting methods (GBM), and neural network (NN)) as $\mathcal{A}$.
+H.1 Theoretical Analysis
+In this section, we theoretically show how the quality of the ACIV estimator (Algorithm˜1) is related to the calibration error of the rational belief estimator $a=\mathcal{A}(\{v_{i},d^{b}_{i},\omega_{i}\}_{i=1}^{n})$.
+We use the regret of the decision maker to represent the quality of the estimation on the rational decision rules, i.e., how much payoff can be improved if we correct the decisions in hindsight [^12] We use the notion of regret to quantify the quality of our ACIV estimator because the perfect decisions are usually unidentifiable with finite signals..
+Because the ACIV is defined as the improvement of the best-attainable performance with the signal and the agent decision over the one with the agent decision alone, the decisions $\hat{d}^{r}$ and $\hat{d}^{rb}$ in Algorithm˜1 should have as small regret as possible.
+We first derive our definition of the regret of Algorithm˜1 using the swap regret by Roth and Shi (2024).
+Definition H.1 (Swap Regret (Roth and Shi, 2024)).
+Given a set of observations $\{(v_{i},d^{b}_{i},\omega_{i})\}_{i=1}^{n}$, a decision rule $d(\cdot)$, and a decision task with payoff function $S$, the swap regret of the DM is:
+$\textsc{Swap}_{S}(d,\{(v_{i},d^{b}_{i},\omega_{i})\}_{i=1}^{n})=\frac{1}{n}\max_{\sigma:\mathbf{D}\rightarrow\mathbf{D}}\sum_{i=1}^{n}\left[S(\sigma(d(v_{i},d^{b}_{i})),\omega_{i})-S(d(v_{i},d^{b}_{i}),\omega_{i})\right]$ (11)
+[…]
+Definition H.2 (Regret of the ACIV estimator).
+Given an empirical distribution as the data-generating process, $\pi=\text{Uniform}(\{v_{i},d^{b}_{i},\omega_{i}\}_{i=1}^{n})$, the estimated decision rules $\hat{d}^{r}:\mathbf{V}\times\mathbf{D}\rightarrow\mathbf{D}$ and $\hat{d}^{rb}:\mathbf{D}\rightarrow\mathbf{D}$ from Algorithm˜1, and a decision task with payoff function $S$, the regret of the ACIV estimator is:
+$\textsc{RegACIV}_{S,\pi}(\hat{d}^{r},\hat{d}^{rb};V,D^{b})=\operatorname{ACIV}^{\pi,S,d^{*r},d^{*rb}}(V;D^{b})-\operatorname{ACIV}^{\pi,S,\hat{d}^{r},\hat{d}^{rb}}(V;D^{b})$ (12)
+where $\operatorname{ACIV}^{\pi,S,\hat{d}^{r},\hat{d}^{rb}}$ denotes the $\operatorname{ACIV}$ estimated under the rational decision rules $\hat{d}^{r}$ and $\hat{d}^{rb}$.
+$d^{*r}$ and $d^{*rb}$ are the optimal decision rules that we can get from $\hat{d}^{r}$ and $\hat{d}^{rb}$:
+$d^{*r}(\cdot,\cdot)=\sigma^{r}(\hat{d}^{r}(\cdot,\cdot))\text{ where }$ $\sigma^{r}=\operatorname*{arg\,max}_{\sigma:\mathbf{D}\rightarrow\mathbf{D}}\sum_{i=1}^{n}\left[S(\sigma(\hat{d}^{r}(v_{i},d^{b}_{i})),\omega_{i})\right]$
+$d^{*rb}(\cdot)=\sigma^{rb}(\hat{d}^{rb}(\cdot))\text{ where }$ $\sigma^{rb}=\operatorname*{arg\,max}_{\sigma:\mathbf{D}\rightarrow\mathbf{D}}\sum_{i=1}^{n}\left[S(\sigma(\hat{d}^{rb}(d^{b}_{i})),\omega_{i})\right]$
+Lemma H.3.
+The regret of the ACIV estimator is upper bounded by the swap regret of $\hat{d}^{r}$:
+$\textsc{RegACIV}_{S,\pi}(\hat{d}^{r},\hat{d}^{rb};V,D^{b})\leq\textsc{Swap}_{S}(\hat{d}^{r},\{(v_{i},d^{b}_{i},\omega_{i})\}_{i=1}^{n})$ (13)
+Proof.
+$\textsc{RegACIV}_{S,\pi}(\hat{d}^{r},\hat{d}^{rb};V,D^{b})$ $=\text{\bf E}_{(v,d^{b},\omega)\sim\pi}\![{S(\sigma^{r}(\hat{d}^{r}(v,d^{b})),\omega_{i})}]-\text{\bf E}_{(v,d^{b},\omega)\sim\pi}\![{S(\hat{d}^{r}(v,d^{b}),\omega_{i})}]$- (14)
+$\quad\left(\text{\bf E}_{(d^{b},\omega)\sim\pi}\![{S(\sigma^{rb}(\hat{d}^{r}(d^{b}_{i})),\omega_{i})}]-\text{\bf E}_{(d^{b},\omega)\sim\pi}\![{S(\hat{d}^{r}(d^{b}_{i}),\omega_{i})}]\right)$
+$\leq\text{\bf E}_{(v,d^{b},\omega)\sim\pi}\![{S(\sigma^{r}(\hat{d}^{r}(v,d^{b})),\omega_{i})}]-\text{\bf E}_{(v,d^{b},\omega)\sim\pi}\![{S(\hat{d}^{r}(v,d^{b}),\omega_{i})}]$
+$=\max_{\sigma:\mathbf{D}\rightarrow\mathbf{D}}\left[\frac{1}{n}\sum_{i=1}^{n}\left[S(\sigma(\hat{d}^{r}(v_{i},d^{b}_{i})),\omega_{i})-S(\hat{d}^{r}(v_{i},d^{b}_{i}),\omega_{i})\right]\right]$
+$=\textsc{Swap}_{S}(\hat{d}^{r},\{(v_{i},d^{b}_{i},\omega_{i})\}_{i=1}^{n})$
+Claim H.4 (Kleinberg et al. (2023), Theorem 12).
+Given a decision task with payoff function $S:\mathbf{D}\times\mathbf{\Omega}\rightarrow[0,1]$, if the DM responds by taking $d(v,d^{b})=d^{*}(a(v,d^{b}))$, where $a$ is an estimator of the probability of the payoff state given the signal and action and $d^{*}(p)=\operatorname*{arg\,max}_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim p}\![{S(d,\omega)}]$ is the best response to the probability $p$, the swap regret of the DM is bounded by the expected calibration error (ECE) of the estimator $a$:
+$\textsc{Swap}_{S}(\hat{d}^{r},\{(v_{i},d^{b}_{i},\omega_{i})\}_{i=1}^{n})\leq 2\text{ECE}(a,\{(v_{i},d^{b}_{i},\omega_{i})\}_{i=1}^{n})$ (15)
+Theorem H.5.
+Given a decision task with bounded payoff $S:\mathbf{D}\times\mathbf{\Omega}\rightarrow[M_{1},M_{2}]$, the regret of the ACIV estimator is bounded by the expected calibration error of the estimator $\hat{a}=\mathcal{A}(\{v_{i},d^{b}_{i},\omega_{i}\}_{i=1}^{n})$:
+$\textsc{RegACIV}_{S,\pi}(\hat{d}^{r},\hat{d}^{rb};V,D^{b})\leq 2(M_{2}-M_{1})\textsc{ECE}(\hat{a},\{(v_{i},d^{b}_{i},\omega_{i})\}_{i=1}^{n})$ (16)
+The proof normalizes the payoff in Lemma˜H.3 into $[0,1]$ and then applies ˜H.4.
+Theorem˜H.5 shows that, in Algorithm˜1, when we choose a predictive algorithm $\mathcal{A}$ that yields low ECE, the ACIV estimated by Algorithm˜1 is close to the ACIV under the optimal decision rules.
+H.2 Empirical Analysis
+In this section, we take the chest X-ray diagnosis task in Appendix˜F as an example to empirically analyze the estimation of the full information value by Algorithm˜1 with different modeling approaches (linear regression (LR), gradient boosting methods (GBM), and neural network (NN)).
+Task and data.
+We use the chest X-ray diagnosis task in Appendix˜F as the example.
+The task is to predict the probability of the presence of a disease given a chest X-ray image.
+The decision space is $\mathbf{D}=\Delta[0,1]$ and the payoff space is $\mathbf{\Omega}=\{0,1\}$.
+The payoff function is $S(d,\omega)=1-(d-\omega)^{2}$.
+The signal space is a high-dimensional feature vector of the chest X-ray image.
+We split the dataset that we used in Appendix˜F into a 70/30 train/test split.
+Predictive Models.
+We use the following models:
+  • Linear Regression (LR) from scikit-learn package.
+  • Extreme Gradient Boosting (GBM) from xgboost package with n_estimators = 100, max_depth = 6, and learning_rate = 0.3.
+  • MLP classifier (NN) from scikit-learn package with 2 hidden layers of 256 and 64 neurons each.
+Evaluation Metrics.
+We report the following evaluation metrics for each estimator:
+  • Brier Score: $\frac{1}{n}\sum_{i=1}^{n}(\hat{p}_{i}-\omega_{i})^{2}$
+  • Accuracy: $\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}(\mathbb{1}(\hat{p}_{i}\geq 0.5)=\omega_{i})$
+  • Expected Calibration Error (ECE): $\frac{1}{n}\sum_{i=1}^{n}\left|\hat{p}_{i}-\text{\bf E}\![{\omega|\hat{p}_{i}}]\right|$
+Table 7: Performance of different algorithms for ACIV estimation in the chest X-ray diagnosis task.
+Metrics for the estimators include Brier Score, Accuracy, F1-Score, and Expected Calibration Error (ECE).
+We report the estimated $\operatorname{ACIV}(V;\emptyset)$ for each estimator with Algorithm˜1.
+Model Brier Score $\uparrow$ Accuracy $\uparrow$ ECE $\downarrow$ $\mathbf{ACIV(V;\emptyset)}\uparrow$
+Linear Regression (LR) 0.834 0.775 0.039 0.086
+Gradient Boosting (GBM) 0.841 0.771 0.026 0.108
+Neural Network (NN) 0.853 0.789 0.036 0.0757
+Results.
+The results are shown in Table˜7.
+We can see that the GBM estimator achieves the best performance in terms of ECE, while the NN estimator achieves the best performance in terms of Brier Score and Accuracy.
+This validates our theoretical result: ECE is a good proxy for the regret of the estimation of the ACIV by Algorithm˜1 compared to the Brier Score and Accuracy.
+Appendix I Robustness Analysis in Demonstration I
 Figure 7: Robust analysis for experiment I on all V-shaped payoff functions.
@@
 The subplots are symmetric along the diagonal, e.g., $(1,2)$ subplot and $(2,1)$ subplot display the same distribution with opposite signs.
-Appendix H Screenshots of the Experiment
+Appendix J Screenshots of the Experiment
 Figure 8: Screenshot of the human-alone trials in the first round of the experiment.